### v1

# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

Abstract
Patterns are words with terminals and variables. The language of a pattern is the set of words obtained by uniformly substituting all variables with words that contain only terminals. In their original definition, patterns only allow for multiple distinct occurrences of some variables to be related by the equality relation, represented by using the same variable multiple times. In an extended notion, called relational patterns and relational pattern languages, variables may be related by arbitrary other relations. We extend the ongoing investigation of the main decision problems for patterns (namely, the membership problem, the inclusion problem, and the equivalence problem) to relational pattern languages under a wide range of individual relations. It is shown show that - even for many much simpler or less restrictive relations - the complexity and (un)decidability characteristics of these problems do not change compared to the classical case where variables are related only by equality.
1 Introduction
A pattern (with variables) is a finite word that consists only of symbols from a finite set of (terminal) letters $\Sigma=\{\mathtt{a}_{1},...,\mathtt{a}_{\sigma}\}$ and from an infinite but countable set of variables $X=\{x_{1},x_{2},$...\} with $\Sigma\cap X=\emptyset$. By applying terminal preserving morphisms, called substitutions, to a pattern, we can obtain words that consist only of terminal letters. The language of a pattern is then just the set of all words that can be obtained through the application of arbitrary substitutions.
Originally, pattern languages as defined and introduced by Angluin [Angluin1980] only considered words that are obtained by what is now called non-erasing substitutions, i.e., substitutions where each variable is substituted by a word of at least length one. Hence, these languages are now also called non-erasing (NE) pattern languages. Shinohara [Shinohara1983] extended this notion by allowing for the empty word to be used to substitute a variable, introducing so called extended/erasing (E) pattern languages.
For example, consider the pattern $\alpha=x_{1}\mathtt{a}\mathtt{a}x_{1}\mathtt{b}x_{2}$. Using the non-erasing substitution $h$ that sets $h(x_{1})=\mathtt{b}\mathtt{b}$ and $h(x_{2})=\mathtt{a}$, we obtain the word $h(\alpha)=\mathtt{b}\mathtt{b}\mathtt{a}\mathtt{a}\mathtt{b}\mathtt{b}\mathtt{b}\mathtt{a}$. If we also considered the E-pattern language of $\alpha$, we could set all variables to the empty word and obtain the word $\mathtt{a}\mathtt{a}\mathtt{b}$ which we could not obtain in the NE pattern language of $\alpha$.
As patterns form a natural and compact device for the definition of formal languages and due to their practical and simple definition, they occur in numerous areas in computer science and discrete mathematics. Among others, these include for example the areas of unavoidable patterns in words [Jiang1994, lothaire1997], word equations [lothaire1997], algorithmic learning theory [Angluin1980, FERNAU201844, Shinohara1995], database theory [FreydenbergerP21, SchmidSchweikardtPODS2022], or the theory of extended regular expressions with back references[FREYDENBERGER20191]. Furthermore, many practical areas such as machine learning, database systems, or bio-informatics deal with the question of identifying or describing patterns in (sets of) strings [Mousawi2024], for example in the analysis of protein data [Arikawa1993], in pattern matching [10.1007/978-3-642-03784-9_29], or the design of algorithms for program synthesis [10.1145/4472.4476].
Three main decision problems emerge when it comes to the investigation of patterns and pattern languages. These are the general membership problem and its variations [gawrychowski_et_al:LIPIcs.MFCS.2021.48, Manea2022, Fleischmann2023], the inclusion problem, and the equivalence problem, all of which are considered in the non-erasing (NE) as well as erasing (E) cases. The membership problem determines whether a word belongs to a pattern’s language. It has been shown to be generally NP-complete in the non-erasing as well as erasing cases [Angluin1980, Jiang1994]. The inclusion problem checks whether the language of a pattern is included as a subset in another pattern’s language. It has been shown to be generally undecidable by Jiang et al. [Jiang1995]. Their result has been improved by Freydenberger and Reidenbach in [Freydenberger2010] as well as by Bremer and Freydenberger in [BREMER201215], where it has been shown to be undecidable for all alphabets $\Sigma$ of fixed size $|\Sigma|\geq 2$, for erasing and non-erasing pattern languages. The equivalence problem determines whether the languages of two given patterns are equal. In Angluin’s original work [Angluin1980], it has been shown to be trivially decidable for all NE pattern languages. For erasing pattern languages, however, the general decidability of this problem has been the most significant open problem in the field of decision problems for patterns [Jiang1995, Reidenbach2004-1, Reidenbach2004-2, Ohlebusch1996, Reidenbach2007]. In the special case of terminal-free patterns (containing only variables), the membership problem remains unchanged in both erasing and non-erasing cases, as does the non-erasing equivalence problem. However, inclusion for terminal-free E-pattern languages is NP-complete, as it reduces to checking for a morphism between patterns [Jiang1995, DBLP:journals/ipl/EhrenfeuchtR79a], while equivalence is decidable. In contrast, Saarela [saarela:LIPIcs.ICALP.2020.140] showed that inclusion for terminal-free NE pattern languages is undecidable over a binary alphabet.
Various extensions to the notion of patterns and pattern languages have been introduced over time either to get closer to an answer for the remaining open problems or to obtain additional expressibility that is usable in some practical context. Some examples are the bounded scope coincidence degree, patterns with bounded treewidth, $k$-local patterns, or strongly-nested patterns (see [Day2018] and references therein). Koshiba [Koshiba1995] introduced so called typed patterns where the substitutions of variables can be restricted by arbitrary recursive languages, so called types. This has been recently extended by Geilke and Zilles in [Geilke2011] to the notion of relational patterns and relational pattern languages. Here, other relations than equality between variables may be used to restrict the valid substitution of patterns.
This model enhances expressiveness, includes all previous models, but also includes arbitrary other constructions that involve relations between variables in a pattern. Consider, for example, the pattern $\beta=x_{1}\mathtt{c}\mathtt{c}x_{2}$. Now let’s say that the variables $x_{1}$ and $x_{2}$ should be the substituted by the reversal of each other, i.e., $(h(x_{1}),h(x_{2}))\in r_{rev}$, where $h$ is a substitution and $r_{rev}$ is the reversal relation defined by $r_{rev}:=\{\ (w,w^{R})\mid w\in\Sigma^{*}\ \}$ (where $w^{R}$ is defined to be the reversal of a word $w$). Then, e.g., a substitution $h$ with $h(x_{1})=\mathtt{a}\mathtt{b}$ and $h(x_{2})=\mathtt{b}\mathtt{a}$ is said to be valid regarding the pattern $\beta$, we get $h(\beta)=\mathtt{a}\mathtt{b}\mathtt{c}\mathtt{c}\mathtt{b}\mathtt{a}$, and we say that $h(\beta)$ is in the relational pattern language of $\beta$ under the relational constraints containing the pair that relates $x_{1}$ and $x_{2}$ by $r_{rev}$. If we considered a relation $h^{\prime}$, however, in which $h^{\prime}(x_{1})=\mathtt{a}\mathtt{b}$ and also $h^{\prime}(x_{2})=\mathtt{a}\mathtt{b}$, then clearly $h^{\prime}(x_{1})\neq h^{\prime}(x_{2})^{R}$ and therefore $h^{\prime}$ is not a valid substitution in this case, resulting in $h^{\prime}(\beta)$ not being in that relational pattern language of $\beta$.
Geilke and Zilles [Geilke2011] show some properties of relational pattern languages in general. Holte et al. [pmlr-v167-holte22a] investigated relational pattern languages under the equal length or reversal relations, respectively. They show that inclusion, by that also equivalence, under equal length is decidable in polynomial time for ternary alphabets. In a very recent paper, Mousawi and Zilles [Mousawi2024] extended this from the perspective of learning theory, introducing the notion of positive characteristic sets. Their results carry over and it is shown that inclusion, therefore also equivalence, in the erasing and non-erasing cases under equal length is decidable in polynomial time for binary alphabets. Also, for the reversal relation, their result implies that in the non-erasing case, equivalence is decidable. Additionally, also in two recent papers, in the cases of two special kinds of relations/constraints, namely regular constraints [Nowotka2024] or length constraints[Nowotka2025], Nowotka and Wiedenhöft have shown, among other results, that the still open equivalence problem for erasing pattern languages actually becomes undecidable if these constraints are allowed in addition to equality between variables. This sets a first upper bound of constraints necessary to obtain undecidability for that problem. In comparison to many other relations, however, these constraints appear to be very strong, hence leaving the question of what happens in cases of weaker relations.
In total, we see that the field of relational pattern languages remains largely unexplored, leaving room to study various other or weaker relations restricting variable substitutions. This paper considers equality ($r_{=}$), equal length ($r_{|w|}$), subsequence ($r_{ssq}$), abelian equivalence ($r_{ab}$), alphabet permutation ($r_{perm}$), reversal ($r_{rev}$), (nonempty) commutation ($r_{com^{+}}$, $r_{com^{*}}$), and the star relation ($r_{*}$). We study the membership, inclusion, and equivalence problems under these relations, providing a comprehensive overview of many cases (Table 3) and showing that hardness and (un)decidability persist even for weaker relations than equality. In particular, it can be shown that the membership problem is NP-complete for all considered relations in all cases. Regarding the inclusion problem, it is shown that undecidability persists even for the relations $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ in the erasing and non-erasing cases. Finally, in the case of the equivalence problem, we show decidability in polynomial time in the non-erasing case for a class of relations with specific properties, containing $r_{ab}$ and $r_{com^{+}}$.
Section 2 covers all the necessary notation and models needed to follow the results of the paper. In Section 3, we present the results for the membership problem. This is followed in Section 4 with an overview of the results for the inclusion problem. After that, in Section 5, we briefly discuss the results regarding the equivalence problem. Finally, in Section 6, we summarize the paper and provide a brief outlook with regards to open questions. Due to space constraints, some proofs were moved into the appendix. The corresponding statements are marked with (✂).
2 Preliminaries
Let $\mathbb{N}$ denote the natural numbers $\{1,2,3,\dots\}$ and let $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $n,m\in\mathbb{N}$ set $m,n]:=\{k\in\mathbb{N}\mid m\leq k\leq n\}$. Denote $[n]:=[1,n]$ and $n]_{0}:=[0,n]$. An alphabet $\Sigma$ is a non-empty finite set whose elements are called letters. A word is a finite sequence of letters from $\Sigma$. Let $\Sigma^{*}$ be the set of all finite words over $\Sigma$ and set $\Sigma^{+}:=\Sigma^{*}\setminus\{\varepsilon\}$. We call the number of letters in a word $w\in\Sigma^{*}$ length of $w$, denoted by $|w|$. Therefore, we have $|\varepsilon|=0$. Denote $\Sigma^{k}:=\{w\in\Sigma^{*}\mid|w|=k\}$. For $w\in\Sigma^{*}$, let $w[i]$ denote $w$’s $i^{th}$ letter for all $i\in[|w|]$. For compactness reasons, we denote $w[i]\cdots w[j]$ by $w[i\cdots j]$ for all $i,j\in[|w|]$ with $i<j$. Set $\operatorname{alph}(w):=\{\mathtt{a}\in\Sigma\mid\exists i\in[|w|]:w[i]=\mathtt{a}\}$ as $w$’s alphabet.
Let $X:=\{x_{1},x_{2},$...\} be a countable set of variables such that $\Sigma\cap X=\emptyset$. A pattern is then a non-empty, finite word over $\Sigma\cup X$. The set of all patterns over $\Sigma\cup X$ is denoted by $Pat_{\Sigma}$. For example, $x_{1}\mathtt{a}x_{2}\mathtt{b}\mathtt{a}x_{2}x_{3}$ is a pattern over $\Sigma=\{\mathtt{a},\mathtt{b}\}$ with $x_{1},x_{2},x_{3}\in X$. For a pattern $\alpha\in Pat_{\Sigma}$, let $\operatorname{var}(\alpha):=\{\ x\in X\ |\ |p|_{x}\geq 1\ \}$ denote the set of variables occurring in $p$. A substitution of $\alpha$ is a morphism $h:(\Sigma\cup X)^{*}\to\Sigma^{*}$ such that $h(\mathtt{a})=\mathtt{a}$ for all $\mathtt{a}\in\Sigma$ and $h(x)\in\Sigma^{*}$ for all $x\in X$. If we have $h(x)\neq\varepsilon$ for all $x\in\operatorname{var}(p)$, we call $h$ a non-erasing substitution for $\alpha$. Otherwise $h$ is an erasing substitution for $\alpha$. The set of all substitutions w.r.t. $\Sigma$ is denoted by $H_{\Sigma}$. If $\Sigma$ is clear from the context, we may write just $H$. Given a pattern $\alpha\in Pat_{\Sigma}$, it’s erasing pattern language $L_{E}(\alpha)$ and its non-erasing pattern language $L_{NE}(\alpha)$ are defined respectively by
$L_{E}(\alpha)$ $:=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha)\},\text{ and }$
$L_{NE}(\alpha)$ $:=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\}.$
Let $RegPat_{\Sigma}:=\{\alpha\in Pat_{\Sigma}\mid x_{i}\in\operatorname{alph}(\alpha)\cap X,|\alpha|_{x_{i}}=1\}$ denote the set of all regular patterns, i.e., the set of all patterns where each variable occurs only once. Let $R$ be some set of relations and denote for each $n\in\mathbb{N}$ by $R_{n}\subset R$ the subset of $n$-ary relations in $R$. Define $Rel_{R}:=\{(r,x_{i_{1}},...,x_{i_{n}})\mid r\in R_{n},x_{i_{1}},...,x_{i_{n}}\in X,i_{1},...i_{n}\in\mathbb{N},n\in\mathbb{N}\}$ to be the set of all combinations to relate variables in $X$ over the set of relations $R$. Then, a pair $(\alpha,r_{\alpha})\in RegPat_{\Sigma}\times Rel_{R}$ is a relational pattern over the terminal alphabet $\Sigma$ and the set of relations $R$. We denote the set of all relational patterns over $\Sigma$ and $R$ by $\mathtt{RelPat}_{\Sigma,R}$. For some $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $h\in H$, we say that $h$ is a $r_{\alpha}$-valid substitution if for all $(r^{\prime},x_{i_{1}},...,x_{i_{n}})\in r_{\alpha}$ we have $(h(x_{i_{1}}),...,h(x_{i_{n}}))\in r^{\prime}$, i.e., all variables are substituted such that the resulting terminal words are related to each other regarding the relation $r^{\prime}$. The set of all $r_{\alpha}$-valid substitutions w.r.t. $\Sigma$ is denoted by $H_{\Sigma,r_{\alpha}}$. If $\Sigma$ and $r_{\alpha}$ are clear from the context, we may also just write $H_{r_{\alpha}}$ or $H$. We extend the notion of pattern languages by the following. For any $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ we denote by
$L_{E}(\alpha,r_{\alpha}):=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha),h\text{ is }r_{\alpha}\text{-valid}\ \}$
the erasing relational pattern language of $(\alpha,r_{\alpha})$ and by
$L_{NE}(\alpha,r_{\alpha}):=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha),h\text{ is }r_{\alpha}\text{-valid}\ \}$
the non-erasing relational pattern language of $(\alpha,r_{\alpha})$.
In this paper, we primarily investigate the considered decision problems for relational pattern languages for binary relations, i.e., $|R|=1$ and if $R=\{\tilde{r}\}$, then $\tilde{r}$ is 2-ary/binary. In particular, we are interested in the following selection of specific customary binary relations. Let $\mathcal{R}$ be the set containing all of the following nine relations.
  1. 1.
    $r_{=}$ : equality of words, i.e., $(u,v)\in R_{=}$ iff $u=v$.
  2. 2.
    $r_{|w|}$ : length equality of words, i.e., $(u,v)\in R_{|w|}$ iff $|u|=|v|$.
  3. 3.
    $r_{ssq}$ : subsequence relation, i.e., $(u,v)\in R_{ssq}$ iff $u\in\operatorname{SubSeq}(v)$.
  4. 4.
    $r_{ab}$ : abelian equivalence of words, i.e., $(u,v)\in R_{ab}$ iff $u\equiv_{ab}v$.
  5. 5.
    $r_{perm}$ : alphabet permutation of words, i.e., $(u,v)\in R_{perm}$ iff $u$ can be obtained by $v$ with a permutation on the alphabet of its letters.
  6. 6.
    $r_{rev}$ : reversal of words, i.e., $(u,v)\in R_{rev}$ iff $u=v^{R}$.
  7. 7.
    $r_{com^{*}}$ : commutation between words, i.e., $(u,v)\in R_{com^{*}}$ iff there exists $z\in\Sigma^{*}$ such that $u,v\in\{z\}^{*}$.
  8. 8.
    $r_{com^{+}}$ : nonempty commutation between words, i.e., $(u,v)\in R_{com^{+}}$ iff there exists $z\in\Sigma^{*}$ such that $u,v\in\{z\}^{+}$.
  9. 9.
    $r_{*}$ : word in star of the other word, i.e., $(u,v)\in r_{*}$ iff $u\in\{v\}^{*}$.
In the proofs of the main results regarding the inclusion problem, we make use of two different automata models that have undecidable language emptiness properties. The first one is the so called nondeterministic 2-counter automaton without input (see, e.g., [Iberra1978]). The second one is a very specific universal Turing machine $U$ with 2 symbols and 15 states (see [Neary2009]). As the technical proofs are very lengthy and based on the constructions provided in other papers, the formal definitions of the models can be found in Appendix  0.B.1 and Appendix  0.B.2 respectively. For the intuitions of the proofs provided in the main body, it suffices to know that these respective models have undecidable language emptiness properties, that all terminating/accepted computations are encoded using finite binary words, and that the number of existing and not existing outgoing transitions per state are bound by a constant. This last property results in a finite number of properties of words that are not encodings of accepting computations.
3 The Membership Problem
In this section, we discuss the membership problem for relational pattern languages for the relations in $\mathcal{R}$. In addition to that, we also consider certain special cases of one-sided relations mentioned at the end of [Angluin1980] that do not immediately translate to the framework of relational pattern languages considered here. As to be seen, the membership problem remains NP-complete for all considered cases. We will begin by covering the cases arising for all relations in $\mathcal{R}$.
For the case of $r_{=}$, we recall the original result by Angluin in [Angluin1980] in which it is shown that the membership problem of non-erasing pattern languages is NP-complete. In the erasing case, the same result has been shown by Jiang et al. in [Jiang1994]. The following theorem states the main result of this section. It is the NP-completeness of the membership problem for relational pattern languages over any of the relations in $\mathcal{R}$. The rest of this section serves as its proof.
Theorem 3.1
Let $r\in\mathcal{R}$ and assume $R=\{r\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{X}(\alpha,r_{\alpha})$ for $X\in\{E,NE\}$ is NP-complete.
Indeed, the NP-containment of the membership problems for any of the relations follows by a rather similar argument as can be made for the case of the equality relation $r_{=}$ (see Appendix 0.A for more details).
Proposition 1(✂)
Let $r\in\mathcal{R}$ and $R=\{r\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{X}(\alpha,r_{\alpha})$ for $X\in\{E,NE\}$ is in NP.
The Non-Erasing Cases: To show NP-hardness, we begin with the non-erasing case and start with the relations for which the exact same construction as in [Angluin1980] can be applied. An overview of the original proof together with respective arguments why it works for all these other relations can be found in Appendix 0.A.
Lemma 1(✂)
Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$. In the case of $r=r_{hd,k}$, assume $k\in\mathbb{N}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
That leaves us only with the cases of commutation. Notice that for the non-erasing cases, $r_{com^{*}}$ and $r_{com^{+}}$ result in the same valid substitutions as the empty word is excluded from valid substitutions.
Proposition 2(✂)
Let $r\in\{r_{com^{*},com^{+}}\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
Again, the full technical proof can be found in Appendix 0.A. To give an intuition, this proof also relies on a reduction from the $3\mathtt{-SAT}$ problem and has a similar nature as the proof used for Lemma˜1 which is based on the proof from [Angluin1980]. For the adapted construction to work, we only use two different kinds of factors instead of three. First, a part that handles the variable’s true and false encodings, by setting the factor in $w$ to 1#11\#1 and the factor in $\alpha$ to $u_{i}v_{i}$, for each $3\mathtt{-SAT}$ variable $X_{i}$ (resulting in only either $u_{i}$ or $v_{i}$ to be substituted to contain a #\#). For the parts handling the clauses, we use the factor $1^{10}\#1^{10}\#1^{10}$ in $w$ and the factor $z_{j,1}\ x_{j,1}\ z_{j,2}\ x_{j,2}\ z_{j,3}\ x_{j,3}\ z_{j,4}$ in $\alpha$, where each $x_{j,i}$ corresponds to some variable related to either some $u_{k}$ or $v_{k}$ and each $z_{j_{i}}$. By that, at least one variable in each clause must not contain a #\#, resulting in the reduction to go through. This concludes the non-erasing cases.
The Erasing Cases: Now, we continue with the erasing case. As in the non-erasing cases, for many relations, the existing construction (this time from [Jiang1994]) can be directly applied to show NP-hardness (see Appendix 0.A for some more details).
Lemma 2(✂)
Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$. In the case of $r=r_{hd,k}$, assume $k\in\mathbb{N}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
That leaves us with ++-commutation and *-commutation. First of all, for the case of $r_{com^{+}}$, we see that the proof used to show Proposition˜2 also applies for the erasing case, as two related variables are never substituted to the empty word (see Appendix 0.A for more details).
Proposition 3(✂)
Let $R=\{r_{com^{+}}\}$. Given some word $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
Notice that $r_{com^{*}}$, however, always allows for one of two related variables to be substituted by the empty word. Therefore, we need a slightly different construction for this to work. Based on the proof of Proposition Proposition˜2, we can change the clause encoding in $w$ to only the letter $1$ instead of $1^{10}\#1^{10}\#1^{10}$ and remove all $z$ variables for the construction to work. The basic idea is that is a variable is substituted to contain the letter #\#, it cannot be substituted to a single $1$, reversing the idea used in Proposition˜2 (See Appendix 0.A).
Proposition 4(✂)
Let $R=\{r_{com^{*}}\}$. Given some word $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{E}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
This concludes the NP-hardness results for the membership problem under all relations considered and, thus, the proof of Theorem 3.1. As mentioned in the beginning of this section, we also obtain NP-hardness in the cases of restricting the constructed patterns to one-sided subsequence or one-sided star relations. We obtain the following results by proofs that are similar to the ones used for the previous results. Due to space constraints and their special nature, the discussion of the following result can be found in Appendix 0.A.1.
Theorem 3.2
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x_{i}$ occurs at most once, and, if it does, the related variables $x^{*}$ (resp. $x^{s}$) [^1] For the discussion of restrictions to one-sided relations, we deviate from the notation used in all other parts of this paper and use a notation similar to the one introduced in the end of the work by Angluin[Angluin1980]. may occur multiple times, it is NP-complete to decide whether $w\in L_{X}(\alpha)$, where $X\in\{E,NE\}$.
4 The Inclusion Problem
In this section, we continue with the inclusion problem for relational pattern languages. We recall that the inclusion problem for non-erasing as well as erasing pattern languages over $r_{=}$ is generally undecidable for all alphabets $\Sigma$ of size $|\Sigma|\geq 2$, as shown and refined in  [Jiang1995, Freydenberger2010, BREMER201215]. As we will see in this section, for several of the considered relations in $\mathcal{R}$, we also obtain that this problem is undecidable in the non-erasing or erasing cases.
In fact, if we consider, first, the relations $r_{ssq}$ and $r_{*}$, we observe that we can simulate the exact behavior of $r_{=}$, similar to the NP-hardness proofs in the membership section. Meaning, if we take any two variables $x,y\in\operatorname{var}(\alpha)$, for some relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, and any of the three relations above, we see that relating these variables in both direction effectively forces two variables to be substituted equally in $r_{\alpha}$ valid substitutions, i.e., if $(x,y),(y,x)\in r_{\alpha}$, then $h\in H_{r_{\alpha}}$ if and only if $h(x)=h(y)$. By the fact that this problem is undecidable for $r_{=}$, this results in the following, for which the formal proof can be found in Appendix 0.B.3 for reference.
Corollary 1
[Jiang1995, Freydenberger2010, BREMER201215] Let $\Sigma$ be any finite alphabet with $|\Sigma|\geq 2$. Let $\tilde{r}\in\{r_{=},r_{ssq},r_{*}\}$ and let $R=\{\tilde{r}\}$. Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$.
Now, we continue with the analysis for relations for which we cannot immediately infer the undecidability from such trivial relationships.
4.1 The Non-Erasing Cases
In the non-erasing case, an interesting observation regarding $r_{ab}$ can be found. Reconstructing the reduction from [BREMER201215] in the setting of relational pattern languages, it can be shown that the inclusion problem for non-erasing relational pattern languages under $r_{ab}$ is actually also undecidable, even though $r_{ab}$ generally implies a significantly less restrictive behavior than $r_{=}$.
Proposition 5(✂)
Let $R:=\{r_{ab}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$. It is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
Proof
This can be shown by reducing the problem of deciding whether the universal Turing machine $U$ has any valid computations to the problem of whether the NE-relational pattern language of one relational pattern under abelian equivalence is included in the NE-relational pattern language of another relational pattern under abelian equivalence. It is done by constructing two such relational patterns $(\alpha,\alpha_{R})$ and $(\beta,\beta_{R})$ for which we have that $L_{NE}(\alpha,R)\subseteq L_{NE}(\beta,\beta_{R})$ if and only if $U$ has no valid computation.
As mentioned above, this result follows from the same reduction as in [BREMER201215] for the inclusion problem of NE-pattern languages over the equality relation. Only some technical modifications have to be made to formally get it to work in the setting of relational pattern languages under $r_{ab}$. The proof covering the technical adaptions can be found in Appendix 0.B.4 for reference. ∎
As the construction used to show Proposition 5 only related variables to either enforce equality between length 1 substitutions (letters) or equality between unary strings, it also not only translates to the case of $r_{ab}$, but also to the reversal relation $r_{rev}$. Indeed, the reversal of a unary word $w$ is $w$ itself and the reversal of a letter $\mathtt{a}_{i}$ is clearly also just the letter $\mathtt{a}_{i}$. So, we can enforce the same behavior to select only one $x_{i,1}$ to become the letter #\# in valid substitutions while all other $x_{j,i}$ take the letter 0, as $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are always substituted only to unary words that are powers of the letter 0. With that, the $v$ parts will look exactly the same. Further, the counter values that are represented in simple predicates with the variables $\hat{x}_{i,j}$ will also only take the shape of unary words, as do the variables in the third predicate $\pi_{3}$ in the remaining cases not covered by the predicates $\pi_{1}$ and $\pi_{2}$. So, using an analogous construction as for Proposition 5, we obtain the following.
Corollary 2
Let $R:=\{r_{rev}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$. It is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
4.2 The Erasing Cases
We continue with the additional cases for erasing relational pattern languages. First, it can be shown that the inclusion problem under the reversal relation $r_{rev}$ is also undecidable. We can use the same framework as in the result from Freydenberger and Reidenbach [Freydenberger2010] that shows undecidability in the erasing case under $r_{=}$, for all alphabets $\Sigma$ with $|\Sigma|\geq 2$, as a starting point, but we need to adapt it at several points in a more involved way to get the following theorem. Similarly to Proposition 5, in Appendix 0.B.5, we provide, for reference and completeness, a full reconstruction of the proof from [Freydenberger2010] in the framework of relational pattern languages with adaptations where they are necessary for the proof to work for $r_{rev}$.
Theorem 4.1(✂)
Let $R=\{r_{rev}\}$. For all alphabets $\Sigma$ with $|\Sigma|\geq 2$, given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is generally undecidable whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
Next, as for the non-erasing case, we can do something similar to show undecidability of the inclusion problem also relational pattern languages under abelian equivalence $r_{ab}$.
Theorem 4.2
Let $\Sigma$ be a finite alphabet with $|\Sigma|\geq 2$ and let $R=\{r_{ab}\}$. Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
Proof
This proof is analogous to the one of Theorem˜4.1. Therefore, we only describe how to get to the analogy and refer for a detailed proof to Theorem˜4.1. Let $A$ be a nondeterministic 2-counter automaton. Consider
$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}0$ and
$\beta_{A}:=x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}\#^{6}\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}\#^{6}\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}}$
with, for all $i\in\{1,\ldots,\mu\}$, and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{4}},x_{1_{5}},\ldots$ are distinct variables with $(x_{i},x_{i^{\prime}}),(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ and all $\eta_{i}\in X^{*}$ are terminal-free patterns for some $\mu$ and some $\hat{\beta}_{i}$ as in Theorem˜4.1. We have $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ with distinct variables $z_{i},z_{i^{\prime}},\hat{z}_{i_{1}},\hat{z}_{i_{2}},\hat{z}_{i_{3}},\hat{z}_{i_{4}},\hat{z}_{i_{5}}$ and $(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in R$.
The main concept of the proof of Theorem˜4.1 is to show that iff $L_{E}(\alpha_{A},r_{\alpha_{A}})\subseteq L_{E}(\beta_{a},r_{\beta_{A}})$ then $A$ has no accepting computation. All possible non accepting computations are listed in the $\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}$ part where each $\hat{\beta}_{i},i\in\{1,\ldots,\mu\}$ encodes one non accepting computation together with selecting variables $x_{i_{j}}$ for $j\in{1,2,3}$. The part of $\beta_{A}$ and $\alpha_{A}$ before the first $^{6}$ ensures that exactly one pair $x_{i},x_{i}^{\prime},i\in\{1,\ldots,\mu\}$ is selected, resulting in that exactly one non accepting computation is selected and that exactly one $\ddot{\beta}_{i},i\in\{1,\ldots,\mu\}$ is selected. With this, we get that $\eta_{i},i\in\{1,\ldots,\mu\}$ matches $u$ and thus, that $z_{i}=0$ and $\hat{z}_{i_{1}}=\#$. We use these two variables to encode the 0 and the #\# in the non accepting computation encodings. If we can show that this selecting process also works in the setting $R=r_{ab}$ and that we get two variables with which we can encode the letters 0 and #\#, then this proof also works in this setting. For two variables $x,y$, we have if $x=0$ (or x=#x=\#) and $(x,y)\in r_{ab}$, then $y=0$ (or y=#y=\# respectively). Thus, we can model with the relation $r_{ab}$ identical variables if we know a one letter substitution of one variable.
[…]
$x_{i}$ $x_{i^{\prime}}$ works?
0 0 $00\neq 0\#$
0#0\# 0#0\# or #0\#0 $0#$, $0\#,\ 0\#^{2}0\neq 0\#^{3}$
$0\#^{2}$ $0\#^{2}$ or #0#\#0\# or ##0\#\#0 $0\#^{2}0\#^{2},\ 0\#^{3}0\#,\ 0\#^{4}0\neq 0\#^{3}00$
$0\#^{3}$ $0\#^{3}$ or $0\#^{2}$ or $^{2}0\#$ or $^{3}0$ $0\#^{3}0\#^{3},\ 0\#^{3}\#0\#^{2},\ 0\#^{5}0\#,\ 0\#^{6}30\neq 0\#^{3}00\#^{2}$
$0\#^{3}0$ $0\#^{3}0$ $0\#^{3}00\#^{3}0=v\ v$
Table 1: Possibilities for the first match with vvv\ v
Thus, the only possibility is that for exactly one $i\in\{1,\ldots,\mu\}$ the variables $x_{i}$ and $x_{i^{\prime}}$ get substituted with $v$ and all other variables $x_{j},x_{j^{\prime}}$ for $j\neq i,j\in\{1,\ldots,\mu\}$ get substituted with $\epsilon$. This together with the matching from vuvv\ u\ v to $\ddot{\beta}_{1}\ldots\ddot{\beta}_{\mu}$ implies that $\eta_{i}$ needs to match $u$. For this matching between $\eta_{i}=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ and $u=0\#^{5}0$, we get the following possibilities. Remember that $(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in R$.
$z_{i}$ $\hat{z}_{i_{1}}$ works?
$\epsilon$ 0 or #\# $0^{5},\ \#^{5}\neq 0\#^{5}0$
0 #\# $0\#^{5}0=0\#^{5}0=u$
0#0\# #\# $0\#\#^{5}\#0\neq 0\#^{5}0$
0##0\#\# #\# $0\#^{2}\#^{5}\#^{2}0\neq 0\#^{5}0$
0###0\#\#\# $\epsilon$ $0\#^{6}0\neq 0\#^{5}0$
Table 2: Possibilities for the match with $u$
Thus, the only possibility is that $z_{i}$ and $z_{i^{\prime}}$ get substituted with 0 and $\hat{z}_{i_{k}},k\in\{1,2,3,4,5\}$ get substituted with #\#. With this, we can encode the 0 and the #\# and thus, the proof of Theorem˜4.1 works in this setting analogously. ∎
This concludes the results regarding the inclusion problem for relational pattern languages covered in this paper.
4.3 Discussion and Challenges Regarding the Other Relations
Currently, we know that for relational patterns under the relations $r_{=}$, $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ the inclusion problem is undecidable. This mainly comes from the fact that either the relations can be used to simulate the behavior of $r_{=}$ generally (as in the case of $r_{ssq}$ and $r_{*}$) or it comes from the fact that essential properties used in relevant reductions can also be emulated using other relations (enforce equal letters or equal unary words using $r_{ab}$ or $r_{rev}$).
When it comes to the remaining relations $r_{|w|}$, $r_{perm}$, $r_{com^{*}}$, and $r_{com^{+}}$, we see that these properties are not trivially obtainable. The relations $r_{|w|}$ and $r_{perm}$ both do not force equal letters or equal unary words in alphabets that are at least of binary size. The relations $r_{com^{*}}$ and $r_{com^{+}}$ do not even force equal length between variables in valid substitutions, making, for example, the comparison of counter values impossible in the way it has been done in the previous reductions. Hence, it remains an worthwhile open question whether the inclusion problem becomes decidable at some point, for very structurally weak relations.
Question 1
Let $R=\{\tilde{r}\}$ for some relation $\tilde{r}\in\{r_{|w|},r_{perm},r_{com^{*}},r_{com^{+}}\}$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. Is the inclusion problem $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$, generally decidable?
5 The Equivalence Problem
In this section, we briefly discuss some results regarding the equivalence problem for non-erasing relational pattern languages. As we will see, we will obtain a series of small results that builds up to the main theorem of this section, stating that the equivalence problem for all non-erasing pattern languages over a binary equivalence relation that is antisymmetric on elements of length 1 (i.e., has equality for single letters) is decidable in polynomial time (covering the relations $\{r_{=},r_{ab},r_{con^{+}}\}$ from this paper’s discussion).
We define $r_{\alpha}]$ to be the transitive, reflexive and symmetric closure on some set of binary related variables $r_{\alpha}$. Additionally, for this section, we assume that variables are always introduced inside a pattern with increasing indices starting from $1$ (i.e., $x_{1}$, $x_{2}$, and so on). Due to space constraints, the technical proofs of almost all steps can be found in Appendix 0.C, while the steps themselves are supposed to give some intuition on why the main result of this section, Theorem 5.1, holds. We begin with a result regarding $r_{\alpha}]$ in the case of $r_{\alpha}$ being an equivalence relation.
Lemma 3
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation and let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. Then we have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$.
Proof
($\Rightarrow$): Let $w\in L_{NE}(\alpha,r_{\alpha})$. Then there exists $h\in H$ such that $h((\alpha,r_{\alpha}))=w$. Let $(\tilde{r},x,y)\in r_{\alpha}$ for $x,y\in\operatorname{var}(\alpha)$. Then $(h(x),h(y))\in\tilde{r}$ by definition. As $\tilde{r}$ is an equivalence relation, we also have $(h(y),h(x))\in\tilde{r}$, $(h(x),h(x))\in\tilde{r}$, and if there exists $(\tilde{r},y,z)\in r_{\alpha}$ such that $(h(y),h(z))\in\tilde{r}$ then also $(h(x),h(z))\in\tilde{r}$. So, $h$ satisfied all constraints set by $r_{\alpha}]$. Hence, $w\in L_{NE}(\alpha,[r_{\alpha}])$.
($\Leftarrow$): Let $w\in L_{NE}(\alpha,[r_{\alpha}])$ and let $h\in H$ such that $h((\alpha,[r_{\alpha}]))=w$. As $r_{\alpha}\subseteq[r_{\alpha}]$, all constraints in $r_{\alpha}$ must be satisfied. Thus, we also have $w\in L_{NE}(\alpha,r_{\alpha})$. ∎
Following on that, we can also obtain the following relationship.
Lemma 4(✂)
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being some equivalence relation and assume $|\Sigma|\geq 2$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. If $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$, then $\alpha=\beta$.
Using the previous lemma, we can show the following result.
Lemma 5(✂)
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation that is antisymmetric on words of length 1 (i.e., equality for letters) and assume $|\Sigma|\geq 2$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two relational patterns. Then $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$ if and only if $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
Finally, taking all previous lemmas into account, we obtain the main result of this section.
Theorem 5.1(✂)
Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation that is antisymmetric on words of length 1 (i.e., equality for letters) and assume $|\Sigma|\geq 2$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two relational patterns. Then $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ if and only if $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
Hence, for the considered relations in this paper, we obtain the following.
Corollary 3(✂)
Let $R=\{\tilde{r}\}$ for some $\tilde{r}\in\{r_{=},r_{ab},r_{con^{+}}\}$. Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, we can check in polynomial time whether $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
6 Conclusion
We considered the three main decision problems membership, inclusion, and equivalence for relational pattern languages under various typical binary relations, i.e., $r_{=}$, $r_{ssq}$, $r_{ab}$, $_{perm}, rr$$_{rev}$, $r_{com^{+}}$, $r_{com^{*}}$, and $r_{*}$.
In particular, for the membership problem, we observed that in all considered cases, that is, for all relations in $\mathcal{R}$ and for erasing as well as non-erasing relational pattern languages, the problem turns out to be NP-complete, even for very simple relations such as $r_{|w|}$. As the reduction relies on rather simple properties to hold between related areas inside the constructed patterns, we expect the problem to be NP-hard for generally almost all typical relations that could also be considered. This allows for the assumption that searching for words in sets of texts that are compactly described by, e.g., a relational pattern of almost any kind, is generally a computationally hard problem.
For the inclusion problem, we observed that many cases remain undecidable. In the erasing as well as the non-erasing cases, the problem remains undecidable for the structurally less restrictive relations $r_{ssq}$, $r_{ab}$, and $r_{*}$ as well as for the relation $r_{rev}$. In the cases of $r_{ssq}$ and $r_{*}$, we can abuse the property that relating two variables in both directions results basically in $r_{=}$. For the other cases, we observe that the key for the reductions to work is, first, finding ways to force certain variables to be substituted by equal single letters to realize structural properties and, second, to force equality of longer unary words to realize the comparison of counter values. It would be interesting to know whether similar results can be obtained using the other relations that enforce even less structure in valid substitutions.
Regarding the equivalence problem, we obtained decidability in polynomial time for $r_{ab}$ and $r_{com^{+}}$ in the non-erasing case, similar to $r_{=}$. We conjecture that this should hold for many more relations in the non-erasing case by checking some set of minimal length substitutions. For the erasing cases, everything except $r_{|w|}$ remains open and we expect the decidability question to be generally rather difficult to answer, as it is the main open question for $r_{=}$.
This concludes this paper. See Table 3 for a visual summary of the results.
$r_{=}$ $r_{|w|}$ $r_{ssq}$ $r_{ab}$ $_{perm} rr$$_{rev}$ $r_{com^{*}}$ $_{com^{+}} rr_{*}$
 E ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC
 E ($\subseteq$)  UD  P UD UD ? UD ? ? UD
 E (==) ?  P ? ? ? ? ? ? ?
 NE ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC
 NE ($\subseteq$)  UD  P UD UD ? UD ? ? UD
 NE (==)  P  P ? P ?  D ? P ?
Table 3: Current state of decision problems for relational pattern languages for various constraints (NPC = NP-Complete, UD = Undecidable, D = Decidable, E = Erasing, NE = Non-Erasing, ($\in$) membership problem, ($\subseteq$) inclusion problem, (==) equivalence problem). Red text marks the results covered in this paper.
Appendix 0.A Additional Content for Membership Problem
Proof of Proposition˜1
Proof
Assume $|\Sigma|\geq 2$. Let $r\in\mathcal{R}$, $R=\{r\}$, $w\in\Sigma^{*}$, and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. We can guess a substitution $h\in H$ which results in $|h(\alpha)|=|w|$, then check whether $(x,y)\in r$ for all $(r,x,y)\in r_{\alpha}$, and finally, if that is the case, return true if $h(\alpha)=w$ and return false in any other case. As all relations $r\in\mathcal{R}$ are polynomial-time decidable relations, this results in a polynomial runtime of the verifier, hence resulting for the problem to be in NP. ∎
Proof of Lemma˜1:
Proof
By the result from Angluin in [Angluin1980], we know that this statement already holds for the case of $r_{=}$. This is shown using a reduction from the $3\mathtt{-SAT}$ problem. We quickly recall the construction used in [Angluin1980] and refer to the original source for the proof of its correctness.
Notice that the pattern constructed here is not a relational pattern by our definition but a pattern using the usual notation allowing for multiple occurrences of the same variable. We will comment on the adaptions needed for the notation of relational pattern languages after giving the construction.
Let $\varphi$ be a $3\mathtt{-SAT}$ formula. Let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the set of the used variables in $\varphi$ for some $m\in\mathbb{N}$. Let $C=\{C_{1},...,C_{n}\}$ be the set of all clauses in $\varphi$ that are in conjunction to each other for some $n\in\mathbb{N}$. For each clause $C_{i}\in C$, refer to its three variables by $X_{i,1}$, $X_{i,2}$, and $X_{i,3}$. Define a function $f:\mathcal{X}\rightarrow V$ by $f(X_{i})=v_{i}$ and $f(\overline{X_{i}})=u_{i}$ for new independent pattern variables $u_{i},v_{i}\in V$ with $i\in[m]$. Now, construct the word
$w=\#\ s_{1}\ \#\ ...\ \#\ s_{m}\ \#\ t_{1}\ \#\ ...\ \#\ t_{n}\ \#\ w_{1}\ \#\ ...\ \#\ w_{n}\ \#$
with $s_{i}=1^{3}$, $t_{j}=1^{7}$, and $w_{j}=1^{4}$ for $i\in[m]$ and $j\in[n]$ as well as a pattern
$\alpha=\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\ r_{1}\ \#\ ...\ \#\ r_{n}\ \#$
with $p_{i}=u_{i}v_{i}$, $q_{j}=f(X_{j,1})f(X_{j,2})f(x_{j,3})z_{j}$, and $r_{j}=z_{j}d_{j}$ for new independent pattern variables $z_{j},d_{j}\in X$ with $j\in[n]$. Notice that the construction uses three different elements. The first element contains blocks $u_{i}v_{i}\#$ for $i\in[m]$ which have to be non-emptily substituted to $1^{3}\#$, resulting in one of the variables to be substituted to $11$ and the other one to be substituted to $1$, representing either $X_{i}$ or $\overline{X_{i}}$ to be mapped to true in $\varphi$, having the other one false. The middle part $f(X_{i,1})f(X_{i,2})f(X_{i,3})z_{i}\#$ needs to be non-emptily mapped to $1^{7}$. Notice that due to the last part $z_{i}d_{i}\#$, which is to be substituted to $1^{4}\#$, the variable $z_{i}$ may only be substituted tp $1$, 1111 or $111$, as $d_{i}$ has do be substituted at least to a word of length $1$, i.e., the word $1$. Because of that, at least one of $f(X_{i,j})$ needs to be mapped to $11$ instead of $1$ to obtain $1^{7}\#$, representing at least one variable in each clause to be mapped to true. As this holds for all clauses, deciding whether $w\in L_{NE}(\alpha)$ also decides the satisfiability of the formula $\varphi$. Consider [Angluin1980] for the formal proof.
To obtain this construction in the notation of relational pattern languages, instead of introducing each variable $k\in\mathbb{N}$ times, introduce $k$ different variables in a pattern $\alpha^{\prime}$ that has exactly the same terminal letters and relate those $k$ variables with equality $r_{=}$ in the relation set $r_{\alpha^{\prime}}$. This results in a relational pattern $(\alpha^{\prime},r_{\alpha})$ with $L_{NE}(\alpha)=L_{NE}(\alpha^{\prime},r_{\alpha^{\prime}})$ for which deciding the membership $w\in L_{NE}(\alpha,r_{\alpha})$ also decides the satisfiablity of $\varphi$.
Now we give arguments why the exact same construction works for the other relations in $\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$.
  1. 1.
    If $R=\{r_{|w|}\}$, then variables have to be substituted with equal length. As any substitution resulting in $w$ only uses unary words for the variables, having equal length also means being the equal word. This concludes this case.
  2. 2.
    If $R=\{r_{ssq}\}$, then relating two variables $x$ and $y$ in both directions, i.e., putting $(r_{ssq},x,y)\in r_{\alpha^{\prime}}$ as well as $(r_{ssq},y,x)\in r_{\alpha^{\prime}}$ results in some substitution $h\in H$ to be valid if and only if $h(x)=h(y)$, as otherwise one can’t be a subsequence of the other. That concludes this case.
  3. 3.
    If $R=\{r_{ab}\}$, then valid substitutions of related variables result in them to have equal length, as abelian equivalence implies that. Hence, this case follows as in the case of the equal length relation $r_{|w|}$.
  4. 4.
    If $R=\{r_{perm}\}$, then we also get that two related variables must be substituted with equal length. As all substitutions must result in unary words consisting only of the letter $1$, this case also follows directly as the case of the equal length relation.
  5. 5.
    Also in the case of $R=\{r_{rev}\}$ we can use the argument that only unary substitutions are relevant in this reduction. As the reversal of a unary word is again the same unary word, this case also follows immediately.
  6. 6.
    Finally, in the case of $R=\{r_{*}\}$, as we said before, we may use this relation in both directions for each pair of variables, resulting effectively in equality of variables as in the case of the subsequence relation $r_{ssq}$.
This concludes all cases. For larger alphabets, the same construction may be used by trivially extending the word and the pattern by terminal letters made of up all additional letters in the alphabet. ∎
Proof of Proposition˜2:
Proof
We show this, also, by reducing from the $3\mathtt{-SAT}$ problem. So, as before, let $\varphi$ be a $3\mathtt{-SAT}$ formula and let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the set of the used variables in $\varphi$ for some $m\in\mathbb{N}$. Let $C=\{C_{1},...,C_{n}\}$ be the set of all clauses in $\varphi$ that are in conjunction to each other for some $n\in\mathbb{N}$. For each clause $C_{j}\in C$ for $j\in[n]$, refer to its three variables by $X_{j,1}$, $X_{j,2}$, and $X_{j,3}$. W.l.o.g. assume that no variable occurs two times in the same clause (a variable and its reversal, however, may still occur together in one clause). Define $n+1$ functions $f_{j^{\prime}}$ with $j_{j^{\prime}}\in[n]_{0}$ such that $f_{j^{\prime}}:\mathcal{X}\rightarrow X$ with $f_{j^{\prime}}(X_{i})=u_{i,j^{\prime}}$ and $f_{j^{\prime}}(\overline{X_{i}})=v_{i,j^{\prime}}$. Assume $R=\{r\}$ for $r\in\{r_{com^{+}},r_{com^{*}}\}$. Construct the word
$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
with
$s_{i}$ $=1\#1\text{, and}$
$t_{j}$ $=1^{10}\#1^{10}\#1^{10}$
for $i\in[m]$ and $j\in[n]$ as well as the relational pattern $(\alpha,r_{\alpha})$ such that $\alpha$
$\alpha=\#\#\ \alpha_{1,1}\ \#\#\ ...\ \#\#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\#\ ...\ \#\#\ \alpha_{2,n}\ \#\#$
for $i\in[m]$ and $j\in[n]$ with
$\alpha_{1,i}$ $=f_{0}(X_{i})f_{0}(\overline{X_{i}}),\text{ and}$
$\alpha_{2,j}$ $=z_{j,1}\ f_{j}(X_{j,1})\ z_{j,2}\ f_{j}(X_{j,2})\ z_{j,3}\ f_{j}(X_{j,3})\ z_{j,4},$
and the set of related variables $r_{\alpha}=\{\ (r,u_{i,j},u_{i,j^{\prime}}),(r,v_{i,j},v_{i,j^{\prime}})\ |\ i\in[m],j,j^{\prime}\in[n+1]\ \}$. Hence, notice that two variables with the same base name $u_{i,j}$ and $u_{i^{\prime},j^{\prime}}$ (resp. $v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$) are related to each other if and only if $i=i^{\prime}$. Also notice, that each variable $z_{i,j}$ is independent and may be substituted to anything.
Now, assume there exists an assignment of the variables in $\mathcal{X}$ such that $\varphi$ is satisfied. We will now construct a $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$. For each $i\in[m]$, if $X_{i}$ is true (hence, $\overline{X_{i}}$ is false), set $h(u_{i,j})=1$ and $h(v_{i,j})=\#1$, otherwise set $h(u_{i,j})=1\#$ and $h(v_{i,j}=1)$ for $j\in[n+1]$. This results in $h(\alpha_{1,i})=1\#1$ for all $i\in[m]$. As there is at least one variable assigned to true for each clause $j\in[n]$, we know that $|h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))|_{\#}\leq 2$. As there is at most two $1$’s between two occurrences of #\# and at most $3$ occurrences of $1$’s before or after the first or last occurrence of #\# in $h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))$, we always find a non-empty substitution of all variables $z_{j,1}$ to $z_{j,4}$ such that $h(\alpha_{2,j})=\#\#\ 1^{10}\#1^{10}\#1^{10}$. All substitutions done by $h$ result in related variables to be substituted equally, hence, this satisfies the relation $r$, and by that there exists some $r_{\alpha}$ valid $h\in H$ such that $h(\alpha)=w$.
Now assume $w\in L_{NE}(\alpha,r_{\alpha})$. Then there exists some $r_{\alpha}$ valid substitution $h\in H$ such that $h(\alpha)=w$. We know that each occurrence of the factor ##\#\# in $w$ must be mapped by one occurrence of the factor ##\#\# in $h(\alpha)$. As the same number of ##\#\# occur as terminal letters in $\alpha$ as there are occurrences of it in $w$, we know that these occurrences must align. Hence, we get that for all $i\in[m]$ and $j\in[n]$ we have
$h(\alpha_{1,i})=h(u_{i,1}v_{i,1})=1\#1=s_{i}$
and
$h(\alpha_{2,j})=h(z_{j,1}\ f_{j}(X_{j,1})\ z_{j,2}\ f_{j}(X_{j,2})\ z_{i,3}\ f_{j}(X_{j,3})\ z_{j,4})=1^{10}\#1^{10}\#1^{10}=t_{j}.$
By the first equality, we obtain that either $h(u_{i,1})=1$ and $h(v_{i,1})=\#1$ or $h(u_{i,1})=1\#$ and $h(v_{i,1})=1$. Notice by the definition of $r$, in this case, that for all $i\in[m]$ and $j\in[n]$ we have that $h(u_{i,j})\in\{1\}^{*}$ if and only if $h(u_{i,0})=1$. The same holds analogously for all $v_{i,j}$. Now select an assignment for the variables in $\mathcal{X}$ for the $3\mathtt{-SAT}$ formula by setting $X_{i}=true$ (hence $\overline{X_{i}}=false$) iff $h(u_{i,1})=1$ and otherwise $\overline{X_{i}}=true$ (hence $X_{i}=false$) iff $h(v_{i,1})=1$. By the argument from before, this is a valid assignment of the variables. If there were only falsely assigned variables in a single clause, using the previous observations, there existed $j\in[n]$ such that $|h(\alpha_{2,j})|_{\#}\geq 3$, which is not the case. Hence, there is at least one variable assigned to true per clause and by that we obtain that $\varphi$ must be satisfiable. ∎
Proof of Lemma˜2:
Proof
Again, the original proof in [Jiang1994] for the case of equality between variables uses a reduction from $3\mathtt{-SAT}$ to the membership problem. The construction is essentially the same as in [Angluin1980] just with different sizes of each part of the word $w$ to construct. Here, we have for all $i\in[m]$ and $j\in[n]$ that $s_{i}=1$, $t_{j}=1^{7}$, and $w_{j}=1^{2}$. By that, as variables are allowed to be substituted by the empty word, the reduction works in exactly the same way. Consider [Angluin1980, Jiang1994] for the formal proof of this construction. As the reasoning about the construction works exactly in the same way as before, and as the considered relations don’t change in their general behavior if the empty word is allowed, we can use the same arguments as in Lemma 1 and obtain NP-hardness for all these cases here as well. ∎
Proof of Proposition˜3:
Proof
Consider the construction from Proposition 2. As mentioned before, because of the definition of $r_{com^{+}}$, two related variables may never be substituted to the empty word, as $\varepsilon\notin\{y\}^{+}$ for all $y\in\Sigma^{*}$. As only the variables with the base name $z$ are independent, only those may be substituted emptily. Those are only used to fill up remaining #\#’s or $1$’s in the encoded clauses in the case of valid assignments for the $3\mathtt{-SAT}$ formula $\varphi$. Thus, the main argument stays the same and can be applied here in exactly the same way. ∎
Proof of Proposition˜4:
Proof
Again, we reduce from $3\mathtt{-SAT}$. Let $\varphi$ be a $3\mathtt{-SAT}$ formula, let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the variables in $\varphi$ for some $m\in\mathbb{N}$, let $C=\{C_{1},...,C_{n}\}$ be the set of clauses in $\varphi$ for some $n\in\mathbb{N}$. For each clause $C_{j}\in C$, refer to its three variables by $X_{j,1}$, $X_{j,2}$, and $X_{j,3}$ for $j\in[n]$. W.l.o.g. assume that no variable occurs two times in the same clause (a variable and its reversal, however, may still occur together in one clause). Define $n+1$ functions $f_{j}$ with $j\in[n]_{0}$ such that $f_{j}:\mathcal{X}\rightarrow X$ with $f_{j}(X_{i})=u_{i,j}$ and $f_{j}(\overline{X_{i}})=v_{i,j}$ for $i\in[m]$. Now, let
$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
with
$s_{i}$ $=1\#1,\text{ and}$
$t_{j}$ $=1$
for $i\in[m]$ and $j\in[n]$ and construct a relational pattern $(\alpha,r_{\alpha})$ by
$\alpha=\#\#\alpha_{1,1}\ \#\#\ ...\ \#\#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\#\ ...\ \#\#\ \alpha_{2,n}\ \#\#$
with
$\alpha_{1,i}$ $=f_{0}(X_{i})f_{0}(\overline{X_{i}})\text{, and}$
$\alpha_{2,j}$ $=f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3})$
and the set of related variables $r_{\alpha}=\{\ (r,u_{i,j},u_{i,j^{\prime}}),(r,v_{i,j},v_{i,j^{\prime}})\ |\ i\in[m],j,j^{\prime}\in[n+1]\ \}$. Hence, notice that two variables with the same base name $u_{i,j}$ and $u_{i^{\prime},j^{\prime}}$ (resp. $v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$) are related to each other if and only if $i=i^{\prime}$.
First, assume there exists a satisfying assignment for $\varphi$. Let $h\in H$ such that for all $i\in[m]$ and $j\in[n]_{0}$ we have $h(u_{i,0})=1$ and $h(v_{i,0})=\#1$ iff $X_{i}$ is true (hence $\overline{X_{i}}$ is false) and otherwise $h(u_{i,0})=1\#$ and $h(v_{i,0})=1$ iff $X_{i}$ is false (hence $\overline{X_{i}}$ is true). As the assignment for $\varphi$ is true, we have for each clause that there is at least one variable $u_{i,j}$ (or $v_{i,j}$) such that $(r_{com^{*}},u_{i,j},u_{i,0})\in r_{\alpha}$ (or $(r_{com^{*}},v_{i,j},v_{i,0})$) and $h(u_{i,0})=1$ (or $h(v_{i,0})=1$). Hence we may set $h(u_{i,j})=1$ (or $v_{i,j}=1$) and for all other variables $z$ in that clause set $h(z)=\varepsilon$. As the empty word commutes with any other word by the definition of $r_{com^{*}}$, this substitution is $r_{\alpha}$-valid. As we can do that for each $j\in[n]$, we get can find a $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$, concluding this direction.
For the other direction assume $w\in L_{E}(\alpha,r_{\alpha})$. Then there exists some $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$. For each $j\in[n]$, we know that $h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))=1$. Hence, w.l.o.g. there exists one variable $u_{i,j}$ for $i\in[m]$, such that $h(u_{i,j})=1$ and $(r_{com^{*}},u_{i,1},u_{i,j})\in r_{\alpha}$. By the definition of $r_{com^{*}}$ and $w$ we know that $h(u_{i,0})\in\{\varepsilon,1\}$ for $h$ to be $r_{\alpha}$-valid. In that case we know that $h(v_{i,0})\in\{1\#1,\#1\}$ and by that we have $h(v_{i,j})\neq 1$ for any $j\in[n]$ by the definition of $r_{\alpha}$. So, we can set an assignment of $X$ for $\varphi$ by setting $X_{i}=true$ iff $h(u_{i,1})\in\{\varepsilon,1\}$ and setting $\overline{X_{1}}=true$ iff $h(v_{i,1})\in\{\varepsilon,1\}$. As for each clause $j$ we find either a variable $u_{i,j}$ or $v_{i,j}$ substituted to $1$, we know that the assignment for $X$ is satisfying for $\varphi$. That concludes the reduction. ∎
0.A.1 The Membership Problem under the One-Sided Star and One-Sided Subsequence Relations
In this subsection, we briefly discuss the special cases mentioned in the end of [Angluin1980], which considers one-sided relations of potential interest. These cases cannot be expressed directly using the framework of this paper.
Consider, for example, the pattern $\alpha=\mathtt{a}x\mathtt{b}x^{*}\mathtt{c}$ that assumes that $h(x^{*})\in\{h(x)\}^{*}$ needs to hold for some valid substitution $h\in H$. This can be mapped using the relation $r_{*}$, but the way Angluin described that case, we should not be able to have the $r_{*}$ hold in both directions between $x$ and $x^{*}$, which, however, is allowed in relational pattern languages over $r_{*}$. Hence, these special cases of one-sided relations are considered for the subsequence and the star relation at the end of this section.
For simplicity, only in this subsection, we denote patterns similar to the notation used in [Angluin1980], i.e., for the one-sided star we write, for example, $\alpha=\#x\#x^{*}\#$ instead of introducing some variables $x_{1}$ and $x_{2}$ separately and relating them externally by some relation $r$. In particular, in this subsection, given $x$ and $x^{*}$, a substitution $h$ is valid if and only if $h(x^{*})\in\{x\}^{*}$, and, given $x$ and $x^{s}$, a substitution $h$ is valid if and only if $h(x^{s})\in\operatorname{SubSeq}(x)$. For the erasing case, we obtain that the same reduction can be used for both cases, one-sided star and one-sided subsequence.
Proposition 6
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{*}$ (resp. $x^{s}$) may occur multiple times, it is NP-complete to decide whether $w\in L_{E}(\alpha)$.
Proof
NP-containment follows from the fact that we can guess a valid substitution $h$ that results in $|h(\alpha)|=|w|$ and check whether $h(\alpha)=w$ holds, all in polynomial time. To establish NP-hardness, we reduce from $3\mathtt{-SAT}$ as before and use the same formal framework as used in the reductions before. Assuming the number of variables in the formula is $m\in\mathbb{N}$ and assuming the number of clauses of the formula is $n\in\mathbb{N}$, the word $w\in\Sigma^{*}$ will be set to
$w=\#\#\underbrace{\ 1\ \#\ ...\ \#\ 1\ }_{m\text{ blocks}}\#\#\underbrace{\ 1\ \#\ ...\ \#\ 1\ }_{n\text{ blocks}}\#\#.$
The corresponding pattern $\alpha$ is given by
$\alpha=\#\#\ \alpha_{1,1}\ \#\ ...\ \#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\ ...\ \#\ \alpha_{2,n}\ \#\#$
where $\alpha_{1,i}=u_{i}v_{i}$, for $i\in[m]$ and $u_{i},v_{i}$ being new pattern variables, and $\alpha_{2,i}=f(X_{i,1})f(X_{i,2})f(X_{i,3})$, where $f(X_{i,j})=u_{k}^{*}$ if $X_{i,j}=X_{k}$ (resp. $f(X_{i,j})=v_{k}^{*}$ if $X_{i,j}=\overline{X_{k}}$) in the one-sided star framework, and $f(X_{i,j})=u_{k}^{s}$ if $X_{i,j}=X_{k}$ (resp. $f(X_{i,j})=v_{k}^{s}$ if $X_{i,j}=\overline{X_{k}}$) in the one-sided subsequence framework.
The correctness of the reduction comes from the fact that, for $w\in L_{E}(\alpha)$ to hold, exactly one variable in each $\alpha_{1,i}$, for $i\in[m]$, must be set to $1$, either $h(u_{i})=1$ or $h(v_{i})=1$. A variable $u_{i}^{*}$ or $v_{i}^{*}$ (resp. $u_{i}^{s}$ or $v_{i}^{s}$), for $i\in[m]$, that occurs in some $\alpha_{2,j}$, for $j\in[m]$, can only be substituted by $1$ if the corresponding variable in $u_{i}$ or $v_{i}$ in $\alpha_{1,i}$ is substituted by $1$. Bringing these two facts together immediately results in that the $3\mathtt{-SAT}$ formula must have an accepting configuration if and only if $w\in L_{E}(\alpha)$, as we can only substitute all $\alpha_{2,i}$, for $i\in[n]$, by $1$ if and only if there exists some configuration of the variables where either $X_{k}$ or $\overline{X_{k}}$ is set to true and each clause contains at least one of these variables that are set to true. ∎
For the non-erasing cases, we can again use similar reductions as for all the other problems before, but the configuration is slightly different for one-sided star and one-sided subsequence, respectively.
Proposition 7
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{*}$ may occur multiple times, it is NP-complete to decide whether $w\in L_{NE}(\alpha)$.
Proof
NP-containment follows from the fact that we can guess a valid substitution $h$ that results in $|h(\alpha)|=|w|$ and check whether $h(\alpha)=w$ holds, all in polynomial time. To establish NP-hardness, we reduce from $3\mathtt{-SAT}$ as before, again, assuming that a given formula $\phi$ has $n$ clauses and uses $m$ distinct variables. We set $w\in\Sigma^{*}$ to
$w=\#\#\underbrace{\ 1^{3}\ \#\ ...\ \#\ 1^{3}\ }_{m\text{ blocks}}\#\#\underbrace{\ 1^{6}\ \#\ ...\ \#\ 1^{7}\ }_{n\text{ blocks}}\#\#$
The pattern $\alpha$ is basically a prefix of the construction used in Lemma 1, i.e.,
$\alpha=\#\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\#,$
where $p_{1}$ to $p_{m}$ and $q_{1}$ to $q_{m}$ are constructed as in the proof of Lemma 1, just with the only difference that every occurrence of $u_{i}$ or $v_{i}$, for $i\in[m]$, in some $q_{1}$ to $q_{n}$, is replaced by $u_{i}^{*}$ and $v_{i}^{*}$ respectively. Then, each $u_{i}$ and $v_{i}$ without a start occurs only once, that is in $p_{i}$. All other occurrences are with a star. The variables $z_{i}$, for $i\in[n]$, are unchanged in $q_{i}$ to $q_{n}$.
To show correctness of this reduction, notice the fact that each $p_{i}$ is still to be substituted to $1^{3}$, resulting in one of $u_{i}$ or $v_{i}$ being substituted by $1$ while the other is substituted by $11$. Assume that a substitution to $1$ represents a true assignment while a substitution by $11$ represents a false assignment. For each clause $i\in[n]$, we know that $q_{i}=x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*}z_{i}$, assuming each variable $x_{i,j}$ is either some $u_{k}$ or $v_{k}$, for $k\in[m]$. We know that $z_{i}$ must be substituted by a word of at least length $1$. Hence, for $h(q_{i})=1^{6}$ to be the case, we then need $|h(x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*})|\leq 5$. If $h(x_{i,j})=11$, for all $j\in[3]$, then we know that $|h(x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*})|\geq 6$, as $h(x_{i,j})=\varepsilon$ is not valid in the non-erasing case. So, for at least one variable $x_{i,j}$ we need $h(x_{i,j})=1$ for $h(q_{i})=1^{6}$ to be possible. As this holds for all clauses, we can only find a substitution $h$ for which $h(\alpha)=w$ if and only if the given $3\mathtt{-SAT}$ is satisfiable. ∎
As we only do slight changed to the constructed word and pattern, we give the proof of the following result in a very short form and assume the same formal framework as in Proposition 7.
Proposition 8
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{s}$ may occur multiple times, it is NP-complete to decide whether $w\in L_{NE}(\alpha)$.
Proof
NP-containment follows as before, e.g. as in Proposition 7. For NP-hardness, we reduce from $3\mathtt{-SAT}$ as before. The word $w\in\Sigma^{*}$ is given by
$w=\#\#\underbrace{\ 1^{3}\ \#\ ...\ \#\ 1^{3}\ }_{m\text{ blocks}}\#\#\underbrace{\ 1^{4}\ \#\ ...\ \#\ 1^{4}\ }_{n\text{ blocks}}\#\#.$
The pattern $\alpha$ is given by
$\alpha=\#\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\#,$
where $p_{i}=u_{i}v_{i}$, for $i\in[m]$ and new variables $u_{i}$ and $v_{i}$, and we have $q_{i}=f(X_{i,1})f(X_{i,2})f(X_{i,3})$ for $f$ being defined as before as well, just using $u_{j}^{s}$ and $v_{j}^{s}$ instead of $u_{j}$ and $v_{j}$, for $j\in[m]$ and $i\in[n]$. Now notice, again, that for each $p_{i}$, for $i\in[m]$, one of $u_{i}$ and $v_{i}$ must be substituted by $1$ while the other one is substituted by $11$. Here, we consider a substitution by $11$ as a true assignment of that variable configuration. For all other variable substitutions we know that they must be a subsequence of either some $u_{i}$ or $v_{i}$. As we are in the non-erasing case, for a substitution $h$ to be valid, that results in w.l.o.g. $h(u_{i}^{s})=1$ if $h(u_{i})=1$ and $h(u_{i}^{s})\in\{1,11\}$ if $h(u_{i})=11$, for $i\in[m]$ (respectively for $v_{i}$). So, we can only obtain $h(q_{i})=1^{4}$, for all $i\in[n]$, if and only if there is at least one variable $u_{i}^{s}$ (or $v_{i}^{s}$) in $q_{i}$, where $h(u_{i})=11$ (or $h(v_{i})=11$). By the same arguments as in all results before, we can only find a substitution $h\in H$ for which $h(\alpha)=w$ if and only if there exists a satisfying assignment of the variable for the given $3\mathtt{-SAT}$ formula, concluding this reduction. ∎
This concludes all results necessary for Theorem 3.2 to hold.
Appendix 0.B Additional Content for Inclusion Problem
0.B.1 Definition of Nondeterminstic 2-Counter Automata without Input
A nondeterministic 2-counter automaton without input (see e.g. [Iberra1978]) is a 4-tuple $A=(Q,\delta,q_{0},F)$ which consists of a set of states $Q$, a transition function $\delta:Q\times\{0,1\}^{2}\rightarrow\mathcal{P}(Q\times\{\-1,0,+1\}^{2})$, an initial state $q_{0}\in Q$, and a set of accepting states $F\subseteq Q$. A configuration of $A$ is defined as a triple $(q,m_{1},m_{2})\in Q\times\mathbb{N}\times\mathbb{N}$ in which $q$ indicates the current state and $m_{1}$ and $m_{2}$ indicate the contents of the first and second counter. We define the relation $\vdash_{A}$ on $Q\times\mathbb{N}\times\mathbb{N}$ by $\delta$ as follows. For two configurations $(p,m_{1},m_{2})$ and $(q,n_{1},n_{2})$ we say that $(p,m_{1},m_{2})\vdash_{A}(q,n_{1},n_{2})$ if and only if there exist $c_{1},c_{2}\in\{0,1\}$ and $r_{1},r_{2}\in\{-1,0,+1\}$ such that
  1. 1.
    if $m_{i}=0$ then $c_{i}=0$, otherwise if $m_{i}>0$, then $c_{i}=1$, for $i\in\{1,2\}$,
  2. 2.
    $n_{i}=m_{i}+r_{i}$ for $i\in\{1,2\}$,
  3. 3.
    $(q,r_{1},r_{2})\in\delta(p,c_{1},c_{2})$, and
  4. 4.
    we assume if $c_{i}=0$ then $r_{i}\neq-1$ for $i\in\{1,2\}$.
Essentially, the machine checks in every state whether the counters equal 0 and then changes the value of each counter by at most one per transition before entering a new state. A computation is a sequence of configurations. An accepting computation of $A$ is a sequence $C_{1},...,C_{n}\in(Q\times\mathbb{N}\times\mathbb{N})^{n}$ with $C_{1}=(q_{0},0,0)$, $C_{i}\vdash_{A}C_{i+1}$ for all $i\in\{1,...,n-1\}$, and $C_{n}\in F\times\mathbb{N}\times\mathbb{N}$ for some $n\in\mathbb{N}$ with $n>0$.
We encode configurations of $A$ by assuming $Q=\{q_{0},...,q_{e}\}$ for some $e\in\mathbb{N}$ and defining a function $\operatorname{enc}$ :: $Q\times\mathbb{N}\times\mathbb{N}\rightarrow\{0,\#\}^{*}$ by
$\operatorname{enc}(q_{i},m_{1},m_{2}):=0^{x+i}\#0^{c_{1}+y_{2}\cdot m_{1}}\#0^{c_{2}+y_{2}\cdot m_{2}}$
for some numbers $x,c_{1},y_{2},c_{1},y_{2}\in\mathbb{N}$. The values for these numbers depend on the construction of the respective proofs and are not specified here. This is extended to encodings of computations by defining for every $n\geq 1$ and every sequence $C_{1},...,C_{n}\in Q\times\mathbb{N}\times\mathbb{N}$
$\operatorname{enc}(C_{1},...,C_{n}):=\#\#\ \operatorname{enc}(C_{1})\ \#\#\ ...\ \#\#\ \operatorname{enc}(C_{n})\ \#\#.$
Encodings of this kind are used to prove different undecidability results regarding the inclusion problem for erasing relational pattern languages for various considered relations, in particular the results in Subsection 4.2. For some nondeterministic 2-counter automaton without input $A$, define the set of encodings of accepting computations
$\mathtt{ValC}(A):=\{\operatorname{enc}(C_{1},...,C_{n})\ |\ C_{1},...,C_{n}\text{ is an accepting computation of }A\}$
and let $\mathtt{InvalC}(A)=\{0,\#\}^{*}\setminus\mathtt{ValC}(A)$. The emptiness problem for deterministic 2-counter-automata is undecidable (cf. e.g. [Iberra1978, Minsky1961]), thus it is also undecidable whether a nondeterministic 2-counter automaton without input has an accepting computation [Freydenberger2010, Jiang1995]. That the emptiness problem for universal Turing machines is undecidable is a known fact.
0.B.2 Definition of the Universal Turing Machine $U$
Here, we define the universal Turing machine $U$ used in the referenced proof from [BREMER201215] and referred to in the proofs for the undecidability of the inclusion problem for NE relational pattern languages under $r_{ab}$ and rev in Proposition 5 and Corollary 2. Let $U=(Q,\Gamma,\delta)$ be the universal Turing machine $U_{15,2}$ wih 2 symbols and 15 states as described by Neary and Woods [Neary2009]. This machine hat the states $Q=\{q_{1},...,q_{15}\}$ and the tape alphabet $\Gamma=\{0,1\}$. The transition function $\delta:\Gamma\times Q\rightarrow(\Gamma\times\{L,R\}\times Q)\cup\mathtt{HALT}$ is given in Table 4.
We follow with the definition of encodings of computations of $U$ as depicted in [BREMER201215]. The following conventions are needed to discuss configurations of $U$. The tape content of any configuration of $U$ is characterized by two infinite sequences $t_{L}=(t_{L,n})_{n\geq 0}$ and $t_{R}=(r_{R,n})_{n\geq 0}$ over $\Gamma$. The sequence $t_{L}$ describes the left side of the tape, the sequence starting at the head position of $U$ (including) and extending to the left. Analogously, $t_{R}$ describes the right side of the tape, the sequence starting directly after the head position and extending to the right. A configuration $C=(q_{i},t_{L},t_{R})$ of $U$ is a triple consisting of a state $q_{i}$, a left side of the tape $t_{L}$ and a right side of the tape $t_{R}$.
Let $e:\Gamma\rightarrow N$ be a function defined by $e(0):=0$, $e(1):=1$, and the extension to to infinite sequences $t=(t_{n})_{n\geq 0}$ over $\Gamma$ by $e(t):=\sum_{i=0}^{\infty}e(t_{i})$. As in each configuration of $U$ only a finite number of cells consist of no blank symbol $(0)$, $e(t)$ is always finite and well-defined. Notice that we can always obtain the symbol that is closest to the head by $e(t)\mod 2$ (the symbol at the head position in the case of $t_{L}$ and the symbol right of the head position in the case of $t_{R}$). By multiplying or dividing the encoding $e(t)$ by $2$, each side can be lengthened or shortened, respectively. The encoding of configurations of $U$ indirectly referred to in this paper is defined by
$\operatorname{enc}_{NE}(q_{i},t_{L},t_{R})=0^{7}0^{e(t_{R})}\#0^{7}0^{e(t_{L})}\#0^{i+6}$
for every configuration $(q_{i},t_{L},t_{R})$. Recall that $i>0$ as $q_{i}\in\{q_{1},...,q_{15}\}$. A computation $\mathcal{C}=(C_{1},...,C_{n})$ on $U$ is a finite sequence of configurations of $U$. It is valid if $C_{1}=I$ (II being some initial configuration), $C_{n}$ is a halting configuration, and $C_{i+1}$ is a valid successor configuration of $C_{i}$, for $i\in[n-1]$, as defined by $\delta$. In [BREMER201215], the notion is adopted that any possible configuration where both tape sides have a finite value under $e$ is a valid successor configuration of a halting configuration. The encoding of computations of $U$ is given analogously to the definition in the case of nondeterministic 2-counter automata without input, i.e., for some computation $\mathcal{C}=(C_{1},...,c_{n})$, we have
$\operatorname{enc}_{NE}(\mathcal{C})=\#\#\operatorname{enc}_{NE}(C_{1})\#\#\operatorname{enc}_{NE}(c_{2})\#\#\ ...\ \#\#\operatorname{enc}_{NE}(C_{n})\#\#.$
Finally, also analogous to nondeterministic 2-counter automata without input, let
$\mathtt{ValC}_{U}(I)=\{\operatorname{enc}_{NE}(\mathcal{C})\mid\mathcal{C}\text{ is a valid computation from }I\ \}.$
$q_{1}$ $q_{2}$ $q_{3}$ $q_{4}$ $q_{5}$ $q_{6}$ $q_{7}$ $q_{8}$
0 $(0,R,q_{2})$ $(1,R,q_{3})$ $(0,L,q_{7})$ $(0,L,q_{6})$ $(1,R,q_{1})$ $(1,L,q_{4})$ $(0,L,q_{8})$ $(1,L,q_{9})$
11 $(1,R,q_{1})$ $(1,R,q_{1})$ $(0,L,q_{5})$ $(1,L,q_{5})$ $(1,L,q_{4})$ $(1,L,q_{4})$ $(1,L,q_{7})$ $(1,L,q_{7})$
$q_{9}$ $q_{10}$ $q_{11}$ $q_{12}$ $q_{13}$ $q_{14}$ $q_{15}$
0 $(0,R,q_{1})$ $(1,L,q_{11})$ $(0,R,q_{12})$ $(0,R,q_{13})$ $(0,L,q_{2})$ $(0,L,q_{3})$ $(0,R,q_{14})$
11 $(1,L,q_{10})$ $\mathtt{HALT}$ $(1,R,q_{14})$ $(1,R,q_{12})$ $(1,R,q_{12})$ $(0,R,q_{15})$ $(1,R,q_{14})$
Table 4: Transition table of $U$, i.e., definition of $\delta$, as it is given in [BREMER201215] or [Neary2009].
0.B.3 Undecidability of the Inclusion Problem for Several Basic Relations
Proof of Corollary˜1:
Proof
Consider the construction of Freydenberger and Reidenbach [Freydenberger2010] and the construction of Bremer and Freydenberger [BREMER201215], respectively. The main differences between E-pattern languages $L_{E}(\alpha)$ (resp. NE pattern languages $L_{NE}(\alpha)$) of a pattern $\alpha\in Pat_{\Sigma}$ in [Freydenberger2010, BREMER201215]and relational pattern languages $L_{E}(\alpha,r_{\alpha})$ (resp. $L_{NE}(\alpha,r_{\alpha})$) that we consider are
  1. 1.
    In relational patterns each variable may occur only once whereas in classical patterns each variable may occur any number of times.
  2. 2.
    In relational pattern languages the substitution must be $r_{\alpha}$-valid whereas in pattern languages there is no relation.
Thus, if we use the pattern given in relational pattern languages to simulate identical variables, there is no significant difference between relational pattern languages and classical pattern languages. With this, we can immediately refer to the proofs in [Freydenberger2010, BREMER201215]. First, consider $r_{=}$. We can use the proofs in [Freydenberger2010, BREMER201215]. For each variable $x$, we replace each multiple occurrence of $x$ by new variables $x_{i},i\in\mathbb{N}$ with $(r_{=},x,x_{i})\in r_{\alpha}$. Since $(x,x_{i})\in r_{\alpha}$, we get $h(x)=h(x_{i})$ for each $r_{\alpha}$-valid substitution $h$. As $x_{i}$ is a new variable, we do not change anything else. Thus, this slightly adapted proof works in exactly the same way as the original proof and we get that the inclusion problem for relational pattern languages with the relation $r_{=}$ is undecidable.
Now, we show that we can do the same with the relations $r_{ssq}$ and $r_{*}$. For $r_{ssq}$, we get $(r_{ssq},x,y),(r_{ssq},y,x)\in r_{\alpha}$ implies $h(x)=h(y)$ for any two variables $x,y$ and any $r_{\alpha}$-valid substitution $h$ by the definition of $r_{ssq}$. For each variable $x$, we replace each multiple occurrence of $x$ by new variables $x_{i},i\in\mathbb{N}$ with $(r_{ssq},x,x_{i}),(r_{ssq},x_{i},x)\in r_{\alpha}$. With this, the proof works analogously to the one of [Freydenberger2010, BREMER201215]. For $r_{*}$, we get the same observation: $(r_{*},x,y),(r_{*},y,x)\in r_{\alpha}$ implies $h(x)=h(y)$ for any two variables $x,y$ and any $r_{\alpha}$-valid substitution $h$. Thus, the proof works here as well.
So, in total, we get that the inclusion problem for relational pattern languages with the relation $\tilde{r}$ for $\tilde{r}\in\{r_{=},r_{ssq},r_{*}\}$ is undecidable.
0.B.4 Undecidability of the Inclusion Problem for Non-Erasing Relational Pattern Languages under Abelian Equivalence
Proof
As this proof works almost identical to the proof of Bremer and Freydenberer in [BREMER201215], we only give an extended sketch and focus on the formal intricacies that occur as we are now considering abelian equivalence instead of the equivalence relation.
Let $\Sigma=\{0,\#\}$ and $R:=\{r_{ab}\}$. To show undecidability of this problem, we reduce the problem of deciding whether $\mathtt{UnivValc}_{NE}(I)=\emptyset$ for some initial configuration $I$ of $U$ to the problem of deciding whether $L_{NE}(p_{1},\alpha_{R})\subseteq L_{NE}(p_{2},\beta_{R})$ for two relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. To be specific, we construct two such patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ such that $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$.
We define the relational pattern $(\beta,r_{\beta})$ by
$\beta:=a_{1}b_{1}\ \#^{5}\ a_{2}x_{1,1}x_{2,1}\cdots x_{\mu,1}b_{2}\ \#^{5}\ r_{1}\hat{\beta}_{1}r_{2}\hat{\beta}_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1}$
where for all $i\in\{1,...,\mu\}$ we have
$\hat{\beta}_{i}:=0x_{i,2}\cdots x_{i,5}0\ \upgamma_{i}\ 0x_{i,6}\cdots x_{i,9}0\ \updelta_{i}\ 0x_{i,10}\cdots x_{i,13}0$
such that $a_{1},a_{2},b_{1},b_{2},r_{i},x_{j,k}$ are variables for all $i\in\{1,...,\mu+1\}$, $j\in\{1,...,\mu\}$, and $k\in\{1,...,13\}$ with $(r_{ab},a_{1},a_{2}),(r_{ab},b_{1},b_{2}),(r_{ab},x_{i,j},x_{i,j^{\prime}})\in r_{\beta}$, for $i\in\{1,...,\mu\}$ and $j,j^{\prime}\in\{1,...,13\}$. Also, $\upgamma_{i}\updelta_{i}\in\mathtt{RelPat}_{\Sigma,R}$ are patterns for all $i\in\{1,...,\mu\}$ which are described later, but for which we already mention that the factor $^{4}$ never appears in them. Additionally, we have that each variable $r_{i}$ is not related to any other variable in $\beta$ and we have that the variables $a_{1},a_{2},b_{1},$ and $b_{2}$ are not related to any other variable than the ones mentioned here. For the relational patterns $\upgamma_{i}$ and $\updelta_{i}$ we also mention that all variables occurring in $\upgamma_{i}\updelta_{i}$ are not related to any other variable occurring in some other $\upgamma_{j}\updelta_{j}$ for $i\neq j$ with $i,j\in\{1,...,\mu\}$ and that each variable $x_{i,j}$ is not related to any other variable in $\upgamma_{k}\updelta_{k}$ for $i,k\in\{1,...,\mu\}$ and $j\in\{1,...,13\}$.
Now, we define the second relational pattern $(\alpha,r_{\alpha})$ by
$\alpha:=0^{\mu+1}\ \#^{5}\ 0^{\mu}\#0^{\mu}\ \#^{5}\ tv0\alpha_{1}0v0\alpha_{2}0vt$
where $v:=0\#^{4}0$, $\alpha_{1}$ is some relational pattern not containing $^{3}$ as a factor, $\alpha_{2}$ is a relational pattern not containing #\#, and $t$ is some other terminal string. As we shall see later, we have $r_{\alpha}=\emptyset$. This will become clear once we specifically define $\alpha_{1}$ and $\alpha_{2}$. We define $t$ using the non-erasing substitution $\psi:(\operatorname{var}(\beta)\cup\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\psi(x)=0$ for all $x\in\operatorname{var}(\beta)$. Using that, we say $t:=\psi(r_{1}\hat{\beta}_{1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$. We see that both patterns are basically constructed in the same way as in [BREMER201215], just using the formalism of relational patterns, i.e. all variables which are equal in the construction in [BREMER201215], are in this construction distinct variables that are, however, related to each other through the sets $r_{\alpha}$ and $r_{\beta}$.
The following lemma follows directly from the proof in [BREMER201215] which can be converted directly to the formalism using substitutions over relational patterns.
Lemma 6
[BREMER201215] All $\psi(\hat{\beta}_{i})$ with $1\leq i\leq\mu$ and $t$ begin and end with the terminal letter 0 and do not contain the terminal word $^{4}$ as a factor.
We recall, that $H_{r_{\alpha}}$ denotes the set of all $r_{\alpha}$-valid substitutions. Let $H_{r_{\alpha}^{+}}$ denote the subset of all non-erasing substitutions of $H_{r_{\alpha}}$. Each pair $(\upgamma_{i},\updelta_{i})$ is interpreted as a predicate $\pi_{i}:H_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ such that any $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{i}$ if there exists some $r_{\beta}$-valid non-erasing substitution $\tau:(\operatorname{var}(\upgamma_{i}\updelta_{i})\cap\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\tau(\upgamma_{i})=h(0\alpha_{1}0)$ and $\tau(\updelta_{i})=h(0\alpha_{2}0)$. We need to say that $\tau$ is $r_{\beta}$-valid for formal reasons, but as all variables in $\updelta_{i}\upgamma_{i}$ are only related to variables in $\updelta_{i}\upgamma_{i}$ itself, we can consider such an isolated substitution as it does not interfere with the substitution of any other variable of $\beta$. Using those predicates, we will see that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})$ contains those words $h(\alpha)$ for $h\in H_{r_{\alpha}^{+}}$ for which we have that $h$ does not satisfy any predicate $\pi_{i}$. Essentially, all predicates $\pi_{1}$ to $\pi_{\mu}$ will give an exhaustive list of criteria for membership of $\overline{\mathtt{ValC}_{U}(I)}$. We proceed with more technical preperations that we need for the easy and dynamic construction of predicates. Once this is settled, as the constructed predicates from [BREMER201215] also work in this case, we are done.
We say that a substitution $h\in H_{r_{\alpha}^{+}}$ is of bad form if $h(0\alpha_{1}0)$ contains a factor $^{3}$ or if $h(0\alpha_{2}0)$ contains the letter #\#. We use the predicates $\pi_{1}$ and $\pi_{2}$ to handle all such substitutions of bad form and construct them exactly the same way as they are constructed in [BREMER201215] by saying
$\upgamma_{1}:=y_{1,1}\#^{k}y_{1,2}$ $\upgamma_{2}:=0y_{2}0$
$\updelta_{1}:=0\hat{y}_{1}0$ $\updelta_{2}:=\hat{y}_{2,1}\#\hat{y}_{2,2}$
for new variables $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2}\in X$ that are not related to any other variable in $\beta$. As they are not related to any other variable, no intricacies regarding abelian equivalence in comparison to normal equality occur and we directly get the following lemma from [BREMER201215] as their proof needs no adaptation.
Lemma 7
[BREMER201215] A substitution $h\in H_{r_{\alpha}^{+}}$ is of bad-form if and only if $h$ satisfies $\pi_{1}$ or $\pi_{2}$.
The following key lemma allows us to focus only on predicates from now on and is independent of the shape of the specific predicates $\pi_{3}$ to $\pi_{\mu}$.
Lemma 8
For every non-erasing substitution $h\in H_{r_{\alpha}^{+}}$, we have that $h(\alpha)\in L_{NE}(\beta,r_{\beta})$ if and only if $h$ satisfies some predicate $\pi_{1}$ to $\pi_{\mu}$.
Proof
Due to the different formalism, we need to rewrite certain parts of the proof from [BREMER201215]. However, the general idea is the same. We just have to show that it also works for abelian equivalence.
($\Rightarrow:$) Assume some $h\in H_{r_{\alpha}^{+}}$ satisfies some predicate $\pi_{i}$ for $i\in\{1,...,\mu\}$. Then there exists a non-erasing $r_{\beta}$-valid substitution $\tau:(\operatorname{var}(\upgamma_{i}\updelta_{i})\cup\Sigma)^{*}\rightarrow\Sigma^{*}$ with $\tau(\upgamma_{i})=h(0\alpha_{1}0)$ and $\tau(\updelta_{i})=h(0\alpha_{2}0)$. We extend $\tau$ to a non-erasing $r_{\beta}$-valid substitution $\tau^{\prime}\in H_{r_{\beta}^{+}}$ by
  1. 1.
[…]
  2. 2.
    $\tau^{\prime}(x_{j,k}):=\begin{cases}\#&\text{ for }j=i,k\in\mathbb{N}\\ 0&\text{ for }j\neq i,k\in\mathbb{N}\end{cases}$
  3. 3.
    $\tau^{\prime}(r_{j}):=\begin{cases}\psi(r_{i}\hat{\beta}_{i}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})&\text{ for }j=i\\ \psi(r_{1}\hat{\beta}_{1}\cdots r_{i}\hat{\beta}_{i}r_{i+1}&\text{ for }j=i+1)\\ 0&\text{ else, }\end{cases}$
  4. 4.
    $\tau^{\prime}(a_{k}):=0^{\mu-i+1}$ for $k\in\{1,2\}$, and
  5. 5.
    $\tau^{\prime}(b_{k}):=0^{i}$ for $k\in\{1,2\}$.
Notice that none of the variables in one factor $\operatorname{var}(\upgamma_{i}\updelta_{i})$ appear outside of $\upgamma_{i}$ and $\updelta_{i}$. By that, we can always define $\tau^{\prime}$ that way. Also note, that the only significant difference to [BREMER201215] lies in the definition of $\tau^{\prime}(x_{j,k})$, $\tau^{\prime}(a_{k})$, and $\tau^{\prime}(b_{k})$ as no variables can occur multiple timesin relational patterns; the rest remains the same.
As in [BREMER201215] we obtain $\tau^{\prime}(\upgamma_{i})=\tau(\gamma_{i})=h(0\alpha_{1}0)$, $\tau^{\prime}(\updelta_{i})=\tau(\updelta_{i})=h(0\alpha_{1}0)$, and
$\tau^{\prime}(a_{1}b_{1}\#^{5}a_{2}x_{1,1}\cdots x_{\mu,1}b_{2}\#^{5})$ $=0^{\mu-i+1}0^{i}\#^{5}0^{\mu-i+1}0^{i-1}\#0^{\mu-i}0^{i}\#^{5}$
$=0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}$
as well as
$\tau^{\prime}(\hat{\beta}_{i})$ $=\tau^{\prime}(0x_{i,2}\cdots x_{i,5}0\upgamma_{i}0x_{i,6}\cdots x_{i,9}0\updelta_{i}0x_{i,10}\cdots x_{i,13}0)$
$=0\#^{4}0\tau^{\prime}(\upgamma_{i})0\#^{4}0\tau^{\prime}(\updelta_{i})0\#^{4}0$
$=vh(0\alpha_{1}0)vh(0\alpha_{2}0)v.$
Also, as $\tau^{\prime}(x)=\psi(x)$ for all $x\in\operatorname{var}(\beta)_{j}$ for $j\neq i$, we get for all $j\neq i$ that $\tau^{\prime}(\hat{\beta}_{j})=\psi(\hat{\beta}_{j})$ and by that we finally obtain
$\tau^{\prime}(\beta)=$ $\ \tau^{\prime}(a_{1}b_{2}\#^{5}a_{2}x_{1,1}\cdots x_{\mu,1}\#^{5}r_{1}\hat{\beta}_{1}r_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ 0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}\tau^{\prime}(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ 0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}\tau^{\prime}(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{i-1}\hat{\beta}_{i-1})\tau^{\prime}(r_{i})\tau^{\prime}(\hat{\beta}_{i})\tau^{\prime}(r_{i+1})$
$\ \tau^{\prime}(\hat{\beta}_{i+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ 0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}\psi(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{i-1}\hat{\beta}_{i-1})\psi(r_{i}\hat{\beta}_{i}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})\tau^{\prime}(\hat{\beta}_{i})$
$\ \psi(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{i}\hat{\beta}_{i}r_{i+1})\psi(\hat{\beta}_{i+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ 0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}\psi(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})\tau^{\prime}(\hat{\beta}_{i})\psi(r_{1}\hat{\beta}_{i}r_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ 0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}t\tau^{\prime}(\hat{\beta}_{i})t$
$=$ $\ 0^{\mu+1}\#^{5}0^{\mu}\#0^{\mu}\#^{5}tvh(0\alpha_{1}0)vh(0\alpha_{2}0)vt$
$=$ $\ h(\alpha).$
By that we get $h(\alpha)\in L_{NE}(\beta,r_{\beta})$.
($\Leftarrow$): Now, assume $h(\alpha)\in L_{NE}(\beta,r_{\beta})$. In the case of $h$ being of bad form, we know by Lemma 7 that $h$ satisfies $\pi_{1}$ or $\pi_{2}$. So, assume $h(0\alpha 0)$ does not contain $^{3}$ as a factor and that $h(0\alpha_{2}0)$ has no occurrence of the letter #\#, hence being of the form $h(0\alpha_{2}0)\in 00^{+}0$. As $h(\alpha)\in L_{NE}(\beta,r_{\beta})$, there exists some non-erasing $r_{\beta}$-valid substitution $\tau\in H_{r_{\beta}^{+}}$ such that $\tau(\beta)=h(\alpha)$.
As $h$ is of good form and as $t$ begins and ends with 0 and does not contain any occurrence of $^{4}$ by Lemma 6, we know that $h(\alpha)$ contains the factor $^{5}$ exactly twice. That is also the case for $\tau(\beta)$ as $h(\alpha)=\tau(\beta)$. Hence, we can decompose the equation $h(\alpha)=\tau(\beta)$ into the following system consisting of the equations
$0^{\mu+1}$ $=\tau(a_{1}b_{1})$ (1)
$0^{\mu}\#0^{\mu}$ $=\tau(a_{2}x_{1,1}\cdots x_{\mu,1}b_{2})$ (2)
$tvh(0\alpha_{1}0)vh(0\alpha_{2}0)vt$ $=\tau(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1}).$ (3)
We see that in equation $2$ each $x_{i,1}$ for $i\in\{1,...,\mu\}$ we have that it has to be substituted by a single symbol, as otherwise equation $1$ would be unsatisfied, because $(r_{ab},a_{1},a_{2}),(r_{ab},b_{1},b_{2})\in r_{\beta}$. Also, we get by that that $a_{2}$ and $b_{2}$ can also be substituted to strings consisting of only $0^{\prime}s$. So, for some single $j\in\{1,...,\mu\}$ we have $\tau(x_{j})=\#$ and for all other $j^{\prime}\neq j$ we have $\tau(x_{j}^{\prime})=0$. Now, as $(r_{ab},x_{j,k},x_{j,k^{\prime}})\in r_{\beta}$ for $k,k^{\prime}\in\{1,...,13\}$, we get for that specific $j$ that
$\tau(\hat{\beta}_{j})$ $=\tau(0x_{j,2}\cdots x_{j,5}0\upgamma_{i}0x_{j,6}\cdots x_{j,9}0\updelta_{j}0x_{j,10}\cdots x_{j,13}0)$
$=0\#^{4}0\tau(\upgamma_{i})0\#^{4}0\tau(\updelta_{i})0\#^{4}0$
$=v\tau(\upgamma_{i})v\tau(\updelta_{j})v.$
This can be used to convert the right side of equation $3$ in the following manner.
$\ \tau(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ \tau(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{j-1}\hat{\beta}_{j-1}r_{j})\tau(\hat{\beta}_{j})\tau(r_{j+1}\hat{\beta}_{j+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
$=$ $\ \tau(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{j-1}\hat{\beta}_{j-1}r_{j})v\tau(\upgamma_{i})v\tau(\updelta_{j})v\tau(r_{j+1}\hat{\beta}_{j+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
This results in the following adaptation of equation $3$.
$\ tvh(0\alpha_{1}0)vh(0\alpha_{2}0)vt$
$=$ $\ \tau(r_{1}\hat{\beta}_{1}r_{2}\cdots r_{j-1}\hat{\beta}_{j-1}r_{j})v\tau(\upgamma_{i})v\tau(\updelta_{j})v\tau(r_{j+1}\hat{\beta}_{j+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
We know that $h$ is of good form and we know that $t$ does not contain any factor $^{4}$. So, the left side of the equation contains exactly $3$ occurrences of $^{4}$ as a factor, that is, in each occurrernce of one $v$. The right side of the equation also contains exactly three occurrences of $v$ which themselves make up $3$ occurrences of $^{4}$. Thus, we obtain the following system of equations.
$t$ $=\tau(r_{1}\hat{\beta}_{1}\cdots r_{j-1}\hat{\beta}_{j-1}r_{j})$
$h(0\alpha_{1}0)$ $=\tau(\upgamma_{j})$
$h(0\alpha_{2}0)$ $=\tau(\updelta_{j})$
$t$ $=\tau(r_{j+1}\hat{\beta}_{j+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
As these equations must hold and especially as $h(0\alpha_{1}0)=\tau(\upgamma_{j})$ and $h(0\alpha_{2}0)=\tau(\updelta_{j})$, we know that $h$ satisfies the predicate $\pi_{j}$. ∎
By that, we can now select appropriate predicates $\pi_{1}$ to $\pi_{\mu}$ such that we obtain $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$. Lemma 8 shows us that if for some $h\in H_{r_{\alpha}^{+}}$ and $\tau\in H_{r_{\beta}^{+}}$ with $h(\alpha)=\tau(\beta)$ we have that $h$ is of good form, then we can pick exactly one $i\in\{1,...,\mu\}$ such that we fulfill $\tau(0x_{i,j}\cdots x_{i,j+3}0)=0\#^{4}0=v$ for $j\in\{2,6,9\}$.
For technical reasons, we define a predicate that sets a lower bound to the length of $h(\alpha_{2})$ if unsatisfied for a substitution $h\in H_{r_{\alpha}^{+}}$ of good form. As in [BREMER201215], the predicate $\pi_{3}$ is defined by
$\upgamma_{3}$ $:=y_{3,1}\hat{y}_{3,1}y_{3,2}\hat{y}_{3,2}y_{3,3}\hat{y}_{3,3}y_{3,4}$
$\updelta_{3}$ $:=0\hat{y}_{3,1}\hat{y}_{3,2}\hat{y}_{3,3}$
for pairwise non-related variables $y_{3,1},...,y_{3,4}$ and $\hat{y}_{3,1},...,\hat{3,3}$. As in [BREMER201215], if some $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{3}$, then $h(\alpha_{2})$ is a concatenation of three non-empty factors of $h(\alpha_{1})$. In other words, if some $h$ does not satisfy any of $\pi_{1}$$\pi_{2}$, and $\pi_{3}$, then $h(\alpha_{2})\in 0^{+}$ has to be longer than the $3$ longest non-overlapping and non-touching factors of 0’s in $h(\alpha_{1})$. This restriction allows for a simpler way to construct all other predicates $\pi_{4}$ to $\pi_{\mu}$.
As we will see, the next part regarding simple predicates is also very similar to the one defined and used in [BREMER201215] but had to be adapted formally to also work in the context of relational patterns. Let $X_{s}:=\{\hat{x}_{1,1},\hat{x}_{2,1},\hat{x}_{3,1},\hat{x}_{1,2},\hat{x}_{2,2},\hat{x}_{3,2},...\}\subset X$ and always assume that $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ in any set of relational constraints $r_{p}$ of any relational pattern $(p,r_{p})\in\mathtt{RelPat}_{\Sigma,R}$ if any variable of $X_{s}$ is used in $p$. Let $G_{r_{\alpha}^{+}}\subset H_{r_{\alpha}^{+}}$ denote the subset of $r_{\alpha}$-valid substitutions that are of good form and let $S$ be the set of all non-erasing substitutions $s:(\Sigma\cup X_{s})^{*}\rightarrow\Sigma^{*}$ for which $s(\hat{x}_{i,j})\in\{0\}^{+}$, for all $i\in\{1,2,3\}$ and $j\in\mathbb{N}$, and for which we have $s(\hat{x}_{i,j})=s(\hat{x}_{i^{\prime},j^{\prime}})$ if $i=i^{\prime}$ for $i,i^{\prime}\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$. For any relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ we define $S(p_{s},r_{p_{s}}):=\{s(p_{s})|s\in S\}$. The next definition is exactly the same as in [BREMER201215], just adapted to the notation of relational patterns.
Definition 1
[BREMER201215] A predicate $\pi:G_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ is called a NE-simple predicate for $0\alpha_{1}0$ if there exists a relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ and $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p_{s}}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ and languages $L_{1}\in\{0\Sigma^{*},\{0\}\}$ and $L_{2}\in\{\Sigma^{*}0,\{0\}\}$ such that any non-erasing substitution $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi$ if and only if $h(0\alpha_{1}0)\in L_{1}S(p_{s},r_{p_{s}})L_{2}$. If $L_{1}=0\Sigma^{*}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ an infix-predicate. If $L_{1}=\{0\}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ a prefix-predicate in if its the other way around, we call $\pi$ a suffix predicate.
We understand elements of $X_{s}$ as numerical paramters that describe (concatenative) powers of 0 with non-erasing substitutions $s:(\Sigma\cup X_{s})^{*}\rightarrow\Sigma^{*}$ acting as assignments. The next lemma is the final step we need to take in order to finish the adaptation of the original proof to the new case. It shows that for each simple predicate $\pi_{s}$, we can construct a predicate $\pi$ defined over some $\upgamma$ and $\updelta$ such that each substitution of good form that would have satisfied $\pi_{s}$ now satisfies either $\pi_{3}$ or $\pi_{s}$ which suffices the demands of the construction. As we can see, the argument is again very similar to the one in [BREMER201215] and had to be adapted mainly to the new formalism.
Lemma 9
For every NE-simple predicate $\pi_{s}$, there exists a predicate $\pi$ defined by relational patterns $(\upgamma,r_{\upgamma})$ and $(\updelta,r_{\updelta})$ which can be embedded in $(\beta,r_{\beta})$ such that for all non-erasing substitutions $h\in G_{r_{\alpha}^{+}}$ we have
  1. 1.
    if $h$ satisfies $\pi_{s}$, then $h$ also satisfies $\pi_{3}$ or the newly constructed $\pi$, and
  2. 2.
    if $h$ satisfied the newly constructed $\pi$, then $h$ also satisfies $\pi_{s}$.
Proof
First, consider the case of $L_{1}=0\Sigma^{*}$ and $L_{2}=\Sigma^{*}0$. Assume $\pi_{s}$ is a NE-simple predicate and that $(p_{s},r_{p_{s}})$ is a relational pattern with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ and $r_{p_{s}}:=\{(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})|\hat{x}_{i,j},\hat{x}_{i,j^{\prime}}\in X_{s};i\in\{1,2,3\};j,j^{\prime}\in\mathbb{N}\}$ such that $h\in G_{r_{\alpha}^{+}}$ satisfies $\pi_{s}$ if and only if $h(0\alpha_{1}0)\in L_{1}S(p_{s},r_{p_{s}})L_{2}$. Then we can define the relational patterns $(\upgamma,r_{\upgamma})$ by $\upgamma=y_{1}p_{s}y_{2}$ for two new and to no other variable related variables $y_{1}$ and $y_{2}$ and set $r_{\upgamma}=r_{p_{s}}$. Additionally, we define the relational pattern $(\updelta,r_{\updelta})$ by $\updelta=0\theta\hat{y}0$ where $\theta:=\{\hat{i,y}|\exists\hat{x}_{i,j}\in\operatorname{var}(p_{s})\text{ and }y=|\operatorname{var}(p_{s})|+1\}$ and $\hat{y}$ is a new to no other variable related variable, and by setting $r_{\updelta}=r_{\upgamma}$ (notice that for as long as we do not merge the relational pattern $(\updelta,r_{\updelta})$ with any other relational pattern, that no variable in $\updelta$ is related to each other, however, if we merge $(\updelta,r_{\updelta})$ with $(\upgamma,r_{\upgamma})$, then the corresponding $\hat{x}_{i,j}$ variables become related to each other. Also notice, that each variable in either of those two relational patterns will not be related to any other variable in $\beta$, allowing for them to be directly embedded into $(\beta,r_{\beta})$ without further steps).
(1.) Assume that $h\in G_{r_{\alpha}^{+}}$ satisfies $\pi_{s}$. Then there exist words $w_{1}\in 0\Sigma^{*}$ and $w_{2}\in\Sigma^{*}0$ and a non-erasing substitution $s\in S$ such that $h(0\alpha_{1}0)=w_{1}s(p_{s})w_{2}$. If $h(\alpha)_{2}$ is shorter than the $3$ longest non-overlapping and non-touching factors of the form $0^{+}$ in $h(\alpha_{1})$ concatenated, then $\pi_{3}$ is satisfied. If that is not the case, we define a substitution $\tau$ by setting $\tau(y_{1}):=w_{1}$, $\tau(y_{2})=w_{2}$ and $\tau(\hat{x}_{i,j}):=s(\hat{x}_{i,j})$ for all $i\in\{1,2,3\}$ and $j\in\mathbb{N}$. We also set $\tau(\hat{y})=0^{m}$ where
[…]
As $h$ does not satisfy $\pi_{3}$, we have $m>0$. We obtain
$\tau(\upgamma)$ $=\tau(y_{1})\tau(p_{s})\tau(y_{2})=w_{1}s(p_{s})w_{2}=h(0\alpha_{1}0)$
$\tau(\updelta)$ $=00^{|h(\alpha_{2})|}0=h(0\alpha_{2}0).$
By that, we see that $h$ also satisfies the constructed predicate $\pi$.
(2.) Now assume that $h\in G_{r_{\alpha}^{+}}$ satisfies the constructed predicate $\pi$ (assuming it is embedded in $\beta$). Hence, there exists a non-erasing substitution $\tau\in H_{r_{\beta}^{+}}$ with $h(0\alpha_{1}0)=\tau(\upgamma)$ and $h(0\alpha_{2}0)=\tau(\updelta)$. We get by the construction that $\tau(y_{1})\in 0\Sigma^{*}$ and $\tau(y_{2})\in\Sigma^{*}0$. Now define $s(\hat{x}_{i,j})=\tau(\hat{x}_{i,j})$ for all $\hat{x}_{i,j}\in\operatorname{var}(\updelta)$. We see that $h(0\alpha_{1}0)\in L_{1}S(p_{s},v_{R})L_{2}$ holds. So, $h$ must satisfy the simple predicate $\pi_{s}$.
The other cases, where $L_{1}$ or $L_{2}$ is chosen to be the language containing only the word 0, are handled analogously by omitting $y_{1}$ or $y_{2}$ and keeping the rest of the construction. ∎
So, in general, if $h$ is of good form but does not satisfy $\pi_{3}$, then $h(\alpha_{2})\in 0^{+}$ is long enough to provide building blocks for NE-simple predicates that use variables from $X_{s}$.
Lemma 9 shows us that the construction of NE-simple predicates in [BREMER201215] can be used in the setting of relational patterns under abelian equivalence in exactly the same way. Mainly, this is caused by the fact that each variable $\hat{x}_{i,j}\in X_{s}$ is substituted by a unary string. By that, each related variable has to be substituted exactly in the same way as they are substituted in [BREMER201215], i.e. equally. Having shown that the same simple predicates from [BREMER201215] also work in this setting in exactly the same way, we define $\pi_{4}$ to $\pi_{\mu}$ exactly as they are defined in [BREMER201215], as all of those predicates are given described as simple predicates.
We set $\alpha_{1}:=\#\#enc_{NE}(I)\#\#x\#0^{6}0^{1}0\#\#$ and $\alpha_{2}:=y00$ for pairwise distinct variables $x$ and $y$ (notice, that $x$ and $y$ are the only variables in the pattern $\alpha$, hence for $(\alpha,r_{\alpha})$ we have $r_{\alpha}=\emptyset$). By the construction of $\pi_{1}$ to $\pi_{\mu}$ and essentially the same argument over the embedding of all invalid computations in $(\beta,r_{\beta})$ from [BREMER201215], we obtain that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$. This concludes the case $|\Sigma|=2$.
For the case of larger alphabets, i.e. alphabets of size $|\Sigma|\geq 3$, the exact construction in [BREMER201215] can also be adapted to the formality of our setting and works by the same arguments as in [BREMER201215]. The relational patterns $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ need to be slightly adapted to work with larger alphabets. Essentially, suffixes to both the patterns $\alpha$ and $\beta$ are added to allow for another new variable $\hat{x}$ to be mapped to any letter of $\Sigma$ (similar to how we ensured that one variable $x_{i,j}$ is mapped to #\#). Then, $2$ new predicates are added to ensure that if this letter occurs in any substitution of $h(\alpha_{1})$ or $h(\alpha_{2})$, then one of these predicates is satisfied.
[…]
$\tilde{\alpha}$ $:=\alpha\#^{5}0\mathtt{a}_{1}\mathtt{a}_{2}\cdots\mathtt{a}_{\sigma}0\#^{5}0\mathtt{a}_{1}\mathtt{a}_{2}\cdots\mathtt{a}_{\sigma}0\#^{5},$
$\tilde{\beta}$ $=\beta\#^{5}\tilde{y}_{1}\tilde{x}_{1,1}\tilde{z}\#^{5}\tilde{y}_{2}\tilde{x}_{2,1}\tilde{z}_{2}\#^{5}$
for new and pairwise non-related variables $\tilde{y}_{1},\tilde{y}_{2},\tilde{x}_{1,1},\tilde{x}_{2,1},\tilde{z}_{1},\tilde{z_{2}}$. We mention here that also the variables $\tilde{x}_{1,2}$ and $\tilde{x}_{2,2}$ will occur in $\tilde{\beta}$. For those, we have $(r_{ab},\tilde{x}_{1,1},\tilde{x}_{1,2}),(r_{ab},\tilde{x}_{2,1},\tilde{x}_{2,2})\in r_{\tilde{\beta}}$. To work with the new construction and the two new predicates defined below, we need an extended version of the definition of $\psi:(\operatorname{var}{\hat{\beta}_{1}...\hat{\beta}_{\mu+2}})\rightarrow\Sigma^{*}$. Now, we also have $\psi(\tilde{x}_{1})=\psi(\tilde{x}_{2})=\mathtt{a}_{1}\cdots\mathtt{a}_{\sigma}$ and $\psi(x)=0$ for all $x\in\operatorname{var}(\hat{\beta}_{1}\cdots\hat{\beta}_{\mu+1})\setminus\{\tilde{x}_{1},\tilde{x}_{2}\}$. Finally, we define an adapted version of the two new predicates $\pi_{\mu+1}$ and $\pi_{\mu+2}$ by
$\upgamma_{\mu+1}:=y_{\mu+1,1}\ \tilde{x}_{1,2}\ y_{\mu+1,2}$ $\upgamma_{2}:=0\ y_{\mu+2}\ 0$
$\updelta_{\mu+1}:=0\ \hat{y}_{\mu+1}\ 0$ $\updelta_{2}:=\hat{y}_{\mu+2,1}\ \tilde{x}_{2,2}\ \hat{y}_{\mu+2,2}$
where $y_{\mu+1,1},y_{\mu+1,2},y_{\mu+2},\hat{y}_{\mu+1},\hat{y}_{\mu+2,1},$ and $\hat{y}_{\mu+2,2}$ are new and pairwise distinct and unrelated variables. Remember, that we have $(r_{ab},\tilde{x}_{1,1},\tilde{x}_{1,2})\in r_{\tilde{\beta}}$ and $(r_{ab},\tilde{x}_{2,1},\tilde{x}_{2,2})\in r_{\tilde{\beta}}$. Now, using these constructions, we obtain the exact argument as in [BREMER201215] to obtain the final result. The main argument why this construction can be used by the same argument is that it works over unary substitutions of $\tilde{x}_{1,1},\tilde{x}_{1,2},\tilde{x}_{2,1},$ and $\tilde{x}_{2,2}$ of length $1$. Hence, using substitutions of this type, abelian equivalence behaves exactly the same as equality between variables. Thus, we omit the details at this point and refer to [BREMER201215] for a more detailed proof.
This concludes the case of larger alphabets of size $|\Sigma|\geq 3$. We obtain undecidability of the inclusion problem for non-erasing relational pattern languages under abelian equivalence for alphabets of size $|\Sigma|\geq 2$. ∎
0.B.5 Undecidability of the Inclusion Problem for Erasing Relational Pattern Languages under the Reversal Relation
Proof of Theorem˜4.1:
Proof
The proof is based on the construction by Freydenberger and Reidenbach [Freydenberger2010]. For better understanding for those who don’t know the construction and for a really detailed verification, we copied the proof of Theorem 2 of [Freydenberger2010] and adapted it at several points. The main difference is the definition of $u$ and the definition of the delimiter. Why these adaptations work is shorter described in the proof of Theorem˜4.2 for those who already are familiar with the construction. The adaptations of Theorem˜4.1 and Theorem˜4.2 are very similar.
We begin with the case $|\Sigma|=2$, so let $\Sigma:=\{0,\#\}$. Let $A:=(Q,\delta,q_{0},F)$ be a nondeterministic 2-counter automaton; w.l.o.g. let $Q:=\{q_{0},\ldots,q_{s}\}$ for some $s\in\mathbb{N}_{0}$. Our goal is to construct patterns $\alpha_{A},\beta_{A}\in\mathtt{RelPat}_{\Sigma,R}$ such that $L_{E}(\alpha_{A},r_{\alpha_{A}})\subseteq L_{E}(\beta_{A},r_{\beta_{A}})$ if and only if $\mathtt{ValC}(A)=\emptyset$. We define
$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}0$. Furthermore, for a yet unspecified $\mu\in\mathbb{N}$ that shall be defined later, let
$\beta_{A}:=x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}\#^{6}\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}\#^{6}\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}},$
with, for all $i\in\{1,\ldots,\mu\}$, $\hat{\beta}_{i}:=x_{i_{1}}\ \gamma_{i}\ x_{i_{2}}\ \delta_{i}\ x_{i_{3}}$ and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{1}},x_{1_{2}},x_{1_{3}},x_{1_{4}},x_{1_{5}},\ldots,x_{\mu},x_{\mu^{\prime}},x_{\mu_{1}},x_{\mu_{2}},x_{\mu_{3}},x_{\mu_{4}},x_{\mu_{5}}$ are distinct variables with $(x_{i},x_{i^{\prime}}),(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ and all
$\gamma_{i},\delta_{i},\eta_{i}\in X^{*}$ are terminal-free patterns. The patterns $\gamma_{i}$ and $\delta_{i}$ shall be defined later; for now, we only mention:
  1. 1.
    $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$,
  2. 2.
    var($\gamma_{i}\delta_{i}\eta_{i}$) $\cap$ var($\gamma_{j}\delta_{j}\eta_{j}$) = $\emptyset$ for all $i,j\in\{1,\ldots,\mu\}$
Thus, for every $i$, the elements of var($\gamma_{i}\delta_{i}\eta_{i}$) appear nowhere but in these three factors. Let $H$ be the set of all substitutions $\sigma:(\Sigma\ \cup\ \{x,y\})^{*}\rightarrow\Sigma^{*}$. We interpret each triple ($\gamma_{i},\delta_{i},\eta_{i}$) as a predicate $\pi_{i}:H\rightarrow\{0,1\}$ in such a way that $\sigma\in H$ satisfies $\pi_{i}$ if there exists a morphism $\tau:\text{var}(\gamma_{i}\delta_{i}\eta_{i})^{*}\rightarrow\Sigma^{*}$ with $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ - in the terminology of word equations (cf. Karhumäki et al. [DBLP:journals/jacm/KarhumakiMP00]), this means that $\sigma$ satisfies $\pi_{i}$ if and only if the system consisting of the three equations $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ has a solution $\tau$. Later, we shall see that $L_{E}(\alpha_{A},r_{\alpha_{A}})\setminus L_{E}(\beta_{A},r_{\beta_{A}})$ exactly contains those $\sigma(\alpha_{A})$ for which $\sigma$ does not satisfy any of $\pi_{1}$ to $\pi_{\mu}$, and choose these predicates to describe $\mathtt{InvalC}(A)$. The encoding of $\mathtt{InvalC}(A)$ shall be handled by $\pi_{4}$ to $\pi_{\mu}$, as each of these predicates describes a sufficient criterium for membership in $\mathtt{InvalC}(A)$. But at first we need a considerable amount of technical preparations. A substitution $\sigma$ is of good form if $\sigma(x)\in\{0,\#\}^{*}$, $\sigma(x)$ does not contain $^{3}$ as a factor, and $\sigma(y)\in 0^{*}$. Otherwise, $\sigma$ is of bad form. The predicates $\pi_{1}$ and $\pi_{2}$ handle all cases where $\sigma$ is of bad form and are defined through
$\gamma_{1}:=y_{1,1}\ \hat{z}_{1_{6}}\ \hat{z}_{1_{7}}\ \hat{z}_{1_{8}}\ y_{1,2},$ $\gamma_{2}:=y_{2},$
$\delta_{1}:=\hat{y}_{1},$ $\delta_{2}:=\hat{y}_{2,1}\ \hat{z}_{2_{6}}\ \hat{y}_{2,2}$
where $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2},\hat{z}_{1_{6}},\hat{z}_{1_{7}},\hat{z}_{1_{8}}$ and $\hat{z}_{2_{6}}$ are pairwise distinct variables with $(\hat{z}_{1_{1}},\hat{z}_{1_{6}}),(\hat{z}_{1_{1}},\hat{z}_{1_{7}}),(\hat{z}_{1_{1}},\hat{z}_{1_{8}}),(\hat{z}_{2_{1}},\hat{z}_{2_{6}})\in R$. Recall that
$\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ for all $i$. It is not very difficult to see that $\pi_{1}$ and $\pi_{2}$ characterize the morphisms that are of bad form:
Lemma 10
A substitution $\sigma\in H$ is of bad form if and only if $\sigma$ satisfies $\pi_{1}$ or $\pi_{2}$.
Proof
We begin with the only if direction. If $\sigma(x)=w_{1}\#^{3}w_{2}$ for some $w_{1},w_{2}\in\Sigma^{*}$, choose $\tau(y_{1,1}):=w_{1},\tau(y_{1,2}):=w_{2},\tau(\hat{z}_{1_{1}}):=\#,\tau(\hat{z}_{1_{2}}):=\#,\tau(\hat{z}_{1_{3}}):=\#,\tau(\hat{z}_{1_{4}}):=\#,\tau(\hat{z}_{1_{5}}):=\#,\tau(\hat{z}_{1_{6}}):=\#,\tau(\hat{z}_{1_{7}}):=\#,\tau(\hat{z}_{1_{8}}):=\#,\tau(\hat{y}_{1}):=\sigma(y),\tau(z_{1}):=0$ and $\tau(z_{1^{\prime}}):=0$. Then $\tau(y_{1})=\sigma(x),\tau(\delta_{1})=\sigma(y)$ and $\tau(\eta_{1})=u$; thus $\sigma$ satisfies $\pi_{1}$.
If $\sigma(y)=w_{1}\#w_{2}$, for some $w_{1},w_{2}\in\Sigma^{*}$, let $\tau(y_{2}):=\sigma(x)$, $\tau(\hat{y}_{2,1}):=w_{1}$, $\tau(\hat{y}_{2,2}):=w_{2}$, and $\tau(\hat{z}_{2_{1}}):=\#$, $\tau(\hat{z}_{2_{2}}):=\#$, $\tau(\hat{z}_{2_{3}}):=\#$, $\tau(\hat{z}_{2_{4}}):=\#$, $\tau(\hat{z}_{2_{5}}):=\#$, $\tau(\hat{z}_{2_{6}}):=\#$, and $\tau(z_{2}):=0,\tau(z_{2^{\prime}}):=0$. It is easy to see that $\sigma$ satisfies $\pi_{2}$.
For the if direction, if $\sigma$ satisfies $\pi_{1}$, then there exists a morphism $\tau$ with $\tau(\gamma_{1})=\sigma(x)$ and $\tau(\eta_{1})=0\#^{5}0$. Thus, $\tau(\hat{z}_{1_{1}})=\tau(\hat{z}_{1_{2}})=\tau(\hat{z}_{1_{3}})=\tau(\hat{z}_{1_{4}})=\tau(\hat{z}_{1_{5}})=\#$ and $\tau(z_{1})=\tau(z_{1^{\prime}})=0$ must hold. Consequently, $\tau(\hat{z}_{1_{6}})=\tau(\hat{z}_{1_{7}})=\tau(\hat{z}_{1_{8}})=\#$ and thus, $\sigma(x)$ contains $^{3}$, and $\sigma$ is of bad form.
Analogously, if $\sigma$ satisfies $\pi_{2}$, then $\sigma(y)$ contains the letter #\#, and $\sigma$ is of bad form.
This allows us to make the following observation, which serves as the central part of the construction and is independent from the exact shape of $\pi_{3}$ to $\pi_{\mu}$:
Lemma 11
For every substitution $\sigma\in H$, $\sigma(\alpha_{A})\in L_{E}(\beta_{A},r_{\beta_{A}})$ if and only if $\sigma$ satisfies one of the predicates $\pi_{1}$ to $\pi_{\mu}$.
Proof
We begin with the if direction. Assume $\sigma\in H$ satisfies some predicate $\pi_{i}$. Then there exists a morphism $\tau:\text{var}(\gamma_{i}\delta_{i}\eta_{i})\rightarrow\Sigma^{*}$ such that $\tau(\gamma_{i})=\sigma(x),\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$. We extend $\tau$ to a substitution $\tau^{\prime}$ defined by
  1. 1.
    $\tau^{\prime}(x):=\tau(x)$ for all $x\in\text{var}(\gamma_{i}\delta_{i}\eta_{i})$,
  2. 2.
    $\tau^{\prime}(x_{i}):=0\#^{3}0=v$ and $\tau^{\prime}(x_{i^{\prime}}):=\tau^{\prime}(x_{i_{1}}):=\tau^{\prime}(x_{i_{2}}):=\tau^{\prime}(x_{i_{3}}):=\tau^{\prime}(x_{i_{4}}):=\tau^{\prime}(x_{i_{5}}):=0\#^{3}0=v$,
  3. 3.
    $\tau^{\prime}(0):=0$ and $\tau^{\prime}(\#):=\#$,
  4. 4.
    $\tau^{\prime}(x):=\epsilon$ in all other cases.
By definition, none of the variables in var($\gamma_{i}\delta_{i}\eta_{i}$) appear outside of these factors. Thus, $\tau^{\prime}$ can always be defined in this way. We obtain
$\tau^{\prime}(\hat{\beta}_{i})$ $=\tau^{\prime}(x_{i_{1}}\ \gamma_{i}\ x_{i_{2}}\ \delta_{i}\ x_{i_{3}})$
$=v\ \tau(\gamma_{i})\ v\ \tau(\delta_{i})\ v$
$=v\ \sigma(x)\ v\ \sigma(y)\ v,$
$\tau^{\prime}(\ddot{\beta_{i}})\$$=\tau^{\prime}(x_{i_{4}}\ \eta_{i}\ x_{i_{5}})$
$=v\ \tau(\eta_{i})\ v$
$=v\ u\ v.$
As $\tau^{\prime}(\gamma_{j})=\tau^{\prime}(\delta_{j})=\tau^{\prime}(\eta_{j})=\tau^{\prime}(\hat{\beta}_{j})=\tau^{\prime}(\ddot{\beta_{j}})=\epsilon$ for all $j\neq i$, this leads to
$\tau^{\prime}(\beta_{A})$ $=\tau^{\prime}(x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}\#^{6}\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}\#^{6}\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}})$
$=\tau^{\prime}(x_{i}x_{i^{\prime}})\#^{6}\tau^{\prime}(\hat{\beta}_{i})\#^{6}\tau^{\prime}(\ddot{\beta_{i}})$
$=v\ v\ \#^{6}\ v\ \sigma(x)\ v\ \sigma(y)\ v\ \#^{6}\ v\ u\ v$
$=\sigma(\alpha_{A})$
This proves $\sigma(\alpha_{A})\in L_{E}(\beta_{A},r_{\beta_{A}})$.
For the other direction, assume that $\sigma(\alpha_{A})\in L_{E}(\beta_{A},r_{\beta_{A}})$. If $\sigma$ is of bad form, then by Lemma˜10, $\sigma$ satisfies $\pi_{1}$ of $\pi_{2}$. Thus, assume $\sigma(x)$ does not contain $^{3}$ as a factor, and $\sigma(y)\in 0^{*}$. Let $\tau$ be a substitution with $\tau(\beta_{A})=\sigma(\alpha_{A})$.
Now, as $\sigma$ is of good form, $\sigma(\alpha_{A})$ contains exactly two occurrences of $^{6}$, and these are non-overlapping. As $\sigma(\alpha_{A})=\tau(\beta_{A})$, the same holds for $\tau(\beta_{A})$. Thus, the equation $\sigma(\alpha_{A})=\tau(\beta_{A})$ can be decomposed into the system consisting of the following three equations:
$0\#^{3}0\ 0\#^{3}0=\tau(x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}),$ (4)
$0\#^{3}0\ \sigma(x)\ 0\#^{3}0\ \sigma(y)\ 0\#^{3}0=\tau(\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}),$ (5)
$0\#^{3}0\ u\ 0\#^{3}0=\tau(\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}})$ (6)
First, consider Equation˜4 and choose the smallest $i$ for which $\tau(x_{i})\neq\epsilon$. Then $\tau(x_{i})$ has to start with 0, and as
$\tau(x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}})=0\#^{3}0\ 0\#^{3}0$
and $(x_{i},x_{i^{\prime}})\in R$ for all $i\in\{1,\ldots,\mu\}$, it is easy to see that $\tau(x_{i})=0\#^{3}0=v=\tau(x_{i^{\prime}})$ and $\tau(x_{j})=\tau(x_{j^{\prime}})=\epsilon$ for all $j\neq i$ must hold. Furthermore, as $(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ for all $i\in\{1,\ldots,\mu\}$, we get $\tau(x_{i_{1}})=\tau(x_{i_{2}})=\tau(x_{i_{3}})=\tau(x_{i_{4}})=\tau(x_{i_{5}})=v$ and $\tau(x_{j_{1}})=\tau(x_{j_{2}})=\tau(x_{j_{3}})=\tau(x_{j_{4}})=\tau(x_{j_{5}})=\epsilon$ for all $j\neq i$.
Note that $u$ does not contain $0\#^{3}0$ as a factor, and does neither begin with $^{3}0$, nor end on $0\#^{3}$. But as $\tau(\ddot{\beta_{i}})$ begins with and ends on $0\#^{3}0$, we can use Equation˜6 to obtain $0\#^{3}0\ u\ 0\#^{3}0=\tau(\ddot{\beta_{i}})$ and $\tau(\ddot{\beta_{j}})=\epsilon$ for all $j\neq i$. As $\ddot{\beta_{i}}=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$ and $\tau(x_{i_{4}})=\tau(x_{i_{5}})=0\#^{3}0$, $\tau(\eta_{i})=u$ must hold.
As $\sigma$ is of good form, $\sigma(0\#^{3}0\ x\ 0\#^{3}0\ y\ 0\#^{3}0)$ contains exactly three occurrences of $^{3}$. But there are already three occurrences of $^{3}$ in $\tau(\hat{\beta}_{i})=0\#^{3}0\ \tau(\gamma_{i})\ 0\#^{3}0\ \tau(\delta_{i})\ 0\#^{3}0$. This, and Equation˜5, lead to $\tau(\hat{\beta}_{j})=\epsilon$ for all $j\neq i$ and, more importantly, $\tau(\gamma_{i})=\sigma(x)$ and $\tau(\delta_{i})=\sigma(y)$. Therefore, $\sigma$ satisfies the predicate $\pi_{i}$.
Thus, we can select predicates $\pi_{1}$ to $\pi_{\mu}$ in such a way that $L_{E}(\alpha_{A},r_{\alpha_{A}})\setminus L_{E}(\beta_{A},r_{\beta_{A}})=\emptyset$ if and only if $\mathtt{ValC}(A)=\emptyset$ by describing $\mathtt{InvalC}(A)$ through a disjunction of predicates on $H$. The proof of Lemma˜11 shows that if $\sigma(\alpha_{A})=\tau(\beta_{A})$ for substitutions $\sigma,\tau$, where $\sigma$ is of good form, there exists exactly one $i,3\leq i\leq\mu$, such that $\tau(x_{i})=0\#^{3}0$.
Due to technical reasons, we need a predicate $\pi_{3}$ that, if unsatisfied, sets a lower bound on the length of $\sigma(y)$, defined by
$\gamma_{3}\$$:=\gamma_{3,1}\hat{\gamma}_{3,1}\gamma_{3,2}\hat{\gamma}_{3,2}\gamma_{3,3}\hat{\gamma}_{3,3}\gamma_{3,4},$
$\delta_{3}\$$:=\hat{\gamma}_{3,1^{\prime}}\hat{\gamma}_{3,2^{\prime}}\hat{\gamma}_{3,3^{\prime}},$
where all of $\gamma_{3,1}$ to $\gamma_{3,4}$ and $\hat{\gamma}_{3,1}$ to $\hat{\gamma}_{3,3}$ and $\hat{\gamma}_{3,1^{\prime}}$ to $\hat{\gamma}_{3,3^{\prime}}$ are pairwise distinct variables with $(\hat{\gamma}_{3,1},\hat{\gamma}_{3,1^{\prime}}),(\hat{\gamma}_{3,2},\hat{\gamma}_{3,2^{\prime}}),(\hat{\gamma}_{3,3},\hat{\gamma}_{3,3^{\prime}})\in R$. Clearly, if some $\sigma\in H$ satisfies $\pi_{3}$, $\sigma(y)$ is a concatenation of three (possibly empty) reversed factors of $\sigma(x)$. Thus, if $\sigma$ satisfies none of $\pi_{1}$ to $\pi_{3}$, $\sigma(y)$ must be longer than the three longest non-overlapping sequences of 0s in $\sigma(x)$. This allows us to identify a class of predicates definable by a rather simple kind of expression, which we use to define $\pi_{4}$ to $\pi_{\mu}$ in a less technical way. The following definitions are adapted to work with our setting of relational patterns.
Let $X^{\prime}:=\{\hat{x}_{1,1},\hat{x}_{2,1},\hat{x}_{3,1},\hat{x}_{1,2},\hat{x}_{2,2},\hat{x}_{3,2},\ldots\}\subset X$ be an infinite set of new variables, let $G_{v_{R}}\subset H_{v_{R}}$ denote the set of all $\alpha$-valid substitutions of good form, and let $S$ be the set of all substitutions $\rho:(\Sigma\ \cup\ X^{\prime})^{*}\rightarrow\Sigma^{*}$ for which $\rho(0)=0,\rho(\#)=\#$ and $\rho(\hat{x}_{i,j})\in 0^{*}$ for $i\in\{1,2,3\}$ and $j\in\mathbb{N}$. For relational patterns $(p,v_{R})$ with $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ and $v_{R}:=\{(r_{rev},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\ |\ i\in\{1,2,3\};j,j^{\prime}\in\mathbb{N}\}$, we define $S(p,v_{R}):=\{\rho(p)\ |\ \rho\in S\cap H_{v_{R}}\}$.
Definition 2
A predicate $\pi:G_{v_{R}}\rightarrow\{0,1\}$ is called a simple predicate if there exists a pattern $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ and languages $L_{1},L_{2}\in\{\Sigma^{*},\{\epsilon\}\}$ such that $\sigma$ satisfies $\pi$ if and only if $\sigma(x)\in L_{1}\ S(p,v_{R})\ L_{2}$.
From a slightly different point of view, the elements of $X^{\prime}$ can be understood as numerical parameters describing (concatenational) powers of 0, with substitutions $\rho\in S$ acting as assignments. For example, if $\sigma\in G_{v_{R}}$ satisfies a simple predicate $\pi$ if and only if $\sigma(x)\in\Sigma^{*}S(\#\hat{x}_{1,1}\#\hat{x}_{2,1}\#\hat{x}_{1,2},v_{R})$, we can also write that $\sigma$ satisfies $\pi$ if and only if $\sigma(x)$ has a suffix of the form $0^{m}\#0^{n}0\#0^{m}$ (with $m,n\in\mathbb{N}_{0}$), which could also be written as $0^{m}\#0^{*}0\#0^{m}$, as $n$ occurs only once in this expression. Using $\pi_{3}$, our construction is able to express all simple predicates:
Lemma 12
For every simple predicate $\pi_{S}$ over variables from $X^{\prime}$, there exists a predicate $\pi$ defined by terminal-free patterns $\gamma,\delta,\eta$ such that for all substitutions $\sigma\in G_{v_{R}}$:
  1. 1.
    if $\sigma$ satisfies $\pi_{s}$, then $\sigma$ also satisfies $\pi$ or $\pi_{3}$,
  2. 2.
    if $\sigma$ satisfies $\pi$, then $\sigma$ also satisfies $\pi_{S}$.
Proof
We first consider the case of $L_{1}=L_{2}=\Sigma^{*}$. Assume that $\pi_{S}$ is a simple predicate, and $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ is a pattern such that $\sigma\in G_{v_{R}}$ satisfies $\pi_{S}$ if and only if $\sigma(x)\in L_{1}\ S(p,v_{R})\ L_{2}$. Then define $\gamma:=y_{1}p^{\prime}y_{2}$, where $p^{\prime}$ is obtained from $p$ by replacing all $k,k\in\mathbb{N}_{0}$ occurrences of 0 with new variables $\tilde{z}_{j},j\in[k]$ with $(\tilde{z}_{1},\tilde{z}_{j})\in R$ for all $j\in[k]$ and all $g,g\in\mathbb{N}_{0}$ occurrences of #\# with different variables $\ddot{z}_{j},j\in[g]$ with $(\ddot{z}_{1},\ddot{z}_{j})\in R$ for all $j\in[g]$, while leaving all present elements of $X^{\prime}$ unchanged. Furthermore, let $\delta:=\hat{x}_{1,j}\hat{x}_{2,j}\hat{x}_{3,j}\hat{y}$ with $j\in\mathbb{N}$ such that $\hat{x}_{1,j},\hat{x}_{2,j},\hat{x}_{3,j}$ are new variables and (in order to stay consistent with the $\eta_{i}$ appearing in $\beta_{A}$) $\eta:=\tilde{z}\ \ddot{z}_{g+1}\ \ddot{z}_{g+2}\ \ddot{z}_{g+3}\ \ddot{z}_{g+4}\ \ddot{z}_{g+5}\ \tilde{z}^{\prime}$ with $(\tilde{z}_{1},\tilde{z}),(\tilde{z}_{1},\tilde{z}^{\prime}),(\ddot{z}_{1},\ddot{z}_{g+1}),(\ddot{z}_{1},\ddot{z}_{g+2}),(\ddot{z}_{1},\ddot{z}_{g+3}),(\ddot{z}_{1},\ddot{z}_{g+4}),(\ddot{z}_{1},\ddot{z}_{g+5})\in R$.
Now, assume that $\sigma\in G_{v_{R}}$ satisfies $\pi_{S}$. Then there exist two words $w_{1},w_{2}\in\Sigma^{*}$ and a substitution $\rho\in S$ such that $\sigma(x)=w_{1}\rho(p)w_{2}$. If $\sigma(y)$ is not longer than any three non-overlapping factors of the form $0^{*}$ of $\sigma(x)$ combined, $\pi_{3}$ is satisfied. Otherwise, we can define $\tau$ by setting $\tau(y_{1}):=w_{1},\tau(y_{2}):=w_{2},\tau(\tilde{z}):=0,\tau(\tilde{z}^{\prime}):=0,\tau(\tilde{z}_{j}):=0$ for all $j\in[k],\tau(\ddot{z}_{j}):=\#$ for all $j\in[g+5],\tau(\hat{x}_{i,j}):=\rho(\hat{x}_{i,j})$ for all $i\in\{1,2,3\},j\in\mathbb{N}$ where $\hat{x}_{i,j}$ appears in $p$ and $\tau(\hat{x}_{i,j}):=\epsilon$ where $\hat{x}_{i,j}$ does not appear in $p$. Finally, let $\tau(\hat{x}_{i,j})=\tau(\hat{x}_{i,j^{\prime}})$ for all other $i\in\{1,2,3\},j\in\mathbb{N}$ for one arbitrary, specific $j^{\prime}\in\mathbb{N}$ such that $\tau(\hat{x}_{i,j^{\prime}})$ is defined and $\tau(\hat{y}):=0^{m}$, where
$m:=|\sigma(y)|-\sum_{\hat{x}\in\text{var}(p)}{|\sigma(\hat{x})|}$
($m>0$ must hold, as $\sigma$ does not satisfy $\pi_{3}$). Then $\tau(p^{\prime})=\rho(p)$, and
$\tau(\gamma)\$$=\ \tau(y_{1})\ \tau(p)\ \tau(y_{2})$
$=w_{1}\ \rho(p)\ w_{2}=\sigma(x),$
[…]
$=0^{|\sigma(y)|}=\sigma(y),$
[…]
$=0\#\#\#\#\#0=u.$
Therefore, $\sigma$ satisfies $\pi$, which concludes this direction.
[…]
$(\tilde{z}_{1},\tilde{z}),(\tilde{z}_{1},\tilde{z}^{\prime}),(\ddot{z}_{1},\ddot{z}_{g+1}),(\ddot{z}_{1},\ddot{z}_{g+2}),(\ddot{z}_{1},\ddot{z}_{g+3}),(\ddot{z}_{1},\ddot{z}_{g+4}),(\ddot{z}_{1},\ddot{z}_{g+5})\in R$ and u=0#####0u=0\#\#\#\#\#0, $\tau(\tilde{z})=\tau(\tilde{z}^{\prime})=0$ and $\tau(\ddot{z}_{j})=\#$ for all $j\in\{g+1,\ldots,g+5\}$ must hold. By definition $\tau(y_{1}),\tau(y_{2})\in\Sigma^{*}$. If we define $\rho(\hat{x}_{i,j}):=\tau(\hat{x}_{i,j})$ for all $i\in\{1,2,3\},j\in\mathbb{N}$ such that $\tau(\hat{x}_{i,j})$ is defined and $\rho(\hat{x}_{i,j}):=\rho(\hat{x}_{i,j^{\prime}})$ for all other $i\in\{1,2,3\},j\in\mathbb{N}$ for one arbitrary, specific $j^{\prime}\in\mathbb{N}$ such that $\rho(\hat{x}_{i,j^{\prime}})$ is defined, we see that $\sigma(x)\in L_{1}S(p,v_{R})L_{2}$ holds. Thus, $\sigma$ satisfies $\pi_{S}$ as well.
The other three cases for choices of $L_{1}$ and $L_{2}$ can be handled analogously by omitting $y_{1}$ or $y_{2}$ as needed. Note that this proof also works in the case $p=\epsilon$.
Roughly speaking, if $\sigma$ does not satisfy $\pi_{3}$, then $\sigma(y)$ (which is in $0^{*}$, due to $\sigma\in G_{v_{R}}$) is long enough to provide building blocks for simple predicates using variables from $X$.
Our next goal is a set of predicates that (if unsatisfied) forces $\sigma(x)$ into a basic shape common to all elements of $\mathtt{ValC}(A)$. We say that a word $w\in\{0,\#\}^{*}$ is of good structure if $w\in(\#\#0^{+}\#0^{+}\#0^{+})^{+}\#\#$. Otherwise, $w$ is of bad structure. Recall that due to the definition of enc, all elements of $\mathtt{ValC}(A)$ are of good structure, thus being of bad structure, we define predicates $\pi_{4}$ to $\pi_{13}$ through simple predicates as follows:
$\pi_{4}:\sigma(x)=\epsilon$, $\pi_{9}:\sigma(x)$ ends on 0,
$\pi_{5}:\sigma(x)=\#$, $\pi_{10}:\sigma(x)$ ends on 0#0\#,
$\pi_{6}:\sigma(x)=\#\#$, $\pi_{11}:\sigma(x)$ contains a factor $0^{*}\#\#$,
$\pi_{7}:\sigma(x)$ begins with 0, $\pi_{12}:\sigma(x)$ contains a factor $0^{*}\#0^{*}\#\#$,
$\pi_{8}:\sigma(x)$ begins with #0\#0, $\pi_{13}:\sigma(x)$ contains a factor $0^{*}\#0^{*}\#0^{*}\#0$.
Due to Lemma˜12, the predicates $\pi_{1}$ to $\pi_{13}$ do not strictly give rise to a characterization of substitutions with images that are of bad structure, as there are $\sigma\in G_{v_{R}}$ where $\sigma(x)$ is of good structure, but $\pi_{3}$ is satisfied due to $\sigma(y)$ being too short. But this problem can be avoided by choosing $\sigma(y)$ long enough to leave $\pi_{3}$ unsatisfied, and the following holds:
Lemma 13
A word $w\in\Sigma^{*}$ is of good structure if and only if there exists a substitution $\sigma\in H_{v_{R}}$ with $\sigma(x)=w$ such that $\sigma$ satisfies none of the predicates $\pi_{1}$ to $\pi_{13}$.
Proof
We begin with the if direction. Assume $\sigma\in H_{v_{R}}$ such that there is no $i\in\{1,\ldots,13\}$ for which $\sigma$ satisfies $\pi_{i}$. Due to Lemma˜10, $\sigma$ is of good form and, thus, $\sigma(y)\in 0^{*}$. As $\pi_{3}$ does not hold, $\sigma(y)$ is also longer than any three non-overlapping factors of $0^{*}$ of $\sigma(x)$. Thus, the structure of $\sigma(x)$ can be inferred by intersecting the complements of the simple predicates given in the definitions $\pi_{4}$ to $\pi_{13}$.
As $\sigma$ does not satisfy $\pi_{4}$, $\sigma(x)\neq\epsilon$. Due to $\pi_{7}$ and $\pi_{9}$, the first and the last letter of $\sigma(x)$ is #\#, and neither is #0\#0 a prefix, nor 0#0\# a suffix of $\sigma(x)$, as otherwise $\pi_{8}$ or $\pi_{10}$ would be satisfied. Therefore, $\sigma(x)$ has ##\#\# as prefix and suffix, but, as $\pi_{6}$ is not satisfied, $\sigma(x)\neq\#\#$. As $\sigma$ is of good form, $\sigma(x)$ does not contain ###\#\#\# as a factor, and
$\sigma(x)\in\#\#0^{+}\Sigma^{*}\#\#$
must hold. But as $\pi_{11}$ is not satisfied, it is possible to refine this observation to
$\sigma(x)\in\#\#0^{+}\#0^{+}\Sigma^{*}\#\#,$
which in turn leads to
$\sigma(x)\in\#\#0^{+}\#0^{+}\#0^{+}\Sigma^{*}\#\#$
by considering $\pi_{12}$ as well. Now, there are two possibilities. If
$\sigma(x)\in\#\#0^{+}\#0^{+}\#0^{+}\#\#,$
then $\sigma(x)$ is of good structure, but if
$\sigma(x)\in\#\#0^{+}\#0^{+}\#0^{+}\#\Sigma^{*}\#\#,$
then $\pi_{13}$ and $\sigma\in G_{v_{R}}$ lead to
$\sigma(x)\in\#\#0^{+}\#0^{+}\#0^{+}\#\#0^{+}\Sigma^{*}\#\#.$
In this case, we can continue referring subsequently to one of $\pi_{11}$ to $\pi_{13}$ together with $\sigma\in G_{v_{R}}$, and conclude
$\sigma(x)\in(\#\#0^{+}\#0^{+}\#0^{+})^{+}\#\#.$
Therefore, if $\sigma$ satisfies none of $\pi_{1}$ to $\pi_{13}$, then $\sigma(x)$ has to be of good structure.
Regarding the only if direction, assume some $w\in\Sigma^{*}$ is of good structure. Define $\sigma$ by $\sigma(x)=w$ and $\sigma(y)=0^{|w|+1}$. As $\sigma$ is of good form, Lemma˜10 demonstrates that $\sigma$ satisfies neither $\pi_{1}$ nor $\pi_{2}$; and as $\sigma(y)$ is longer than any word which results from concatenating any number of non-overlapping factors of the form $0^{*}$ of $w$, $\pi_{3}$ cannot be satisfied either. By looking at the cases used above to define $\pi_{4}$ to $\pi_{13}$, we see that none of these predicates can be satisfied.
For every $w$ of good structure, there exist uniquely determined
$n,i_{1},j_{1},k_{1},\ldots,i_{n},j_{n},k_{n}\in\mathbb{N}$ such that
$w=\#\#0^{i_{1}}\#0^{j_{1}}\#0^{k_{1}}\#\#\ldots\#\#0^{i_{n}}\#0^{j_{n}}\#0^{k_{n}}\#\#$. Thus, if $\sigma\in H_{v_{R}}$ does not satisfy any of $\pi_{1}$ to $\pi_{13}$, $\sigma(x)$ can be understood as an encoding of a sequence $T_{1},\ldots,T_{n}$ of triples $T_{i}\in(\mathbb{N})^{3}$, and for every sequence of that form, there is a $\sigma\in H_{v_{R}}$ such that $\sigma(x)$ encodes a sequence of triples of positive integers, and $\sigma$ does not satisfy any of $\pi_{1}$ to $\pi_{13}$.
[…]
All that remains is to choose an appropriate set of predicates that describe all cases where $C_{1}$ is not the initial configuration, $C_{n}$ is not an accepting configuration, or there are configurations $C_{i},C_{i+1}$ such that $C_{i}\vdash_{A}C_{i+1}$ does not hold (thus, the exact value of $\mu$ depends on the number of invalid transitions in $A$).
To ensure $C_{1}=(q_{0},0,0)$, we define a predicate
  1. 1.
    $\sigma(x)$ has a prefix of the form ##00\#\#00,
that is satisfied if $C_{1}$ has a state $q_{i}$ with $i>0$, and the two predicates
  1. 2.
    $\sigma(x)$ has a prefix of the form $0^{*}\#00$,
  2. 3.
    $\sigma(x)$ has a prefix of the form $0^{*}\#0^{*}\#00$,
to cover all cases where one of the counters is set to a value other than 0. Next, we handle the cases where the last state is not an accepting state. For every $i$ with $q_{i}\in Q\setminus F$, we define a predicate that is satisfied if
  1. 4.
    $\sigma(x)$ has a suffix of the form $00^{i}\#0^{*}\#\#$.
Thus, if $\sigma\in H_{v_{R}}$ satisfies none of the predicates defined up to this point, $\sigma(x)=\text{enc}(C_{1},\ldots,C_{n})$ for some computation $C_{1},\ldots,C_{n}$ with $C_{1}=(q_{0},0,0)$ and $C_{n}\in F\times\mathbb{N}_{0}\times\mathbb{N}_{0}$, there is a $\sigma\in H_{v_{R}}$ with $\sigma(x)=\text{enc}(C_{1},\ldots,C_{n})$, and $\sigma$ satisfies none of these predicates.
All that remains is to define a set of predicates that describe those $C_{i},C_{i+1}$ for which $C_{i}\vdash_{A}C_{i+1}$ does not hold. To simplify this task, we define the following four predicates that are satisfied if one of the counters is changed by more than 1:
  1. 5.
[…]
  2. 6.
[…]
  3. 7.
[…]
  4. 8.
[…]
Here, the first two predicates cover incrementing (or decrementing) the first counter by 2 or more; the other two do the same for the second counter. Then, for all $i,j\in\{1,\ldots,s\}$, all $c_{1},c_{2}\in\{0,1\}$ and all $r_{1},r_{2}\in\{-1,0,+1\}$ for which $(q_{j},r_{1},r_{2})\notin\delta(q_{i},c_{1},c_{2})$, we define a predicate that is satisfied if $\sigma(x)$ contains such a transition. We demonstrate this only for the exemplary case $c_{1}=0,c_{2}=1,r_{1}=+1,r_{2}=0$ without naming $i$ or $j$ explicitly. The predicate covering non-existing transitions of this form is
  1. 9.
    $\sigma(x)$ contains a factor of the form $0^{i+1}\#0\#000^{m}\#\#0^{j+1}\#00\#000^{m}\#\#$ for some $m\in\mathbb{N}_{0}$.
All other predicates for illegal transitions are defined analogously. Note that we can safely assume that none of the counters is changed by more than 1, as these errors are covered by the predicates we defined under points 5-8. The number of predicates required for these points and point 9 determine the exact value of $\mu$.
Now, if there is a substitution $\sigma$ that does not satisfy any of $\pi_{1}$ to $\pi_{\mu}$, then $\sigma(x)=\text{enc}(C_{1},\ldots,C_{n})$ for a computation $C_{1},\ldots,C_{n}$, where $C_{1}$ is the initial and $C_{n}$ a final configuration, and for all $i\in\{1,\ldots,n-1\},C_{i}\vdash_{A}C_{i+1}$. Thus, if $\sigma(\alpha_{A})\notin L_{E}(\beta_{A},r_{\beta_{A}})$, then $\sigma(x)\in\mathtt{ValC}(A)$, which means that $A$ has an accepting computation.
Conversely, if there is some accepting computation $C_{1},\ldots,C_{n}$ of $A$, we can define $\sigma$ through $\sigma(x):=\text{enc}(C_{1},\ldots,C_{n})$, and choose $\sigma(y)$ to be an appropriately long sequence from $0^{*}$. Then $\sigma$ does not satisfy any of the predicates $\pi_{1}$ to $\pi_{\mu}$ defined above, thus $\sigma(\alpha_{A})\notin L_{E}(\beta_{A},r_{\beta_{A}})$, and $L_{E}(\alpha_{A},r_{\alpha_{A}})\nsubseteq L_{E}(\beta_{A},r_{\beta_{A}})$.
We conclude that $A$ has an accepting computation iff $L_{E}(\alpha_{A},r_{\alpha_{A}})$ is not a subset of $L_{E}(\beta_{A},r_{\beta_{A}})$. Therefore, any algorithm deciding the inclusion problem of Theorem˜4.1 can be used to decide whether a nondeterministic 2-counter automaton without input has an accepting computation. As this problem is known to be undecidable, the inclusion problem of Theorem˜4.1 is also undecidable.
This proof can be extended to larger (finite) alphabets. Assume that $\Sigma=\{0,\#,a_{1},\ldots,a_{n}\}$ for some $n\geq 1$. We extend $H$ to the set of all substitutions $\sigma:(\Sigma\cup\{x,y\})^{*}\rightarrow\Sigma^{*}$, but do not extend the definition of substitutions of good form to our new and larger alphabet. Thus, $\sigma\in H$ is of good form if $\sigma(x)\in\{0,\#\}^{*},\sigma(y)\in 0^{*}$ and $\sigma(x)$ does not contain $^{3}$ as a factor. In addition to the predicates $\pi_{1}$ to $\pi_{\mu}$, for each new letter $a_{i}$, we define a predicate $\pi_{\mu+2i-1}$ which implies that $\sigma(x)$ contains an occurrence of $a_{i}$, and a predicate $\pi_{\mu+2i}$ which implies that $\sigma(y)$ contains an occurrence of $a_{i}$. To this end, we define
$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}a_{1}^{5}...a_{n}^{5}0$ (instead of $u=0\#^{5}0$), add the new predicates $\pi_{\mu+1}$ to $\pi_{\mu+2n}$ (which we still leave unspecified for a moment) to $\beta_{A}$ and use
$\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ \ddot{z}_{i,1_{1}}\ \ddot{z}_{i,1_{2}}\ \ldots\ \ddot{z}_{i,n_{1}}\ \ddot{z}_{i,n_{2}}\ z_{i^{\prime}}$
instead of $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$, where all $z_{i},z_{i^{\prime}},\hat{z}_{i_{j}},\ddot{z}_{i,k_{j}}$ are pairwise different variables with
$(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}}),(\ddot{z}_{i,k_{1}},\ddot{z}_{i,k_{2}})\in R.$
Referring to the new shape of $u$, we can make the following observation:
Lemma 14
Let $n\geq 2$, $\{x_{1},x_{1}^{\prime},x_{2_{1}},x_{2_{2}},x_{2_{3}},x_{2_{4}},x_{2_{5}},x_{3_{1}},x_{3,2},\ldots,x_{n_{1}},x_{n_{2}}\}\subset X$ and $\{a_{1},\ldots,a_{n}\}\subseteq\Sigma$. If
$\alpha=x_{1}\ x_{2_{1}}\ x_{2_{2}}\ x_{2_{3}}\ x_{2_{4}}\ x_{2_{5}}\ x_{3_{1}}\ x_{3_{2}}\ldots x_{n_{1}}\ x_{n_{2}}\ x_{1}^{\prime}$
with $(x_{1},x_{1}^{\prime}),(x_{2_{1}},x_{2_{2}}),(x_{2_{1}},x_{2_{3}}),(x_{2_{1}},x_{2_{4}}),(x_{2_{1}},x_{2_{5}}),(x_{i_{1}},x_{i_{2}})\in R$ for all $i\in\{3,\ldots,n\}$ and there is a morphism $\sigma:X^{*}\rightarrow\Sigma^{*}$ with
$\sigma(\alpha)=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$, then $\sigma(x_{1})=\sigma(x_{1}^{\prime})=a_{1}$ and $\sigma(x_{i_{j}})=a_{i}$ for each $i\in\{1,\ldots,n\}$.
Proof
Assume
$\sigma(x_{1}\ x_{2_{1}}\ x_{2_{2}}\ x_{2_{3}}\ x_{2_{4}}\ x_{2_{5}}\ x_{3_{1}}\ x_{3_{2}}\ldots x_{n_{1}}\ x_{n_{2}}\ x_{1}^{\prime})=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$
. If $\sigma(x_{1})=\varepsilon$, then
$\sigma(x_{2_{1}}\ x_{2_{2}}\ x_{2_{3}}\ x_{2_{4}}\ x_{2_{5}}\ x_{3_{1}}\ x_{3_{2}}\ldots x_{n_{1}}\ x_{n_{2}})=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$
leads to an immediate contradiction. But $\sigma(x_{1})\neq\varepsilon$ implies $\sigma(x_{1})=a_{1}$. Therefore,
$\sigma(x_{2_{1}}\ x_{2_{2}}\ x_{2_{3}}\ x_{2_{4}}\ x_{2_{5}}\ x_{3_{1}}\ x_{3_{2}}\ldots x_{n_{1}}\ x_{n_{2}})=(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}$
must hold. Now, $\sigma(x_{2_{j}})=a_{2}$ must hold for every $j\in\{1,2,3,4,5\}$ since
$(x_{2_{1}},x_{2_{2}}),(x_{2_{1}},x_{2_{3}}),(x_{2_{1}},x_{2_{4}}),(x_{2_{1}},x_{2_{5}}),(x_{i_{1}},x_{i_{2}})\in R$ for all $i\in\{3,\ldots,n\}$. Thus, we also get that $\sigma(x_{i_{1}})=\sigma(x_{i_{2}})=a_{i}$ for every $i\in\{3,\ldots,n\}$.
Lemma˜14 allows $\pi_{\mu+1}$ to $\pi_{\mu+2n}$ to be analogously constructed to $\pi_{2}$. To this end, we define
$\gamma_{\mu+2i-1}:=y_{\mu+2i-1,1}\ \ddot{z}_{\mu+2i-1,i}^{\prime}\ y_{\mu+2i-1,2},$ $\gamma_{\mu+2i}:=y_{\mu+2i},$
$\delta_{\mu+2i-1}:=\hat{y}_{\mu+2i-1},$ $\delta_{\mu+2i}:=\hat{y}_{\mu+2i,1}\ \ddot{z}_{\mu+2i,i}^{\prime}\ \hat{y}_{\mu+2i,2}$
for each $i\in[n]$. Again, all $y_{j,k},\hat{y}_{j,k},\ddot{z}_{j,k}^{\prime}$ are pairwise different variables with $(\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i-1,i}^{\prime}),(\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i,i}^{\prime})\in R$. Now Lemma˜10 applies (mutatis mutandis) as for binary alphabets, and since all substitutions of good form behave for $\Sigma$ as for the binary alphabet, we can use the very same predicates and the same reasoning as before to prove undecidability of the inclusion problem of Theorem˜4.1.
This concludes the proof of Theorem˜4.1.
Appendix 0.C Additional Content for Equivalence Problem
Proof of Lemma˜4:
Proof
Assume $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$. First of all, we know $|\alpha|=|\beta|$ as otherwise, if w.l.o.g. $|\alpha|>|\beta|$, then no shortest word $w\in L_{NE}(\beta,[r_{\beta}])$ obtained by substitutions $h$ where variables are replaced with letters can be obtained from $(\alpha,r_{\alpha})$. So, from now on, assume $|\alpha|=|\beta|$. Suppose $\alpha\neq\beta$. Then there exists a first position $i\in[|\alpha|]$ such that $\alpha[i]\neq\beta[i]$. First, if $\alpha[i],\beta[i]\in\Sigma$ are both terminal letters, then using a substitution $h$ that replaces all variables with letters in $\alpha$ and $\beta$ always results in words $h(\alpha)$ and $h(\beta)$ for which $h(\alpha)[i]\neq h(\beta)[i]$, a contradiction. So, assume w.l.o.g. that $\alpha[i]\in\Sigma$ is a terminal letter and $\beta[i]\in X$ is a variable. Then, again, any substitution $h\in H_{r_{\beta}}$ that replaces all variables in $\beta$ with letters and sets $h(\beta[i])$ to a letter which is not $\alpha[i]$ produces a word that cannot be in $L_{NE}(\alpha,[r_{\alpha}])$, another contradiction to $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$. As we assume all variables to be introduced in an increasing manner, $\alpha=\beta$. ∎
Proof of Lemma˜5:
Proof
($\Leftarrow$): Assume $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$. Then $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$. So, we trivially get $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$.
($\Rightarrow$) Assume $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$. By Lemma 4 we have $\alpha=\beta$. The relational constraints $r_{\alpha}]$ and $r_{\beta}]$ imply blocks of equivalently substituted variables under $\tilde{r}$. Suppose $r_{\alpha}]\neq[r_{\beta}]$. Let $\alpha=\alpha_{p}x_{i}\alpha_{s}=\beta$ for some $\alpha_{p},\alpha_{s}\in\mathtt{RelPat}_{\Sigma,R}$ such that $x_{i}$ is the first variable for which there is a difference in the relations, i.e., we get that w.l.o.g. $x_{i}\in\operatorname{var}(\alpha)$ is related to some other variable $x_{j}\in\operatorname{var}(\alpha)$ in $r_{\alpha}]$ but not in $r_{\beta}]$ (note that $\operatorname{var}(\alpha)=\operatorname{var}(\beta)$). So, there exists $(\tilde{r},x_{i},x_{j})\in[r_{\alpha}]$ such that $(\tilde{r},x_{i},x_{j})\notin[r_{\beta}]$. Select $h\in H_{r_{\beta}}$ such that $h$ substitutes all variables in $\beta$ to some letter, i.e., $|h(x)|=1$ for all $x\in\operatorname{var}(\beta)$. In particular, let $h(x_{i})=h(x_{k})=\mathtt{a}$, for some $\mathtt{a}\in\Sigma$ and all $(\tilde{r},x_{i},x_{k})\in[r_{\beta}]$. Finally, set $h(x_{j})=h(x_{k^{\prime}})=\mathtt{b}\neq\mathtt{a}=h(x_{i})$, for all other variables $x_{k^{\prime}}\in\operatorname{var}(\beta)$ which are not already set to $h(x_{i})=\mathtt{a}$ and some letter $\mathtt{b}\in\Sigma$ with $\mathtt{a}\neq\mathtt{b}$. Then, we have that $h(\beta)\in L_{NE}(\beta,[r_{\beta}])$ but $h(\beta)\notin L_{NE}(\alpha,[r_{\alpha}])$ as there cannot exist any substitution $h^{\prime}\in H_{r_{\alpha}}$ with $h^{\prime}(x_{i})\neq h^{\prime}(x_{j})$ as $(\tilde{r},x_{i},x_{j})\in[r_{\alpha}]$ and by the assumption that $\tilde{r}$ is an equivalence relation. This is a contradiction to $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$. So, we must have $r_{\alpha}]=[r_{\beta}]$. ∎
Proof of Theorem˜5.1:
Proof
($\Rightarrow$): Assume $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$. As $\tilde{r}$ is an equivalence relation, we get by Lemma 3 that $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$ and $L_{NE}(\beta,r_{\beta})=L_{NE}(\beta,[r_{\beta}])$. Hence, $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$, and we get by Lemma 5 that $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$. Thus, we have $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
($\Leftarrow$): Assume $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$. Then, by Lemma 5 we have $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$ and by Lemma 3 we get that $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ using the same arguments from the other direction. ∎
Proof of Corollary˜3:
Proof
First, we notice that all relations in $\{r_{=},r_{ab},r_{con^{+}}\}$ are equivalence relations that are antisymmetric on elements of length $1$. Now, we normalize $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ such that all variables are introduced with an index in increasing manner, starting with $x_{1}$ in both relational patterns. Then, we check in linear time whether $\alpha=\beta$. If that is the case, we check whether $r_{\alpha}]=[r_{\beta}]$ by constructing the sets representing reflexive, symmetric, and transitive closure of related variables. This can be done in linear time in the lengths of the patterns by scanning them and filling up the resulting sets. The total number of elements in these sets will not exceed the total number of variables in the patterns. If these sets are equal for both relational patterns, we return true, applying Theorem˜5.1. In any other case, we return false. ∎

### v2

@@
 In an extended notion, called relational patterns and relational pattern languages, variables may be related by arbitrary other relations.
-We extend the ongoing investigation of the main decision problems for patterns (namely, the membership problem, the inclusion problem, and the equivalence problem) to relational pattern languages under a wide range of individual relations.
-It is shown show that - even for many much simpler or less restrictive relations - the complexity and (un)decidability characteristics of these problems do not change compared to the classical case where variables are related only by equality.
+We extend the ongoing investigation of the main decision problems for patterns (namely, the equivalence problem, the inclusion problem, and the membership problem) to relational pattern languages under a wide range of relevant individual relations, providing a comprehensive foundation in all three research directions.
 1 Introduction
@@
 The language of a pattern is then just the set of all words that can be obtained through the application of arbitrary substitutions.
-Originally, pattern languages as defined and introduced by Angluin [Angluin1980] only considered words that are obtained by what is now called non-erasing substitutions, i.e., substitutions where each variable is substituted by a word of at least length one.
+Originally, pattern languages as defined and introduced by Angluin [3] only considered words that are obtained by what is now called non-erasing substitutions, i.e., substitutions where each variable is substituted by a word of at least length one.
 Hence, these languages are now also called non-erasing (NE) pattern languages.
-Shinohara [Shinohara1983] extended this notion by allowing for the empty word to be used to substitute a variable, introducing so called extended/erasing (E) pattern languages.
+Shinohara [42] extended this notion by allowing for the empty word to be used to substitute a variable, introducing so-called extended/erasing (E) pattern languages.
 For example, consider the pattern $\alpha=x_{1}\mathtt{a}\mathtt{a}x_{1}\mathtt{b}x_{2}$.
@@
 As patterns form a natural and compact device for the definition of formal languages and due to their practical and simple definition, they occur in numerous areas in computer science and discrete mathematics.
-Among others, these include for example the areas of unavoidable patterns in words [Jiang1994, lothaire1997], word equations [lothaire1997], algorithmic learning theory [Angluin1980, FERNAU201844, Shinohara1995], database theory [FreydenbergerP21, SchmidSchweikardtPODS2022], or the theory of extended regular expressions with back references[FREYDENBERGER20191].
-Furthermore, many practical areas such as machine learning, database systems, or bio-informatics deal with the question of identifying or describing patterns in (sets of) strings [Mousawi2024], for example in the analysis of protein data [Arikawa1993], in pattern matching [10.1007/978-3-642-03784-9_29], or the design of algorithms for program synthesis [10.1145/4472.4476].
+Among others, these include for example the areas of unavoidable patterns in words [22, 26], word equations [26], algorithmic learning theory [3, 9, 43], database theory [14, 41], or the theory of extended regular expressions with back references[15].
+Furthermore, many practical areas such as machine learning, database systems, or bio-informatics deal with the question of identifying or describing patterns in (sets of) strings [30], for example in the analysis of protein data [4], in pattern matching [6], or the design of algorithms for program synthesis [32].
 Three main decision problems emerge when it comes to the investigation of patterns and pattern languages.
-These are the general membership problem and its variations [gawrychowski_et_al:LIPIcs.MFCS.2021.48, Manea2022, Fleischmann2023], the inclusion problem, and the equivalence problem, all of which are considered in the non-erasing (NE) as well as erasing (E) cases.
-The membership problem determines whether a word belongs to a pattern’s language.
-It has been shown to be generally NP-complete in the non-erasing as well as erasing cases [Angluin1980, Jiang1994].
+These are the general equivalence problem, the inclusion problem, and the membership problem and its variations [16, 17, 12], all of which are considered in the non-erasing (NE) as well as erasing (E) cases.
+The equivalence problem determines whether the languages of two given patterns are equal.
+In Angluin’s original work [3], it has been shown to be trivially decidable for all NE pattern languages.
+For erasing pattern languages, however, the general decidability of this problem has been the most significant open problem in the field of decision problems for patterns [23, 38, 37, 35, 39].
 The inclusion problem checks whether the language of a pattern is included as a subset in another pattern’s language.
 It has been shown to be generally undecidable by Jiang et al.
-[Jiang1995].
-Their result has been improved by Freydenberger and Reidenbach in [Freydenberger2010] as well as by Bremer and Freydenberger in [BREMER201215], where it has been shown to be undecidable for all alphabets $\Sigma$ of fixed size $|\Sigma|\geq 2$, for erasing and non-erasing pattern languages.
-The equivalence problem determines whether the languages of two given patterns are equal.
-In Angluin’s original work [Angluin1980], it has been shown to be trivially decidable for all NE pattern languages.
-For erasing pattern languages, however, the general decidability of this problem has been the most significant open problem in the field of decision problems for patterns [Jiang1995, Reidenbach2004-1, Reidenbach2004-2, Ohlebusch1996, Reidenbach2007].
-In the special case of terminal-free patterns (containing only variables), the membership problem remains unchanged in both erasing and non-erasing cases, as does the non-erasing equivalence problem.
-However, inclusion for terminal-free E-pattern languages is NP-complete, as it reduces to checking for a morphism between patterns [Jiang1995, DBLP:journals/ipl/EhrenfeuchtR79a], while equivalence is decidable.
-In contrast, Saarela [saarela:LIPIcs.ICALP.2020.140] showed that inclusion for terminal-free NE pattern languages is undecidable over a binary alphabet.
-Various extensions to the notion of patterns and pattern languages have been introduced over time either to get closer to an answer for the remaining open problems or to obtain additional expressibility that is usable in some practical context.
-Some examples are the bounded scope coincidence degree, patterns with bounded treewidth, $k$-local patterns, or strongly-nested patterns (see [Day2018] and references therein).
-Koshiba [Koshiba1995] introduced so called typed patterns where the substitutions of variables can be restricted by arbitrary recursive languages, so called types.
-This has been recently extended by Geilke and Zilles in [Geilke2011] to the notion of relational patterns and relational pattern languages.
-Here, other relations than equality between variables may be used to restrict the valid substitution of patterns.
+[23].
+Their result has been improved by Freydenberger and Reidenbach in [13] as well as by Bremer and Freydenberger in [5], where it has been shown to be undecidable for all alphabets $\Sigma$ of fixed size $|\Sigma|\geq 2$, for erasing and non-erasing pattern languages.
+The membership problem determines whether a word belongs to a pattern’s language.
+It has been shown to be generally NP-complete in the non-erasing as well as erasing cases [3, 22] and has been investigated extensively in various communities [2, 36, 11, 10, 20, 28, 27].
+In the special case of terminal-free patterns (containing only variables), the membership problem remains hard in both erasing and non-erasing cases, as does the non-erasing equivalence problem.
+However, inclusion for terminal-free E-pattern languages is NP-complete, as it reduces to checking for a morphism between patterns [23, 8], resulting also in decidability of equivalence.
+In contrast, Saarela [40] showed that inclusion for terminal-free NE pattern languages is undecidable over a binary alphabet.
+Various extensions to the notion of pattern languages have been introduced over time either to get closer to an answer for the remaining open problems or to obtain additional expressibility that is usable in some practical context.
+Some examples include the bounded scope coincidence degree, patterns with bounded treewidth, $k$-local patterns, or strongly-nested patterns (see [7] and references therein).
+Koshiba [25] introduced so-called typed patterns where the substitutions of variables can be restricted by arbitrary recursive languages, so-called types.
+This has been recently extended by Geilke and Zilles in [18] to the notion of relational patterns and relational pattern languages.
+Here, relations other than equality may be used to restrict the valid substitution of patterns.
 This model enhances expressiveness, includes all previous models, but also includes arbitrary other constructions that involve relations between variables in a pattern.
 Consider, for example, the pattern $\beta=x_{1}\mathtt{c}\mathtt{c}x_{2}$.
-Now let’s say that the variables $x_{1}$ and $x_{2}$ should be the substituted by the reversal of each other, i.e., $(h(x_{1}),h(x_{2}))\in r_{rev}$, where $h$ is a substitution and $r_{rev}$ is the reversal relation defined by $r_{rev}:=\{\ (w,w^{R})\mid w\in\Sigma^{*}\ \}$ (where $w^{R}$ is defined to be the reversal of a word $w$).
-Then, e.g., a substitution $h$ with $h(x_{1})=\mathtt{a}\mathtt{b}$ and $h(x_{2})=\mathtt{b}\mathtt{a}$ is said to be valid regarding the pattern $\beta$, we get $h(\beta)=\mathtt{a}\mathtt{b}\mathtt{c}\mathtt{c}\mathtt{b}\mathtt{a}$, and we say that $h(\beta)$ is in the relational pattern language of $\beta$ under the relational constraints containing the pair that relates $x_{1}$ and $x_{2}$ by $r_{rev}$.
-If we considered a relation $h^{\prime}$, however, in which $h^{\prime}(x_{1})=\mathtt{a}\mathtt{b}$ and also $h^{\prime}(x_{2})=\mathtt{a}\mathtt{b}$, then clearly $h^{\prime}(x_{1})\neq h^{\prime}(x_{2})^{R}$ and therefore $h^{\prime}$ is not a valid substitution in this case, resulting in $h^{\prime}(\beta)$ not being in that relational pattern language of $\beta$.
-Geilke and Zilles [Geilke2011] show some properties of relational pattern languages in general.
+Assume $x_{1}$ and $x_{2}$ should be substituted by the reversal of each other, i.e., $(h(x_{1}),h(x_{2}))\in r_{rev}$, where $h$ is a substitution and $r_{rev}$ is the reversal relation defined by $r_{rev}:=\{\ (w,w^{R})\mid w\in\Sigma^{*}\ \}$ (where $w^{R}$ is the reversal of a word $w$).
+Then, a substitution $h$ with $h(x_{1})=\mathtt{a}\mathtt{b}$ and $h(x_{2})=\mathtt{b}\mathtt{a}$ is called valid regarding the pattern $\beta$, we get $h(\beta)=\mathtt{a}\mathtt{b}\mathtt{c}\mathtt{c}\mathtt{b}\mathtt{a}$, and we say that $h(\beta)$ is in the relational pattern language of $\beta$ under the relational constraints that relate $x_{1}$ and $x_{2}$ by $r_{rev}$.
+If we considered a substitution $h^{\prime}$, however, in which $h^{\prime}(x_{1})=\mathtt{a}\mathtt{b}$ and also $h^{\prime}(x_{2})=\mathtt{a}\mathtt{b}$, then clearly $h^{\prime}(x_{1})\neq h^{\prime}(x_{2})^{R}$ and therefore $h^{\prime}$ is not a valid substitution in this case, resulting in $h^{\prime}(\beta)$ not being in that relational pattern language of $\beta$.
+Geilke and Zilles [18] show some properties of relational pattern languages in general.
 Holte et al.
-[pmlr-v167-holte22a] investigated relational pattern languages under the equal length or reversal relations, respectively.
-They show that inclusion, by that also equivalence, under equal length is decidable in polynomial time for ternary alphabets.
-In a very recent paper, Mousawi and Zilles [Mousawi2024] extended this from the perspective of learning theory, introducing the notion of positive characteristic sets.
+[19] investigated relational pattern languages under the equal length or reversal relations, respectively.
+They show that inclusion, and thus also equivalence, under equal length is decidable in polynomial time for ternary alphabets.
+Recently, Mousawi and Zilles [30] extended this from the perspective of learning theory, introducing the notion of positive characteristic sets.
 Their results carry over and it is shown that inclusion, therefore also equivalence, in the erasing and non-erasing cases under equal length is decidable in polynomial time for binary alphabets.
 Also, for the reversal relation, their result implies that in the non-erasing case, equivalence is decidable.
-Additionally, also in two recent papers, in the cases of two special kinds of relations/constraints, namely regular constraints [Nowotka2024] or length constraints[Nowotka2025], Nowotka and Wiedenhöft have shown, among other results, that the still open equivalence problem for erasing pattern languages actually becomes undecidable if these constraints are allowed in addition to equality between variables.
+Additionally, in the cases of two special kinds of relations/constraints, namely regular constraints [33] or length constraints[34], it was recently shown, among other results, that the still open equivalence problem for erasing pattern languages actually becomes undecidable if these constraints are allowed in addition to equality between variables.
 This sets a first upper bound of constraints necessary to obtain undecidability for that problem.
-In comparison to many other relations, however, these constraints appear to be very strong, hence leaving the question of what happens in cases of weaker relations.
-In total, we see that the field of relational pattern languages remains largely unexplored, leaving room to study various other or weaker relations restricting variable substitutions.
-This paper considers equality ($r_{=}$), equal length ($r_{|w|}$), subsequence ($r_{ssq}$), abelian equivalence ($r_{ab}$), alphabet permutation ($r_{perm}$), reversal ($r_{rev}$), (nonempty) commutation ($r_{com^{+}}$, $r_{com^{*}}$), and the star relation ($r_{*}$).
-We study the membership, inclusion, and equivalence problems under these relations, providing a comprehensive overview of many cases (Table 3) and showing that hardness and (un)decidability persist even for weaker relations than equality.
-In particular, it can be shown that the membership problem is NP-complete for all considered relations in all cases.
-Regarding the inclusion problem, it is shown that undecidability persists even for the relations $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ in the erasing and non-erasing cases.
-Finally, in the case of the equivalence problem, we show decidability in polynomial time in the non-erasing case for a class of relations with specific properties, containing $r_{ab}$ and $r_{com^{+}}$.
+In comparison to other relations, however, these constraints appear to be very strong, leaving the question of what happens in cases of weaker relations.
+Hence, the field of relational pattern languages remains largely unexplored, leaving room to study various other relations.
+This paper considers equality ($r_{=}$), equal length ($r_{|w|}$), subsequence ($r_{ssq}$), abelian equivalence ($r_{ab}$), alphabet permutation ($r_{perm}$), reversal ($r_{rev}$), (nonempty) commutation ($r_{com^{+}}$, $r_{com^{*}}$), the star relation ($r_{*}$), and the common (non)emptiness relation ($r_{\varepsilon_{=}}$).
+We study the equivalence, inclusion, and membership problems under these relations, providing a comprehensive overview of many cases (Table 1), with a primary focus on general patterns that may contain terminals.
+We show that in some cases interesting decidability results emerge.
+In many other cases, however, it is shown that hardness and (un)decidability persist even for structurally weaker relations than equality.
+In particular, it can be shown that, first, the equivalence problem in the non-erasing cases is decidable in polynomial time for $r_{\varepsilon_{=}}$, $r_{ab}$ and $r_{com^{+}}$ and that equivalence of pattern languages over the relations $_{ssq}, rr$$_{rev}$, and $r_{*}$ emit strong structural restrictions.
+For erasing cases, i.e.
+the most important open question for $r_{=}$, that problem turns out to be decidable for $r_{\varepsilon_{=}}$.
+On the other hand, it is shown that this problem must be at least as hard as for $r_{=}$ for the relations $r_{rev}$, $r_{ssq}$, and $r_{*}$.
+Regarding the inclusion problem, we obtain decidability in the case of $r_{\varepsilon_{=}}$ while also observing that undecidability persists even for the relations $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ in the erasing and non-erasing cases.
+Finally, it can be shown that the membership problem is NP-complete for all considered relations in all cases, except for $r_{\varepsilon_{=}}$ in the non-erasing case.
 Section 2 covers all the necessary notation and models needed to follow the results of the paper.
-In Section 3, we present the results for the membership problem.
-This is followed in Section 4 with an overview of the results for the inclusion problem.
-After that, in Section 5, we briefly discuss the results regarding the equivalence problem.
-Finally, in Section 6, we summarize the paper and provide a brief outlook with regards to open questions.
-Due to space constraints, some proofs were moved into the appendix.
-The corresponding statements are marked with (✂).
+In Section 3, we discuss the equivalence problem.
+This is followed in Section 4 with a discussion of the inclusion problem.
+After that, in Section 5, we briefly discuss the results for the membership problem.
+Finally, in Section 6, we provide a final discussion of the established results and worthwhile research direction that could follow from the open cases.
+Due to space constraints, some proofs were moved into the appendix, marked with (✂) and a reference to the respective subsection.
+$r_{=}$ $r_{|w|}$ $r_{ssq}$ $r_{ab}$ $_{perm} rr$$_{rev}$ $r_{com^{*}}$ $_{com^{+}} rr_{*}$ $r_{\varepsilon_{=}}$
+ E ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC NPC
+ E ($\subseteq$)  UD  P UD UD ? UD ? ? UD PS
+ E (==) ?  P ? ? ? ? ? ? ? PS
+ NE ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC P
+ NE ($\subseteq$)  UD  P UD UD ? UD ? ? UD P
+ NE (==)  P  P ? P ?  D P P ? P
+Table 1: Current state of decision problems for relational pattern languages for various constraints (PS = PSPACE, NPC = NP-Complete, P = Poly.
+time decidable, UD = Undecidable, D = Decidable, E = Erasing, NE = Non-Erasing, ($\in$) membership problem, ($\subseteq$) inclusion problem, (==) equivalence problem).
+Red text marks the results covered in this paper.
 2 Preliminaries
@@
 For compactness reasons, we denote $w[i]\cdots w[j]$ by $w[i\cdots j]$ for all $i,j\in[|w|]$ with $i<j$.
+Denote by $|w|_{\mathtt{a}}=|\{\ i\in[|w|]\mid w[i]=\mathtt{a}\}|$ the number of times the letter $\mathtt{a}$ occurs in $w$.
 Set $\operatorname{alph}(w):=\{\mathtt{a}\in\Sigma\mid\exists i\in[|w|]:w[i]=\mathtt{a}\}$ as $w$’s alphabet.
+A subsequence $u$ of a word $w$ is a word obtained by deleting arbitrary letters of $w$.
+Denote by $\operatorname{SubSeq}(w)$ the set of subsequences of $w$.
+Two words $u,v\in\Sigma^{*}$ are abelian equivalent, denoted by $u\equiv_{ab}v$, iff $|u|_{\mathtt{a}}=|v|_{\mathtt{a}}$ for all $\mathtt{a}\in\operatorname{alph}(u)\cup\operatorname{alph}(v)$.
 Let $X:=\{x_{1},x_{2},$...\} be a countable set of variables such that $\Sigma\cap X=\emptyset$.
@@
 If $\Sigma$ is clear from the context, we may write just $H$.
-Given a pattern $\alpha\in Pat_{\Sigma}$, it’s erasing pattern language $L_{E}(\alpha)$ and its non-erasing pattern language $L_{NE}(\alpha)$ are defined respectively by
-$L_{E}(\alpha)$ $:=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha)\},\text{ and }$
-$L_{NE}(\alpha)$ $:=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\}.$
-Let $RegPat_{\Sigma}:=\{\alpha\in Pat_{\Sigma}\mid x_{i}\in\operatorname{alph}(\alpha)\cap X,|\alpha|_{x_{i}}=1\}$ denote the set of all regular patterns, i.e., the set of all patterns where each variable occurs only once.
+Given a pattern $\alpha\in Pat_{\Sigma}$, its erasing pattern language $L_{E}(\alpha)$ and its non-erasing pattern language $L_{NE}(\alpha)$ are defined respectively by
+$L_{E}(\alpha)$ $=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha)\},\text{ and }$
+$L_{NE}(\alpha)$ $=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\}.$
+Let $RegPat_{\Sigma}=\{\alpha\in Pat_{\Sigma}\mid x_{i}\in\operatorname{alph}(\alpha)\cap X,|\alpha|_{x_{i}}=1\}$ denote the set of all regular patterns, i.e., the set of all patterns where each variable occurs only once.
 Let $R$ be some set of relations and denote for each $n\in\mathbb{N}$ by $R_{n}\subset R$ the subset of $n$-ary relations in $R$.
-Define $Rel_{R}:=\{(r,x_{i_{1}},...,x_{i_{n}})\mid r\in R_{n},x_{i_{1}},...,x_{i_{n}}\in X,i_{1},...i_{n}\in\mathbb{N},n\in\mathbb{N}\}$ to be the set of all combinations to relate variables in $X$ over the set of relations $R$.
-Then, a pair $(\alpha,r_{\alpha})\in RegPat_{\Sigma}\times Rel_{R}$ is a relational pattern over the terminal alphabet $\Sigma$ and the set of relations $R$.
+Define $Rel_{R}=\{(r,x_{i_{1}},...,x_{i_{n}})\mid r\in R_{n},x_{i_{1}},...,x_{i_{n}}\in X,i_{1},...i_{n}\in\mathbb{N},n\in\mathbb{N}\}$ to be the set of all combinations to relate variables in $X$ over the set of relations $R$.
+Then, a pair $(\alpha,r_{\alpha})\in RegPat_{\Sigma}\times 2^{Rel_{R}}$ is a relational pattern over the terminal alphabet $\Sigma$ and the set of relations $R$.
 We denote the set of all relational patterns over $\Sigma$ and $R$ by $\mathtt{RelPat}_{\Sigma,R}$.
-For some $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $h\in H$, we say that $h$ is a $r_{\alpha}$-valid substitution if for all $(r^{\prime},x_{i_{1}},...,x_{i_{n}})\in r_{\alpha}$ we have $(h(x_{i_{1}}),...,h(x_{i_{n}}))\in r^{\prime}$, i.e., all variables are substituted such that the resulting terminal words are related to each other regarding the relation $r^{\prime}$.
+For some $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $h\in H$, we say that $h$ is a $r_{\alpha}$-valid substitution if for all $(r^{\prime},x_{i_{1}},...,x_{i_{n}})\in r_{\alpha}$ we have $(h(x_{i_{1}}),...,h(x_{i_{n}}))\in r^{\prime}$, i.e., all variables are substituted by words related to each other regarding the relation $r^{\prime}$.
 The set of all $r_{\alpha}$-valid substitutions w.r.t.
 $\Sigma$ is denoted by $H_{\Sigma,r_{\alpha}}$.
-If $\Sigma$ and $r_{\alpha}$ are clear from the context, we may also just write $H_{r_{\alpha}}$ or $H$.
-We extend the notion of pattern languages by the following.
-For any $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ we denote by
-$L_{E}(\alpha,r_{\alpha}):=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{*}\text{ for all }x\in\operatorname{var}(\alpha),h\text{ is }r_{\alpha}\text{-valid}\ \}$
-the erasing relational pattern language of $(\alpha,r_{\alpha})$ and by
-$L_{NE}(\alpha,r_{\alpha}):=\{\ h(\alpha)\ |\ h\in H,h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha),h\text{ is }r_{\alpha}\text{-valid}\ \}$
-the non-erasing relational pattern language of $(\alpha,r_{\alpha})$.
-In this paper, we primarily investigate the considered decision problems for relational pattern languages for binary relations, i.e., $|R|=1$ and if $R=\{\tilde{r}\}$, then $\tilde{r}$ is 2-ary/binary.
+If $\Sigma$ or $r_{\alpha}$ are apparent, we may also just write $H_{r_{\alpha}}$ or $H$.
+For any $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, we denote the erasing relational pattern language $L_{E}(\alpha,r_{\alpha})$ and the non-erasing relational pattern language $L_{NE}(\alpha,r_{\alpha})$ of $(\alpha,r_{\alpha})$ respectively by
+$L_{E}(\alpha,r_{\alpha})$ $=\{\ h(\alpha)\ |\ h\in H_{\Sigma,r_{\alpha}}\ \}$
+$L_{NE}(\alpha,r_{\alpha})$ $=\{\ h(\alpha)\ |\ h\in H_{\Sigma,r_{\alpha}}\text{ and }h(x)\in\Sigma^{+}\text{ for all }x\in\operatorname{var}(\alpha)\ \}.$
+In this paper, we primarily investigate the considered decision problems under binary relations, i.e., $|R|=1$ and, if $R=\{r\}$, then $r$ is 2-ary/binary.
 In particular, we are interested in the following selection of specific customary binary relations.
-Let $\mathcal{R}$ be the set containing all of the following nine relations.
+Let $\mathcal{R}$ be the set containing all of the following relations.
   1. 1.
@@
     $r_{*}$ : word in star of the other word, i.e., $(u,v)\in r_{*}$ iff $u\in\{v\}^{*}$.
-In the proofs of the main results regarding the inclusion problem, we make use of two different automata models that have undecidable language emptiness properties.
-The first one is the so called nondeterministic 2-counter automaton without input (see, e.g., [Iberra1978]).
-The second one is a very specific universal Turing machine $U$ with 2 symbols and 15 states (see [Neary2009]).
-As the technical proofs are very lengthy and based on the constructions provided in other papers, the formal definitions of the models can be found in Appendix  0.B.1 and Appendix  0.B.2 respectively.
-For the intuitions of the proofs provided in the main body, it suffices to know that these respective models have undecidable language emptiness properties, that all terminating/accepted computations are encoded using finite binary words, and that the number of existing and not existing outgoing transitions per state are bound by a constant.
-This last property results in a finite number of properties of words that are not encodings of accepting computations.
-3 The Membership Problem
-In this section, we discuss the membership problem for relational pattern languages for the relations in $\mathcal{R}$.
-In addition to that, we also consider certain special cases of one-sided relations mentioned at the end of [Angluin1980] that do not immediately translate to the framework of relational pattern languages considered here.
-As to be seen, the membership problem remains NP-complete for all considered cases.
-We will begin by covering the cases arising for all relations in $\mathcal{R}$.
-For the case of $r_{=}$, we recall the original result by Angluin in [Angluin1980] in which it is shown that the membership problem of non-erasing pattern languages is NP-complete.
-In the erasing case, the same result has been shown by Jiang et al.
-in [Jiang1994].
-The following theorem states the main result of this section.
-It is the NP-completeness of the membership problem for relational pattern languages over any of the relations in $\mathcal{R}$.
-The rest of this section serves as its proof.
-Theorem 3.1
-Let $r\in\mathcal{R}$ and assume $R=\{r\}$.
-Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{X}(\alpha,r_{\alpha})$ for $X\in\{E,NE\}$ is NP-complete.
-Indeed, the NP-containment of the membership problems for any of the relations follows by a rather similar argument as can be made for the case of the equality relation $r_{=}$ (see Appendix 0.A for more details).
-Proposition 1(✂)
-Let $r\in\mathcal{R}$ and $R=\{r\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{X}(\alpha,r_{\alpha})$ for $X\in\{E,NE\}$ is in NP.
-The Non-Erasing Cases: To show NP-hardness, we begin with the non-erasing case and start with the relations for which the exact same construction as in [Angluin1980] can be applied.
-An overview of the original proof together with respective arguments why it works for all these other relations can be found in Appendix 0.A.
-Lemma 1(✂)
-Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$.
-In the case of $r=r_{hd,k}$, assume $k\in\mathbb{N}$.
-Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
-That leaves us only with the cases of commutation.
-Notice that for the non-erasing cases, $r_{com^{*}}$ and $r_{com^{+}}$ result in the same valid substitutions as the empty word is excluded from valid substitutions.
-Proposition 2(✂)
-Let $r\in\{r_{com^{*},com^{+}}\}$.
-Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
-Again, the full technical proof can be found in Appendix 0.A.
-To give an intuition, this proof also relies on a reduction from the $3\mathtt{-SAT}$ problem and has a similar nature as the proof used for Lemma˜1 which is based on the proof from [Angluin1980].
-For the adapted construction to work, we only use two different kinds of factors instead of three.
-First, a part that handles the variable’s true and false encodings, by setting the factor in $w$ to 1#11\#1 and the factor in $\alpha$ to $u_{i}v_{i}$, for each $3\mathtt{-SAT}$ variable $X_{i}$ (resulting in only either $u_{i}$ or $v_{i}$ to be substituted to contain a #\#).
-For the parts handling the clauses, we use the factor $1^{10}\#1^{10}\#1^{10}$ in $w$ and the factor $z_{j,1}\ x_{j,1}\ z_{j,2}\ x_{j,2}\ z_{j,3}\ x_{j,3}\ z_{j,4}$ in $\alpha$, where each $x_{j,i}$ corresponds to some variable related to either some $u_{k}$ or $v_{k}$ and each $z_{j_{i}}$.
-By that, at least one variable in each clause must not contain a #\#, resulting in the reduction to go through.
-This concludes the non-erasing cases.
-The Erasing Cases: Now, we continue with the erasing case.
-As in the non-erasing cases, for many relations, the existing construction (this time from [Jiang1994]) can be directly applied to show NP-hardness (see Appendix 0.A for some more details).
-Lemma 2(✂)
-Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$.
-In the case of $r=r_{hd,k}$, assume $k\in\mathbb{N}$.
-Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
-That leaves us with ++-commutation and *-commutation.
-First of all, for the case of $r_{com^{+}}$, we see that the proof used to show Proposition˜2 also applies for the erasing case, as two related variables are never substituted to the empty word (see Appendix 0.A for more details).
-Proposition 3(✂)
-Let $R=\{r_{com^{+}}\}$.
-Given some word $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
-Notice that $r_{com^{*}}$, however, always allows for one of two related variables to be substituted by the empty word.
-Therefore, we need a slightly different construction for this to work.
-Based on the proof of Proposition Proposition˜2, we can change the clause encoding in $w$ to only the letter $1$ instead of $1^{10}\#1^{10}\#1^{10}$ and remove all $z$ variables for the construction to work.
-The basic idea is that is a variable is substituted to contain the letter #\#, it cannot be substituted to a single $1$, reversing the idea used in Proposition˜2 (See Appendix 0.A).
-Proposition 4(✂)
-Let $R=\{r_{com^{*}}\}$.
-Given some word $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{E}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
-This concludes the NP-hardness results for the membership problem under all relations considered and, thus, the proof of Theorem 3.1.
-As mentioned in the beginning of this section, we also obtain NP-hardness in the cases of restricting the constructed patterns to one-sided subsequence or one-sided star relations.
-We obtain the following results by proofs that are similar to the ones used for the previous results.
-Due to space constraints and their special nature, the discussion of the following result can be found in Appendix 0.A.1.
-Theorem 3.2
-Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x_{i}$ occurs at most once, and, if it does, the related variables $x^{*}$ (resp.
-$x^{s}$) [^1] For the discussion of restrictions to one-sided relations, we deviate from the notation used in all other parts of this paper and use a notation similar to the one introduced in the end of the work by Angluin[Angluin1980].
-may occur multiple times, it is NP-complete to decide whether $w\in L_{X}(\alpha)$, where $X\in\{E,NE\}$.
+  10. 10.
+    $r_{\varepsilon_{=}}$ : common (non)emptiness, i.e., $(u,v)\in r_{\varepsilon_{=}}$ iff $u=\varepsilon=v$ or $u,v\in\Sigma^{+}$.
+3 The Equivalence Problem
+In this first main section, we cover the equivalence problem for relational pattern languages.
+We recall that in the erasing case, this problem is one of the main open problems for the $r_{=}$ relation.
+We obtain a series of distinct results.
+For the non-erasing case, we give a characterization of a set of relations under which equivalence is characterized by equality of patterns and equality of the transitive, reflexive, and symmetric closures of the given relational constraints.
+Related to the considered relations in this paper, this answers the cases for $r_{ab}$, $r_{com^{+}}$, and clearly also $r_{=}$.
+In contrast to that, we observe for other relations, namely $r_{rev}$, $r_{ssq}$ and $r_{*}$, that deciding equivalence in the non-erasing case cannot be characterized merely by such an equality property, showing a clear boundary to the already known trivial decidability of equivalence for $r_{=}$.
+For the erasing case, most notably, we can show the existence of another relation besides $r_{|w|}$, namely $r_{\varepsilon_{=}}$, that allows for general decidability of the equivalence problem.
+On the other hand, for some of the considered relations, namely $_{rev}, rr_{*}$, and $r_{ssq}$, it can be shown that the equivalence problem must be at least as hard as the equivalence problem for $r_{=}$.
+3.1 The Non-Erasing Case
+Given any relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, we denote by $r_{\alpha}]$ the reflexive, symmetric, and transitive closure on the set of binary related variables in $r_{\alpha}$.
+For example, if $(r_{=},x_{1},x_{2}),(r_{=},x_{2},x_{3})\in r_{\alpha}$, then we would have $(r_{=},x_{i},x_{j})\in[r_{\alpha}]$, for any combination $i,j\in\{1,2,3\}$.
+For the remainder of this section, we assume w.l.o.g.
+a normal form on the patterns such that all variables are always introduced inside a pattern with increasing indices, starting from 1 (i.e.
+$x_{1}$, $x_{2}$, $x_{3}$, and so on).
+The first results gives us a strong condition for relations $r$ that is reflexive on all letters of the alphabet and does not relate two distinct letters.
+Lemma 1.
+Assume $|\Sigma|\geq 2$ and let $R=\{r\}$, for a relation $r$ with the following properties: For all $\mathtt{a}\in\Sigma$, we have $(\mathtt{a},\mathtt{a})\in r$, and if $(\mathtt{a},\mathtt{b})\in r$, for two letters $\mathtt{a},\mathtt{b}\in\Sigma$, then $\mathtt{a}=\mathtt{b}$.
+Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two patterns in normal form.
+If $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ then $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
+Proof.
+First, we see that $|\alpha|=|\beta|$ must hold.
+Suppose otherwise and assume w.l.o.g.
+that $|\alpha|<|\beta|$.
+Take some $r_{\alpha}$-valid substitution $h$ that substitutes each variable by just a single letter.
+Then $|h(\alpha)|=|\alpha|$.
+In the NE-case, there can’t exist any $r_{\beta}$-valid substitution $h^{\prime}$ with $|h^{\prime}(\beta)|=|\alpha|<|\beta|$.
+A contradiction.
+Now suppose $\alpha\neq\beta$.
+As both are regular patterns in normal form, that means there exists a first position $i\in[|\alpha|]$ such that w.l.o.g.
+$\alpha[i]=\mathtt{a}$, for some letter $\mathtt{a}\in\Sigma$, and $\beta[i]=x_{j}$, for some variable $x_{j}\in X$.
+Now let $h$ be a $r_{\beta}$-valid substitution that substitutes each variable by just a single letter and, in particular, set $h(x_{j})=\mathtt{b}$, for some letter $\mathtt{b}$ with $\mathtt{a}\neq\mathtt{b}$.
+As $|\alpha|=|\beta|$, any $r_{\alpha}$-valid substitution $h^{\prime}$ with $h^{\prime}(\alpha)0h(\beta)$ would have to match $h^{\prime}(\alpha[k])=h(\beta[k])$, for all $k\in[|\alpha|]$, and, in particular, would need to match $h^{\prime}(\alpha[i])=h^{\prime}(\mathtt{a})=\mathtt{a}\neq\mathtt{b}=h(x_{j})=h(\beta[i])$.
+A contradiction.
+That concludes $\alpha=\beta$.
+Now suppose $r_{\alpha}]\neq[r_{\beta}]$.
+Then there exist two position $i,j\in[|\alpha|]$ such that $\alpha[i]=\beta[i]=x_{i^{\prime}}$, for some $i^{\prime}\in[|\alpha|]$, and $\alpha[j]=\beta[j]=x_{j^{\prime}}$, for some $j^{\prime}\in[|\alpha|]$, and w.l.o.g.
+$(r,x_{i^{\prime}},x_{j^{\prime}})\in[r_{\alpha}]$ but $(r,x_{i^{\prime}},x_{j^{\prime}})\notin[r_{\beta}]$.
+Now choose a substitution $h$ that substituted each variable just by a single letter such that $h(x_{i^{\prime}})=\mathtt{a}$ and $h(x_{j^{\prime}})=\mathtt{b}$, for two distinct letters $\mathtt{a},\mathtt{b}\in\Sigma$.
+As $r$ forces all blocks of symmetrically and transitively connected variables to be substituted by the same letter, in the case of length 1 substitutions, we know that $h$ cannot be $r_{\beta}$-valid.
+However, there exists such an $h$ that is $r_{\alpha}$-valid by choosing substitutions of the other variables accordingly.
+Hence, there exists such an $h(\alpha)\in L_{NE}(\alpha,r_{\beta})$, but we always have $h(\beta)\notin L_{NE}(\beta,r_{\beta})$.
+A contradiction.
+∎
+This result already confirms a very specific structure for all relations with the above stated properties.
+In particular, this covers the relations $r_{=}$, $r_{ab}$, $_{ssq}, rr$$_{rev}$, $r_{com^{*}}/r_{com^{*}}$ (equivalent for non-erasing pattern languages) and $r_{*}$.
+This property, however, is not exhaustive to characterize equivalence.
+So, adding to the previous property, if we now consider equivalence relations in general, we obtain the following result.
+Lemma 2(✂0.A.1).
+Let $R=\{r\}$ for an equivalence relation $r$.
+Let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ be some relational pattern in normal form.
+We have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$.
+Bringing the previous two kinds of relations together, we can immediately identify a set of relations for which equivalence of non-erasing languages can be characterized by equality of patterns and the corresponding closures.
+Theorem 3.1(✂0.A.2).
+Let $R=\{r\}$ for $r$ being an equivalence relation with the following properties: For all $\mathtt{a}\in\Sigma$, we have $(\mathtt{a},\mathtt{a})\in r$, and if $(\mathtt{a},\mathtt{b})\in r$, for two letters $\mathtt{a},\mathtt{b}\in\Sigma$, then $\mathtt{a}=\mathtt{b}$.
+Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two patterns in normal form.
+$L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ if and only if $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
+As the relations $r_{=}$, $r_{ab}$, and $r_{com^{+}}$ (also $r_{com^{*}}$ as it is the NE case) have the properties required in Theorem˜3.1, we can say the following.
+Corollary 1(✂0.A.3).
+Let $R=\{r\}$ for some $r\in\{r_{=},r_{ab},r_{com^{+}}\}$.
+Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, we can check in polynomial time whether $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
+Remark 1.
+Notice that the relation $r_{\varepsilon_{=}}$ has no effect in the non-erasing case.
+In particular, for any relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r_{\varepsilon_{=}}\}$ we have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha)$ for the regular pattern $\alpha$.
+Hence, equivalence between two such patterns is decidable in linear time in this case [3].
+The other relations considered in Lemma˜1 that are not covered in Corollary˜1, namely $_{rev}, rr$$_{rev}$, $r_{*}$, are clearly not equivalence relations.
+Hence, decidability is not immediately clear in these cases.
+Here, blocks of related variables imply interesting structures on their own.
+Thus, they probably require a deeper investigation to answer (efficient) decidability in general (except the already known general decidability for $r_{rev}$ [30]).
+We conjecture they are all decidable.
+3.2 The Erasing Case
+In this part, our primary focus lies on the relation $r_{\varepsilon_{=}}$.
+First, note that $r_{\varepsilon_{=}}$ is reflexive, symmetric, and transitive.
+We extend the notation of the prior section by defining equivalence classes $\mathcal{C}_{\alpha}=\{C_{\alpha,1},\dots,C_{\alpha,k_{\alpha}}\}$ of variables $C_{\alpha,i}\subseteq\operatorname{var}(\alpha)$, $i\in[k_{\alpha}]$ that are in relation with one another.
+In the following, we prove that this relation results in decidability in the context of the erasing equivalence problem.
+For general patterns, we first present an polynomial space approach before giving linear-time and polynomial-time algorithms for patterns that satisfy specific properties.
+Theorem 3.2(✂0.A.4).
+Let $R=\{r_{\varepsilon_{=}}\}$ and $\Sigma$ be some alphabet.
+Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, checking $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ is decidable in polynomial space.
+Please consider Appendix 0.A.4 for a more detailed proof.
+Here, we provide a rough sketch.
+First, consider $(\alpha,r_{\alpha})$ and for each subset $S\subseteq\mathcal{C}_{\alpha}$, replace the variables $v\in\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by any word in $\Sigma^{+}$ and any other variable $v\notin\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by $\varepsilon$.
+We call the resulting language $L_{E}(\alpha,r_{\alpha},S)$.
+Using that, we obtain $L_{E}(\alpha,r_{\alpha})=\bigcup_{S\subseteq\mathcal{C}_{\alpha}}L_{E}(\alpha,r_{\alpha},S)$.
+Do the same for $\beta$.
+Note there exist at most $2^{k_{\alpha}}\leq 2^{|\alpha|}$ subsets $S$ for $\alpha$ and at most $2^{k_{\beta}}\leq 2^{|\beta|}$ subsets for $\beta$.
+For each $S\subseteq\mathcal{C}_{\alpha}$, we can take a sub-pattern $\alpha_{S}$ of $\alpha$ that only contains the variables that are not substituted by the empty word.
+As $\alpha_{S}$ is a regular pattern, it can be observed that $L_{E}(\alpha,r_{\alpha},S)=L_{NE}(\alpha_{s})$ and we can construct an NFA $A_{\alpha_{S}}$ with $L(A_{\alpha_{S}})=L_{NE}(\alpha_{S})=L_{NE}(\alpha,r_{\alpha},S)$.
+Doing that for each such $S$, combining all these NFAs into one big NFA $A_{\alpha}$ for $L_{E}(\alpha,r_{\alpha})$ and doing the same for $(\beta,r_{\beta})$ in an NFA $A_{\beta}$, we can compare the languages $L_{E}(\alpha,r_{\alpha})$ and $L_{E}(\beta,r_{\beta})$ by comparing $L(A_{\alpha})=L(A_{\beta})$.
+The latter is well known to be PSAPCE-complete (e.g., see [1]), concluding the result.
+Next, we introduce approaches with an improved running time for some pattern structures.
+Lemma 3(✂0.A.5).
+Let $R=\{r_{\varepsilon_{=}}\}$ and $\Sigma$ be some alphabet. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be some relational patterns.
+  1. 1.
+    If $\alpha$ and $\beta$ are terminal-free, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
+  2. 2.
+    If $|\Sigma|=1$, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
+  3. 3.
+    If there exists $c\in\Sigma$ that is not a terminal in $\alpha$ and $\beta$, we can check in polynomial time whether $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
+Finally, we consider some other relations in the context of the equivalence problem for erasing relational pattern languages.
+In particular, we can show that this problem must be at least as hard as in the case of $r_{=}$ for the relations $r_{*}$, $r_{ssq}$, and $r_{rev}$.
+While the first two follow trivially by simulating $r_{=}$, the last one requires a more technical construction (given in Appendix 0.A.6).
+Proposition 1(✂0.A.6).
+Let $R=\{r\}$ for $r\in\{r_{ssq},r_{*},r_{rev}\}$.
+Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
+Deciding $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ is at least as hard as deciding the equivalence problem for erasing pattern languages.
 4 The Inclusion Problem
 In this section, we continue with the inclusion problem for relational pattern languages.
-We recall that the inclusion problem for non-erasing as well as erasing pattern languages over $r_{=}$ is generally undecidable for all alphabets $\Sigma$ of size $|\Sigma|\geq 2$, as shown and refined in  [Jiang1995, Freydenberger2010, BREMER201215].
+We recall that the inclusion problem for non-erasing as well as erasing pattern languages over $r_{=}$ is generally undecidable for all alphabets $\Sigma$ of size $|\Sigma|\geq 2$, as shown and refined in  [23, 13, 5].
 As we will see in this section, for several of the considered relations in $\mathcal{R}$, we also obtain that this problem is undecidable in the non-erasing or erasing cases.
-In fact, if we consider, first, the relations $r_{ssq}$ and $r_{*}$, we observe that we can simulate the exact behavior of $r_{=}$, similar to the NP-hardness proofs in the membership section.
-Meaning, if we take any two variables $x,y\in\operatorname{var}(\alpha)$, for some relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, and any of the three relations above, we see that relating these variables in both direction effectively forces two variables to be substituted equally in $r_{\alpha}$ valid substitutions, i.e., if $(x,y),(y,x)\in r_{\alpha}$, then $h\in H_{r_{\alpha}}$ if and only if $h(x)=h(y)$.
-By the fact that this problem is undecidable for $r_{=}$, this results in the following, for which the formal proof can be found in Appendix 0.B.3 for reference.
-Corollary 1
-[Jiang1995, Freydenberger2010, BREMER201215] Let $\Sigma$ be any finite alphabet with $|\Sigma|\geq 2$.
-Let $\tilde{r}\in\{r_{=},r_{ssq},r_{*}\}$ and let $R=\{\tilde{r}\}$.
+However, there also exist relations for which the problem is decidable.
+First, consider the relations $r_{ssq}$ and $r_{*}$.
+We observe that we can simulate the exact behavior of $r_{=}$.
+Meaning, if we take any two variables $x,y\in\operatorname{var}(\alpha)$, for some relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, and any of the three relations above, we see that relating these variables in both directions effectively forces two variables to be substituted equally in $r_{\alpha}$-valid substitutions, i.e., if $(x,y),(y,x)\in r_{\alpha}$, then $h\in H_{r_{\alpha}}$ if and only if $h(x)=h(y)$.
+By the fact that this problem is undecidable for $r_{=}$, this results in the following.
+Corollary 2.
+[23, 13, 5] Let $\Sigma$ be any finite alphabet with $|\Sigma|\geq 2$.
+Let $r\in\{r_{=},r_{ssq},r_{*}\}$ and let $R=\{r\}$.
 Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$.
@@
 4.1 The Non-Erasing Cases
-In the non-erasing case, an interesting observation regarding $r_{ab}$ can be found.
-Reconstructing the reduction from [BREMER201215] in the setting of relational pattern languages, it can be shown that the inclusion problem for non-erasing relational pattern languages under $r_{ab}$ is actually also undecidable, even though $r_{ab}$ generally implies a significantly less restrictive behavior than $r_{=}$.
-Proposition 5(✂)
-Let $R:=\{r_{ab}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$.
-It is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
-Proof
-This can be shown by reducing the problem of deciding whether the universal Turing machine $U$ has any valid computations to the problem of whether the NE-relational pattern language of one relational pattern under abelian equivalence is included in the NE-relational pattern language of another relational pattern under abelian equivalence.
-It is done by constructing two such relational patterns $(\alpha,\alpha_{R})$ and $(\beta,\beta_{R})$ for which we have that $L_{NE}(\alpha,R)\subseteq L_{NE}(\beta,\beta_{R})$ if and only if $U$ has no valid computation.
-As mentioned above, this result follows from the same reduction as in [BREMER201215] for the inclusion problem of NE-pattern languages over the equality relation.
-Only some technical modifications have to be made to formally get it to work in the setting of relational pattern languages under $r_{ab}$.
-The proof covering the technical adaptions can be found in Appendix 0.B.4 for reference.
+In the non-erasing case, first notice that, similar to the equivalence problem, in the case of $r_{\varepsilon_{=}}$, we basically compare two regular patterns (as $\varepsilon$ substitutions are excluded anyways).
+Hence, this case is clearly decidable.
+Additionally, an interesting observation regarding $r_{ab}$ can be found.
+Reconstructing the reduction from [5] in the setting of relational pattern languages, it can be shown that the inclusion problem for non-erasing relational pattern languages under $r_{ab}$ or $r_{rev}$ are actually also undecidable, even though $r_{ab}$ generally implies a significantly less restrictive behavior than $r_{=}$ and $r_{rev}$ implies a generally distinct structure.
+Theorem 4.1(✂0.B.1).
+Let $R:=\{r\}$ for $r\in\{r_{ab},r_{rev}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$.
+For relational patterns for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$.
+.
+Indeed, this can be shown by reducing the problem of deciding whether a very specific universal Turing machine $U$ has any valid computations to the problem of whether the NE-relational pattern language of one relational pattern under abelian equivalence is included in the NE-relational pattern language of another relational pattern under abelian equivalence.
+It is done by constructing two such relational patterns $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ for which we have that $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,\beta_{\alpha})$ if and only if $U$ has no valid computation.
+As mentioned above, this result follows from the same reduction as in [5] for the inclusion problem of NE-pattern languages over the equality relation.
+This is due to the fact that related variables under $r_{ab}$ and $r_{rev}$ result in equal substitutions if only letters or longer but unary words are considered.
+Hence, the general reasoning of the construction carries over.
+Only some technical modifications have to be made to formally get it to work in the setting of relational pattern languages under $r_{ab}$ and $r_{rev}$.
+For the sake of completeness and for verification purposes, an extended sketch of the proof idea can be found in Appendix 0.B.1.
+4.2 The Erasing Cases
+We continue with erasing relational pattern languages.
+We show that the inclusion problem is decidable under $r_{\varepsilon_{=}}$ and undecidable under the reversal relation $r_{rev}$ and under abelian equivalence $r_{ab}$.
+Indeed, in the case of $r_{\varepsilon_{=}}$, we can use the exact same approach as for the equivalence problem from Theorem 3.2 to obtain decidability in polynomial space using the construction of two NFAs that represent $L_{E}(\alpha,r_{\alpha})$ and $L_{E}(\beta,r_{\beta})$ respectively.
+Theorem 4.2.
+Let $R=\{r_{\varepsilon_{=}}\}$.
+Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, checking $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ is decidable in polynomial space.
+Again, for some pattern structures, we introduce some approaches with improved polynomial running time.
+Lemma 4(✂0.B.2).
+Let $R=\{r_{\varepsilon_{=}}\}$. Given some alphabet $\Sigma$ and relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
+  1. 1.
+    If $\alpha$ and $\beta$ are terminal-free, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
+  2. 2.
+    If $|\Sigma|=1$, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
+For the undecidability proofs, we use the same framework as in the result from Freydenberger and Reidenbach [13] that shows undecidability in the erasing case under $r_{=}$, for all alphabets $\Sigma$ with $|\Sigma|\geq 2$, as a starting point, but we need to adapt it at several points in a more involved way to get the following theorem.
+Theorem 4.3(✂0.B.3).
+Let $\bar{r}\in\{r_{rev},r_{ab}\}$ and let $R=\{\bar{r}\}$.
+For all alphabets $\Sigma$ with $|\Sigma|\geq 2$, given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is generally undecidable whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
+Analogously to the proof of Theorem˜4.1, a full proof sketch for Theorem˜4.3 is given in Section˜0.B.3.
+5 The Membership Problem
+In this final section, we briefly discuss the membership problem – also known as the matching problem – for relational pattern languages for all relations in $\mathcal{R}$ and general patterns.
+For the case of $r_{=}$, NP-completeness of the membership was originally shown in [3] and [22] for the non-erasing and erasing cases respectively.
+In fact, independent communities have shown this under several considered restrictions (cf.
+[11, 28]).
+As shown in this section, all but one cases remain NP-complete.
+Proposition 2.
+Let $r\in\mathcal{R}$ and $R=\{r\}$. Let $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. Deciding $w\in L_{X}(\alpha,r_{\alpha})$, for $X\in\{E,NE\}$, is in NP.
+Indeed, NP-containment follows by the known arguments from the cases arising in $r_{=}$.
+We can guess a substitution $h$ with $|h(\alpha)|=|w|$, check the relational constraints in polynomial time, and finally check whether $h(\alpha)=w$.
+Regarding NP-hardness, we revisit the exposition given in [28], where a reduction is given from the known NP-complete positive 1-in-3-SAT (no negated variables, 3-CNF clauses, satisfiable with exactly one literal per clause set to true) problem.
+For example, assuming $\Sigma=\{0,\#\}$, the formula
+$\{\{X_{1},X_{2},X_{3}\},\{X_{2},X_{4},X_{5}\},\{X_{1},X_{4},X_{5}\},\{X_{3},X_{4},X_{6}\}\}$
+is in 3-CNF and 1-in-3 satisfiable if and only if the word w=0#0#0#0w=0\#0\#0\#0 is in the erasing pattern language of the pattern $\alpha=x_{1}x_{2}x_{3}\#x_{2}x_{4}x_{5}\#x_{1}x_{4}x_{5}\#x_{3}x_{4}x_{6}$.
+Using $w=0^{4}\#0^{4}\#0^{4}\#0^{4}$, this construction works for the NE case as well.
+For most of the relations considered in this paper, this framework suffices.
+Lemma 5(✂0.C.1).
+Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*},r_{com^{+}},r_{\varepsilon_{=}}\}$ s.t.
+$R=\{r\}$ and let $|\Sigma|\geq 2$.
+For $w\in\Sigma^{*}$, $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding $w\in L_{E}(\alpha,r_{\alpha})$ is NP-hard.
+If $r\notin\{r_{com^{+}},r_{\varepsilon_{=}}\}$, deciding $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard.
+This leaves us with the cases of $r_{com^{+}}(r_{com^{*}})$ in the non-erasing case and with $r_{com^{*}}$ in the erasing case.
+For these, a slightly different reduction is necessary but possible to achieve.
+In Appendix 0.C.2, a reduction from the 3-SAT problem is given, concluding the following result.
+Theorem 5.1(✂0.C.2).
+Let $r\in\mathcal{R}$ s.t.
+$R=\{r\}$ and assume $|\Sigma|\geq 2$.
+For $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding $w\in L_{E}(\alpha,r_{\alpha})$ is NP-complete.
+If $r\neq r_{\varepsilon_{=}}$, deciding $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-complete.
+Finally, we recall from Remark˜1 that the relation $r_{\varepsilon_{=}}$ has no effect in the non-erasing case, resulting in each pattern language $L_{NE}(\alpha,r_{\alpha})$, for $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ and $R=\{r_{\varepsilon_{=}}\}$, being equal to the pattern language $L_{NE}(\alpha)$ of the regular pattern $\alpha$.
+The membership problem for regular patterns has been shown to be decidable in linear time by Manea et al.
+in [27].
+Hence, the following immediately holds.
+Corollary 3.
+Let $R=\{r_{\varepsilon_{=}}\}$. For $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, we can decide in linear time whether $w\in L_{NE}(\alpha,r_{\alpha})$.
+6 Final Discussion
+We considered the three main decision problems equivalence, inclusion, and membership problems for relational pattern languages under various typical binary relations, i.e., $r_{=}$, $r_{ssq}$, $r_{ab}$, $_{perm}, rr$$_{rev}$, $r_{com^{+}}$, $_{com^{*}}, rr_{*}$, and $r_{\varepsilon_{=}}$.
+While obtaining a significant number of results, different research directions building on the established foundation can be worthwhile pursuing.
+Regarding the equivalence problem, we obtained decidability in polynomial time for $r_{\varepsilon_{=}}$, $r_{ab}$ and $r_{com^{+}}$ in the non-erasing case, similar to the known cases $r_{=}$ and $r_{|w|}$.
+For the relations $r_{rev}$, $r_{ssq}$, and $r_{*}$ we obtained strict structural implications of patterns with equal languages.
+A clear answer regarding (efficient) decidability (except the known general decidabilit for $r_{rev}$), however, probably requires a more in depth analysis.
+For the erasing cases, most significantly, the relation $r_{\varepsilon_{=}}$ results in decidability.
+Together with the known result for $r_{|w|}$, this marks one of the few relations with a clear positive answer with respect to decidability in the erasing case.
+We recall that this problem for $r_{=}$ remains the main open problem for decision problems in pattern languages.
+In addition to that, we could establish that this problem under the relations $_{rev}, rr_{*}$, and $r_{ssq}$ must be at least as hard to solve as for $r_{=}$, extending the decidability question in that direction.
+For the inclusion problem, we showed that $r_{\varepsilon_{=}}$ results in decidability.
+Together with $r_{|w|}$, this is the only other relation currently known to do so.
+In contrast, inclusion is undecidable for the relational patterns under $r_{=}$, $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$.
+These undecidability results arise either because the relations can simulate $r_{=}$ in general, as with $r_{ssq}$ and $r_{*}$, or because they still allow for the key properties used in established reductions to be enforced, such as equality of letters or unary words, as with $r_{ab}$ and $r_{rev}$.
+It remains open whether similar results can be obtained for structurally weaker relations.
+For the remaining cases $r_{perm}$, $r_{com^{*}}$, and $r_{com^{+}}$, the crucial reduction properties do not seem to be achievable in an obvious way.
+The relation $r_{perm}$ does not enforce equal letters or equal unary words over alphabets of size at least two, while $r_{com^{+}}/r_{com^{*}}$ do not even enforce equal lengths of substituted words.
+Hence, the boundary between (un)-decidability for weaker relations remains an interesting open question.
+Finally, for the membership problem, we observed that in all considered cases (except $r_{\varepsilon_{=}}$ in the non-erasing case), that is, for all relations in $\mathcal{R}$ in the erasing case and all relations in $\mathcal{R}\setminus\{r_{\varepsilon_{=}}\}$ in the non-erasing case, the problem turns out to be NP-complete.
+Interestingly, this even holds for structurally very weak relations such as $r_{\varepsilon_{=}}$ or $r_{com^{*}}/r_{com^{+}}$.
+As generally assumed, searching for words in sets of texts that are compactly described by, e.g., a relational pattern of almost any kind, turns out to be a generally computationally hard problem.
+This concludes this paper. Again, for a visual summary, please see Table 1.
+Acknowledgements.
+We thank the anonymous reviewers of the previous version of this paper for their valuable comments and suggestions.
+In particular, their remarks on the membership problem helped us to better position our results within the context of more recent research.
+Appendix 0.A Additional Content for the Equivalence Problem
+0.A.1 Proof of Lemma˜2
+Proof.
+($\Rightarrow$): First, let $w\in L_{NE}(\alpha,r_{\alpha})$.
+Then there exists some $r_{\alpha}$-valid substitution $h$ with $h(\alpha)=w$.
+Let $(r,x_{i},x_{j})\in r_{\alpha}$, for two variables $x_{i},x_{j}\in\operatorname{var}(\alpha)$.
+Then $(h(x_{i}),h(x_{j}))\in r$ by definition.
+As $r$ is an equivalence relation.
+we also have $(h(x_{j}),h(x_{i}))\in r$, $(h(x_{i}),h(x_{i}))\in r$, and, if there exists some $(r,x_{j},x_{k})\in r_{\alpha}$ (hence $(h(x_{j}),h(x_{k}))\in r$), we also have $(h(x_{i}),h(x_{k}))\in r$.
+Hence, $h$ satisfies all constraints in $r_{\alpha}]$ and thereby is $r_{\alpha}]-$valid, resulting in $w\in L_{NE}(\alpha,[r_{\alpha}])$.
+($\Leftarrow$): For the other direction, let $w\in L_{NE}(\alpha,[r_{\alpha}])$ and let $h$ be an $r_{\alpha}]-$valid substitution such that $h(\alpha)=w$.
+Notice that $r_{\alpha}\subseteq[r_{\alpha}]$.
+Hence, all constraints in $r_{\alpha}$ are satisfied by $h$, resulting in $h$ being $r_{\alpha}$-valid.
+Thus, $w\in L_{NE}(\alpha,r_{\alpha})$.
 ∎
-As the construction used to show Proposition 5 only related variables to either enforce equality between length 1 substitutions (letters) or equality between unary strings, it also not only translates to the case of $r_{ab}$, but also to the reversal relation $r_{rev}$.
-Indeed, the reversal of a unary word $w$ is $w$ itself and the reversal of a letter $\mathtt{a}_{i}$ is clearly also just the letter $\mathtt{a}_{i}$.
-So, we can enforce the same behavior to select only one $x_{i,1}$ to become the letter #\# in valid substitutions while all other $x_{j,i}$ take the letter 0, as $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ are always substituted only to unary words that are powers of the letter 0.
-With that, the $v$ parts will look exactly the same.
-Further, the counter values that are represented in simple predicates with the variables $\hat{x}_{i,j}$ will also only take the shape of unary words, as do the variables in the third predicate $\pi_{3}$ in the remaining cases not covered by the predicates $\pi_{1}$ and $\pi_{2}$.
-So, using an analogous construction as for Proposition 5, we obtain the following.
-Corollary 2
-Let $R:=\{r_{rev}\}$ and $|\Sigma|\geq 2$ for some alphabet $\Sigma$.
-It is undecidable whether $L_{NE}(\alpha,r_{\alpha})\subseteq L_{NE}(\beta,r_{\beta})$ for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
-4.2 The Erasing Cases
-We continue with the additional cases for erasing relational pattern languages.
-First, it can be shown that the inclusion problem under the reversal relation $r_{rev}$ is also undecidable.
-We can use the same framework as in the result from Freydenberger and Reidenbach [Freydenberger2010] that shows undecidability in the erasing case under $r_{=}$, for all alphabets $\Sigma$ with $|\Sigma|\geq 2$, as a starting point, but we need to adapt it at several points in a more involved way to get the following theorem.
-Similarly to Proposition 5, in Appendix 0.B.5, we provide, for reference and completeness, a full reconstruction of the proof from [Freydenberger2010] in the framework of relational pattern languages with adaptations where they are necessary for the proof to work for $r_{rev}$.
-Theorem 4.1(✂)
-Let $R=\{r_{rev}\}$.
-For all alphabets $\Sigma$ with $|\Sigma|\geq 2$, given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is generally undecidable whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
-Next, as for the non-erasing case, we can do something similar to show undecidability of the inclusion problem also relational pattern languages under abelian equivalence $r_{ab}$.
-Theorem 4.2
-Let $\Sigma$ be a finite alphabet with $|\Sigma|\geq 2$ and let $R=\{r_{ab}\}$.
-Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$.
-Proof
-This proof is analogous to the one of Theorem˜4.1.
-Therefore, we only describe how to get to the analogy and refer for a detailed proof to Theorem˜4.1.
-Let $A$ be a nondeterministic 2-counter automaton.
-Consider
-$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
-where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}0$ and
-$\beta_{A}:=x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}\#^{6}\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}\#^{6}\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}}$
-with, for all $i\in\{1,\ldots,\mu\}$, and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{4}},x_{1_{5}},\ldots$ are distinct variables with $(x_{i},x_{i^{\prime}}),(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ and all $\eta_{i}\in X^{*}$ are terminal-free patterns for some $\mu$ and some $\hat{\beta}_{i}$ as in Theorem˜4.1.
-We have $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ with distinct variables $z_{i},z_{i^{\prime}},\hat{z}_{i_{1}},\hat{z}_{i_{2}},\hat{z}_{i_{3}},\hat{z}_{i_{4}},\hat{z}_{i_{5}}$ and $(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in R$.
-The main concept of the proof of Theorem˜4.1 is to show that iff $L_{E}(\alpha_{A},r_{\alpha_{A}})\subseteq L_{E}(\beta_{a},r_{\beta_{A}})$ then $A$ has no accepting computation.
-All possible non accepting computations are listed in the $\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}$ part where each $\hat{\beta}_{i},i\in\{1,\ldots,\mu\}$ encodes one non accepting computation together with selecting variables $x_{i_{j}}$ for $j\in{1,2,3}$.
-The part of $\beta_{A}$ and $\alpha_{A}$ before the first $^{6}$ ensures that exactly one pair $x_{i},x_{i}^{\prime},i\in\{1,\ldots,\mu\}$ is selected, resulting in that exactly one non accepting computation is selected and that exactly one $\ddot{\beta}_{i},i\in\{1,\ldots,\mu\}$ is selected.
-With this, we get that $\eta_{i},i\in\{1,\ldots,\mu\}$ matches $u$ and thus, that $z_{i}=0$ and $\hat{z}_{i_{1}}=\#$.
-We use these two variables to encode the 0 and the #\# in the non accepting computation encodings.
-If we can show that this selecting process also works in the setting $R=r_{ab}$ and that we get two variables with which we can encode the letters 0 and #\#, then this proof also works in this setting.
-For two variables $x,y$, we have if $x=0$ (or x=#x=\#) and $(x,y)\in r_{ab}$, then $y=0$ (or y=#y=\# respectively).
-Thus, we can model with the relation $r_{ab}$ identical variables if we know a one letter substitution of one variable.
-[…]
-$x_{i}$ $x_{i^{\prime}}$ works?
-0 0 $00\neq 0\#$
-0#0\# 0#0\# or #0\#0 $0#$, $0\#,\ 0\#^{2}0\neq 0\#^{3}$
-$0\#^{2}$ $0\#^{2}$ or #0#\#0\# or ##0\#\#0 $0\#^{2}0\#^{2},\ 0\#^{3}0\#,\ 0\#^{4}0\neq 0\#^{3}00$
-$0\#^{3}$ $0\#^{3}$ or $0\#^{2}$ or $^{2}0\#$ or $^{3}0$ $0\#^{3}0\#^{3},\ 0\#^{3}\#0\#^{2},\ 0\#^{5}0\#,\ 0\#^{6}30\neq 0\#^{3}00\#^{2}$
-$0\#^{3}0$ $0\#^{3}0$ $0\#^{3}00\#^{3}0=v\ v$
-Table 1: Possibilities for the first match with vvv\ v
-Thus, the only possibility is that for exactly one $i\in\{1,\ldots,\mu\}$ the variables $x_{i}$ and $x_{i^{\prime}}$ get substituted with $v$ and all other variables $x_{j},x_{j^{\prime}}$ for $j\neq i,j\in\{1,\ldots,\mu\}$ get substituted with $\epsilon$.
-This together with the matching from vuvv\ u\ v to $\ddot{\beta}_{1}\ldots\ddot{\beta}_{\mu}$ implies that $\eta_{i}$ needs to match $u$.
-For this matching between $\eta_{i}=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ and $u=0\#^{5}0$, we get the following possibilities.
-Remember that $(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in R$.
-$z_{i}$ $\hat{z}_{i_{1}}$ works?
-$\epsilon$ 0 or #\# $0^{5},\ \#^{5}\neq 0\#^{5}0$
-0 #\# $0\#^{5}0=0\#^{5}0=u$
-0#0\# #\# $0\#\#^{5}\#0\neq 0\#^{5}0$
-0##0\#\# #\# $0\#^{2}\#^{5}\#^{2}0\neq 0\#^{5}0$
-0###0\#\#\# $\epsilon$ $0\#^{6}0\neq 0\#^{5}0$
-Table 2: Possibilities for the match with $u$
-Thus, the only possibility is that $z_{i}$ and $z_{i^{\prime}}$ get substituted with 0 and $\hat{z}_{i_{k}},k\in\{1,2,3,4,5\}$ get substituted with #\#.
-With this, we can encode the 0 and the #\# and thus, the proof of Theorem˜4.1 works in this setting analogously.
+0.A.2 Proof of Theorem˜3.1
+Proof.
+We obtain the first direction immediately from Lemma˜1.
+For the other direction, assume $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
+By Lemma˜2 we know $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$ and $L_{NE}(\beta,r_{\beta})=L_{NE}(\beta,[r_{\beta}])$.
+As $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$, we trivially have $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$.
+Thus, $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
 ∎
-This concludes the results regarding the inclusion problem for relational pattern languages covered in this paper.
-4.3 Discussion and Challenges Regarding the Other Relations
-Currently, we know that for relational patterns under the relations $r_{=}$, $r_{ssq}$, $_{ab}, rr$$_{rev}$, and $r_{*}$ the inclusion problem is undecidable.
-This mainly comes from the fact that either the relations can be used to simulate the behavior of $r_{=}$ generally (as in the case of $r_{ssq}$ and $r_{*}$) or it comes from the fact that essential properties used in relevant reductions can also be emulated using other relations (enforce equal letters or equal unary words using $r_{ab}$ or $r_{rev}$).
-When it comes to the remaining relations $r_{|w|}$, $r_{perm}$, $r_{com^{*}}$, and $r_{com^{+}}$, we see that these properties are not trivially obtainable.
-The relations $r_{|w|}$ and $r_{perm}$ both do not force equal letters or equal unary words in alphabets that are at least of binary size.
-The relations $r_{com^{*}}$ and $r_{com^{+}}$ do not even force equal length between variables in valid substitutions, making, for example, the comparison of counter values impossible in the way it has been done in the previous reductions.
-Hence, it remains an worthwhile open question whether the inclusion problem becomes decidable at some point, for very structurally weak relations.
-Question 1
-Let $R=\{\tilde{r}\}$ for some relation $\tilde{r}\in\{r_{|w|},r_{perm},r_{com^{*}},r_{com^{+}}\}$.
-Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
-Is the inclusion problem $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$, generally decidable?
-5 The Equivalence Problem
-In this section, we briefly discuss some results regarding the equivalence problem for non-erasing relational pattern languages.
-As we will see, we will obtain a series of small results that builds up to the main theorem of this section, stating that the equivalence problem for all non-erasing pattern languages over a binary equivalence relation that is antisymmetric on elements of length 1 (i.e., has equality for single letters) is decidable in polynomial time (covering the relations $\{r_{=},r_{ab},r_{con^{+}}\}$ from this paper’s discussion).
-We define $r_{\alpha}]$ to be the transitive, reflexive and symmetric closure on some set of binary related variables $r_{\alpha}$.
-Additionally, for this section, we assume that variables are always introduced inside a pattern with increasing indices starting from $1$ (i.e., $x_{1}$, $x_{2}$, and so on).
-Due to space constraints, the technical proofs of almost all steps can be found in Appendix 0.C, while the steps themselves are supposed to give some intuition on why the main result of this section, Theorem 5.1, holds.
-We begin with a result regarding $r_{\alpha}]$ in the case of $r_{\alpha}$ being an equivalence relation.
-Lemma 3
-Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation and let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$. Then we have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$.
-Proof
-($\Rightarrow$): Let $w\in L_{NE}(\alpha,r_{\alpha})$.
-Then there exists $h\in H$ such that $h((\alpha,r_{\alpha}))=w$.
-Let $(\tilde{r},x,y)\in r_{\alpha}$ for $x,y\in\operatorname{var}(\alpha)$.
-Then $(h(x),h(y))\in\tilde{r}$ by definition.
-As $\tilde{r}$ is an equivalence relation, we also have $(h(y),h(x))\in\tilde{r}$, $(h(x),h(x))\in\tilde{r}$, and if there exists $(\tilde{r},y,z)\in r_{\alpha}$ such that $(h(y),h(z))\in\tilde{r}$ then also $(h(x),h(z))\in\tilde{r}$.
-So, $h$ satisfied all constraints set by $r_{\alpha}]$.
-Hence, $w\in L_{NE}(\alpha,[r_{\alpha}])$.
-($\Leftarrow$): Let $w\in L_{NE}(\alpha,[r_{\alpha}])$ and let $h\in H$ such that $h((\alpha,[r_{\alpha}]))=w$.
-As $r_{\alpha}\subseteq[r_{\alpha}]$, all constraints in $r_{\alpha}$ must be satisfied.
-Thus, we also have $w\in L_{NE}(\alpha,r_{\alpha})$.
+0.A.3 Proof of Corollary˜1
+Proof.
+Since every relation in $r_{=},r_{ab},r_{con^{+}}$} is an equivalence relation enforcing equality of related letters, we first normalize $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ by renaming variables in order of first occurrence to $x_{1},x_{2},\dots$.
+We then test whether $\alpha=\beta$.
+If not, we return false.
+Otherwise, we compute $r_{\alpha}]$ and $r_{\beta}]$ via their reflexive, symmetric, and transitive closures by a single scan of the patterns.
+These classes have total size at most the number of variables, so this also takes linear time.
+We return true exactly if $r_{\alpha}]=[r_{\beta}]$, and false otherwise, by Theorem˜3.1.
 ∎
-Following on that, we can also obtain the following relationship.
-Lemma 4(✂)
-Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being some equivalence relation and assume $|\Sigma|\geq 2$.
-Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
-If $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$, then $\alpha=\beta$.
-Using the previous lemma, we can show the following result.
-Lemma 5(✂)
-Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation that is antisymmetric on words of length 1 (i.e., equality for letters) and assume $|\Sigma|\geq 2$.
-Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two relational patterns.
-Then $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$ if and only if $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
-Finally, taking all previous lemmas into account, we obtain the main result of this section.
-Theorem 5.1(✂)
-Let $R:=\{\tilde{r}\}$ with $\tilde{r}$ being an equivalence relation that is antisymmetric on words of length 1 (i.e., equality for letters) and assume $|\Sigma|\geq 2$.
-Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two relational patterns.
-Then $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ if and only if $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
-Hence, for the considered relations in this paper, we obtain the following.
-Corollary 3(✂)
-Let $R=\{\tilde{r}\}$ for some $\tilde{r}\in\{r_{=},r_{ab},r_{con^{+}}\}$.
-Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, we can check in polynomial time whether $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
-6 Conclusion
-We considered the three main decision problems membership, inclusion, and equivalence for relational pattern languages under various typical binary relations, i.e., $r_{=}$, $r_{ssq}$, $r_{ab}$, $_{perm}, rr$$_{rev}$, $r_{com^{+}}$, $r_{com^{*}}$, and $r_{*}$.
-In particular, for the membership problem, we observed that in all considered cases, that is, for all relations in $\mathcal{R}$ and for erasing as well as non-erasing relational pattern languages, the problem turns out to be NP-complete, even for very simple relations such as $r_{|w|}$.
-As the reduction relies on rather simple properties to hold between related areas inside the constructed patterns, we expect the problem to be NP-hard for generally almost all typical relations that could also be considered.
-This allows for the assumption that searching for words in sets of texts that are compactly described by, e.g., a relational pattern of almost any kind, is generally a computationally hard problem.
-For the inclusion problem, we observed that many cases remain undecidable.
-In the erasing as well as the non-erasing cases, the problem remains undecidable for the structurally less restrictive relations $r_{ssq}$, $r_{ab}$, and $r_{*}$ as well as for the relation $r_{rev}$.
-In the cases of $r_{ssq}$ and $r_{*}$, we can abuse the property that relating two variables in both directions results basically in $r_{=}$.
-For the other cases, we observe that the key for the reductions to work is, first, finding ways to force certain variables to be substituted by equal single letters to realize structural properties and, second, to force equality of longer unary words to realize the comparison of counter values.
-It would be interesting to know whether similar results can be obtained using the other relations that enforce even less structure in valid substitutions.
-Regarding the equivalence problem, we obtained decidability in polynomial time for $r_{ab}$ and $r_{com^{+}}$ in the non-erasing case, similar to $r_{=}$.
-We conjecture that this should hold for many more relations in the non-erasing case by checking some set of minimal length substitutions.
-For the erasing cases, everything except $r_{|w|}$ remains open and we expect the decidability question to be generally rather difficult to answer, as it is the main open question for $r_{=}$.
-This concludes this paper. See Table 3 for a visual summary of the results.
-$r_{=}$ $r_{|w|}$ $r_{ssq}$ $r_{ab}$ $_{perm} rr$$_{rev}$ $r_{com^{*}}$ $_{com^{+}} rr_{*}$
- E ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC
- E ($\subseteq$)  UD  P UD UD ? UD ? ? UD
- E (==) ?  P ? ? ? ? ? ? ?
- NE ($\in$)  NPC NPC NPC NPC NPC NPC NPC NPC NPC
- NE ($\subseteq$)  UD  P UD UD ? UD ? ? UD
- NE (==)  P  P ? P ?  D ? P ?
-Table 3: Current state of decision problems for relational pattern languages for various constraints (NPC = NP-Complete, UD = Undecidable, D = Decidable, E = Erasing, NE = Non-Erasing, ($\in$) membership problem, ($\subseteq$) inclusion problem, (==) equivalence problem).
-Red text marks the results covered in this paper.
-Appendix 0.A Additional Content for Membership Problem
-Proof of Proposition˜1
-Proof
-Assume $|\Sigma|\geq 2$.
-Let $r\in\mathcal{R}$, $R=\{r\}$, $w\in\Sigma^{*}$, and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$.
-We can guess a substitution $h\in H$ which results in $|h(\alpha)|=|w|$, then check whether $(x,y)\in r$ for all $(r,x,y)\in r_{\alpha}$, and finally, if that is the case, return true if $h(\alpha)=w$ and return false in any other case.
-As all relations $r\in\mathcal{R}$ are polynomial-time decidable relations, this results in a polynomial runtime of the verifier, hence resulting for the problem to be in NP.
+0.A.4 Proof of Theorem˜3.2
+Proof.
+Again, consider $(\alpha,r_{\alpha})$ and for each subset $S\subseteq\mathcal{C}_{\alpha}$, replace the variables $v\in\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by any word in $\Sigma^{+}$ and any other variable $v\notin\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by $\varepsilon$.
+We call the resulting language $L_{E}(\alpha,r_{\alpha},S)$.
+Using that, we obtain $L_{E}(\alpha,r_{\alpha})=\bigcup_{S\subseteq\mathcal{C}_{\alpha}}L_{E}(\alpha,r_{\alpha},S)$.
+Do the same for $\beta$.
+Note there exist at most $2^{k_{\alpha}}\leq 2^{|\alpha|}$ subsets $S$ for $\alpha$ and at most $2^{k_{\beta}}\leq 2^{|\beta|}$ subsets for $\beta$.
+Here, we provide a more detailed construction of the previously mentioned NFAs: Starting with $(\alpha,r_{\alpha})$, for each such $S$ defined above, we can take a sub-pattern $\alpha_{S}$ of $\alpha$ that only contains the variables that are not substituted by the empty word.
+As $r_{\varepsilon_{=}}$ imposes no structural restriction other than being substituted (non)emptily commonly, essentially $\alpha_{S}$ is a regular pattern and we obtain the relationship $L_{NE}(\alpha_{S})=L_{NE}(\alpha,r_{\alpha},S)$.
+So, for each such regular sub-pattern $\alpha_{S}$, we can create an NFA $A_{\alpha_{S}}=(Q,\Sigma,\delta,q_{0}),F$ with $L(A_{\alpha_{S}})=L_{NE}(\alpha_{S})$ using a straightforward construction that represents that regular pattern.
+Consider the following sketch.
+We set $Q=\{q_{\alpha_{S},0},...,q_{\alpha_{S},|\alpha_{S}|}\}$ where each state $q_{\alpha_{S},i}$ represents a position between some letters in the pattern $\alpha_{S}$ (or a position in front of the first letter or after the last letter).
+Then we set the state corresponding to the last position in $\alpha_{S}$ (i.e., $q_{\alpha_{S},|\alpha_{S}|}$) to the only final state.
+Hence, $F=\{q_{\alpha_{S},|\alpha_{S}|}\}$.
+For all other positions $i<|\alpha_{S}|$, if $\alpha_{S}[i+1]$ is some terminal letter $\mathtt{a}\in\Sigma$, we add only one outgoing transition and set $\delta(q_{\alpha_{S},i},\mathtt{a})=\{q_{\alpha_{S},i+1}\}$.
+On the other hand, if $\alpha_{S}[i+1]\in\operatorname{var}(\alpha_{S})$ is some variable, add two outgoing transitions for each letter, one going to the state representing the next position, the other one staying in the same state.
+Thus, set $\delta(q_{\alpha_{S},i},\mathtt{a})=\{q_{\alpha_{S},i},q_{\alpha_{S},i+1}\}$, for all $\mathtt{a}\in\Sigma$.
+As $\alpha_{S}$ is a regular pattern as established before, the relationship $L(A_{\alpha_{S}})=L_{NE}(\alpha_{S})$ is apparent.
+Now, for each $S\subseteq\mathcal{C}_{\alpha}$, we can construct such an NFA $A_{\alpha_{S}}$.
+Finally, we can combine all of these NFA into one common NFA $A_{\alpha}$ where a new initial state is introduced that is connected to each initial state of the sub-automata $A_{\alpha_{S}}$ by an $\varepsilon$-transition, and each final state is preserved.
+So, using one outgoing transition of the initial state of $A_{\alpha}$, we choose which blocks of variables $S$ are to be substituted emptily together and which are not.
+Then, the respective sub-automaton $A_{\alpha_{S}}$ produces some word of the corresponding sub-language $L_{E}{\alpha,r_{\alpha_{S}}}$.
+In total, we see that the NFA $A_{\alpha}$ can be constructed and represents $L(A_{\alpha})=L_{E}{\alpha,r_{\alpha}}$.
+So the same for $(\beta,r_{\beta})$ by constructing an NFA $A_{\beta}$ analogously.
+It is well known that checking equivalence of the languages of two NFA is PSPACE-complete (see, e.g., [1]).
+Hence, we can check $L(A_{\alpha})=L(A_{\beta})$ and use the result to decide $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
 ∎
-Proof of Lemma˜1:
-Proof
-By the result from Angluin in [Angluin1980], we know that this statement already holds for the case of $r_{=}$.
-This is shown using a reduction from the $3\mathtt{-SAT}$ problem.
-We quickly recall the construction used in [Angluin1980] and refer to the original source for the proof of its correctness.
-Notice that the pattern constructed here is not a relational pattern by our definition but a pattern using the usual notation allowing for multiple occurrences of the same variable.
-We will comment on the adaptions needed for the notation of relational pattern languages after giving the construction.
-Let $\varphi$ be a $3\mathtt{-SAT}$ formula.
-Let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the set of the used variables in $\varphi$ for some $m\in\mathbb{N}$.
-Let $C=\{C_{1},...,C_{n}\}$ be the set of all clauses in $\varphi$ that are in conjunction to each other for some $n\in\mathbb{N}$.
-For each clause $C_{i}\in C$, refer to its three variables by $X_{i,1}$, $X_{i,2}$, and $X_{i,3}$.
-Define a function $f:\mathcal{X}\rightarrow V$ by $f(X_{i})=v_{i}$ and $f(\overline{X_{i}})=u_{i}$ for new independent pattern variables $u_{i},v_{i}\in V$ with $i\in[m]$.
-Now, construct the word
-$w=\#\ s_{1}\ \#\ ...\ \#\ s_{m}\ \#\ t_{1}\ \#\ ...\ \#\ t_{n}\ \#\ w_{1}\ \#\ ...\ \#\ w_{n}\ \#$
-with $s_{i}=1^{3}$, $t_{j}=1^{7}$, and $w_{j}=1^{4}$ for $i\in[m]$ and $j\in[n]$ as well as a pattern
-$\alpha=\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\ r_{1}\ \#\ ...\ \#\ r_{n}\ \#$
-with $p_{i}=u_{i}v_{i}$, $q_{j}=f(X_{j,1})f(X_{j,2})f(x_{j,3})z_{j}$, and $r_{j}=z_{j}d_{j}$ for new independent pattern variables $z_{j},d_{j}\in X$ with $j\in[n]$.
-Notice that the construction uses three different elements.
-The first element contains blocks $u_{i}v_{i}\#$ for $i\in[m]$ which have to be non-emptily substituted to $1^{3}\#$, resulting in one of the variables to be substituted to $11$ and the other one to be substituted to $1$, representing either $X_{i}$ or $\overline{X_{i}}$ to be mapped to true in $\varphi$, having the other one false.
-The middle part $f(X_{i,1})f(X_{i,2})f(X_{i,3})z_{i}\#$ needs to be non-emptily mapped to $1^{7}$.
-Notice that due to the last part $z_{i}d_{i}\#$, which is to be substituted to $1^{4}\#$, the variable $z_{i}$ may only be substituted tp $1$, 1111 or $111$, as $d_{i}$ has do be substituted at least to a word of length $1$, i.e., the word $1$.
-Because of that, at least one of $f(X_{i,j})$ needs to be mapped to $11$ instead of $1$ to obtain $1^{7}\#$, representing at least one variable in each clause to be mapped to true.
-As this holds for all clauses, deciding whether $w\in L_{NE}(\alpha)$ also decides the satisfiability of the formula $\varphi$.
-Consider [Angluin1980] for the formal proof.
-To obtain this construction in the notation of relational pattern languages, instead of introducing each variable $k\in\mathbb{N}$ times, introduce $k$ different variables in a pattern $\alpha^{\prime}$ that has exactly the same terminal letters and relate those $k$ variables with equality $r_{=}$ in the relation set $r_{\alpha^{\prime}}$.
-This results in a relational pattern $(\alpha^{\prime},r_{\alpha})$ with $L_{NE}(\alpha)=L_{NE}(\alpha^{\prime},r_{\alpha^{\prime}})$ for which deciding the membership $w\in L_{NE}(\alpha,r_{\alpha})$ also decides the satisfiablity of $\varphi$.
-Now we give arguments why the exact same construction works for the other relations in $\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$.
-  1. 1.
-    If $R=\{r_{|w|}\}$, then variables have to be substituted with equal length.
-As any substitution resulting in $w$ only uses unary words for the variables, having equal length also means being the equal word.
-This concludes this case.
-  2. 2.
-    If $R=\{r_{ssq}\}$, then relating two variables $x$ and $y$ in both directions, i.e., putting $(r_{ssq},x,y)\in r_{\alpha^{\prime}}$ as well as $(r_{ssq},y,x)\in r_{\alpha^{\prime}}$ results in some substitution $h\in H$ to be valid if and only if $h(x)=h(y)$, as otherwise one can’t be a subsequence of the other.
-That concludes this case.
-  3. 3.
-    If $R=\{r_{ab}\}$, then valid substitutions of related variables result in them to have equal length, as abelian equivalence implies that.
-Hence, this case follows as in the case of the equal length relation $r_{|w|}$.
-  4. 4.
-    If $R=\{r_{perm}\}$, then we also get that two related variables must be substituted with equal length.
-As all substitutions must result in unary words consisting only of the letter $1$, this case also follows directly as the case of the equal length relation.
-  5. 5.
-    Also in the case of $R=\{r_{rev}\}$ we can use the argument that only unary substitutions are relevant in this reduction.
-As the reversal of a unary word is again the same unary word, this case also follows immediately.
-  6. 6.
-    Finally, in the case of $R=\{r_{*}\}$, as we said before, we may use this relation in both directions for each pair of variables, resulting effectively in equality of variables as in the case of the subsequence relation $r_{ssq}$.
-This concludes all cases.
-For larger alphabets, the same construction may be used by trivially extending the word and the pattern by terminal letters made of up all additional letters in the alphabet.
+0.A.5 Proof of Lemma˜3
+Proof.
+(1): Let $\alpha=x_{1}x_{2}\dots x_{n_{\alpha}}$ and $\beta=y_{1}y_{2},\dots y_{n_{\beta}}$ be terminal-free and let $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\beta}$ be the equivalence classes.
+Note that the empty word can be obtained by both patterns.
+Thus, in the following, we focus on non-empty words.
+Consider the pattern $\alpha$.
+Let $S\subseteq\mathcal{C}_{\alpha}$ with $S\neq\emptyset$ be the subset of equivalence classes chosen to be non-empty.
+Assume that all variables in the remaining classes are replaced by $\varepsilon$.
+Since each non-empty variable contributes at least one letter to the resulting word but has no other restriction, we can build all words of length at least $\sum_{C_{\alpha,i}\in S}|C_{\alpha,i}|$.
+Now, when taking all possible subsets $S$ into account, no word $w$ of positive length below $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|$ is possible.
+Thus, the language of accepted words by $\alpha$ is
+$L_{E}(\alpha,r_{\alpha})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}.$
+Respectively, the language of accepted words by $\beta$ is
+$L_{E}(\beta,r_{\beta})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
+Now, $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|=\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$ if and only if both patterns are equivalent.
+This property can be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
+(2): Assume $|\Sigma|=1$ and $a\in\Sigma$.
+Note that the patterns $\alpha$ and $\beta$ must have the same number of terminals in order to be equivalent.
+Otherwise we could set all variables to $\varepsilon$ and obtain distinct words.
+Let $\ell$ be the number of terminals in each pattern.
+By the same arguments as in (1), either all variables are empty, i.e., $\{a^{\ell}\}\in L_{E}(\alpha,r_{\alpha})$ and $\{a^{\ell}\}\in L_{E}(\beta,r_{\beta})$, or a non-empty subset $S$ of equivalence classes contributes $\sum_{C_{\alpha,i}\in S}|C_{\alpha,i}|$ (respectively $\sum_{C_{\beta,i}\in S}|C_{\beta,i}|$) $a$’s to the words.
+As we consider a single-letter alphabet, the language of accepted words by $\alpha$ is
+$L_{E}(\alpha,r_{\alpha})=\{a^{\ell}\}\cup\{w\mid|w|\geq\ell+\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}.$
+Respectively, the language of accepted words by $\beta$ is
+$L_{E}(\beta,r_{\beta})=\{a^{\ell}\}\cup\{w\mid|w|\geq\ell+\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
+We obtain that both patterns are equivalent if and only if they have the same number of terminals and $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|=\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$.
+This can again be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
+(3): Assume there exists a letter $c\in\Sigma$ that does not appear as a terminal in either $\alpha$ or $\beta$.
+If all variables are set to $\varepsilon$, both patterns generate a word that matches their terminals.
+Hence, a necessary condition for equivalence is that the terminal words of $\alpha$ and $\beta$ are identical.
+This, we assume in the following and focus on words where at least one variable is non-empty.
+Let $m$ be the number of terminals in $\alpha$ or $\beta$. We write
+$\alpha=V_{\alpha,0}T_{1}V_{\alpha,1}T_{2}\dots T_{m}V_{\alpha,m},$
+where $T_{j},j\in\{1,\dots,m\}$ are blocks terminals and $V_{\alpha,j},j\in\{0,\dots,m\}$ are blocks of variables.
+Since the terminals of both patterns are identical, $\beta$ has the same decomposition but the blocks of variables may be different, i.e.,
+$\beta=V_{\beta,0}T_{1}V_{\beta,1}T_{2}\dots T_{m}V_{\beta,m}.$
+First, consider the pattern $\alpha$. For each equivalence class $C_{\alpha,i}$, let $\mathbf{occ}_{\alpha,i}(j)$ be the number of occurrences of that class in block $V_{\alpha,j}$ and define
+$\mathbf{occ}_{\alpha,i}:=(\mathbf{occ}_{\alpha,i}(0),\mathbf{occ}_{\alpha,i}(1),\dots,\mathbf{occ}_{\alpha,i}(m)).$
+If a subset $S\subseteq\mathcal{C}_{\alpha}$ is non-empty, the minimal number of letters in block $V_{j}$ is $\sum_{C_{\alpha,i}\in S}\mathbf{occ}_{\alpha,i}(j)$.
+Thus, every word of length at least $\sum_{C_{\alpha,i}\in S}\mathbf{occ}_{\alpha,i}(j)$ can be generated in that block.
+We say that an equivalence class $C_{\alpha,i}$ is dominated if there exists a set $S^{\prime}$ of other classes, i.e., $S^{\prime}\subset\mathcal{C}_{\alpha}$ and $C_{\alpha,i}\notin S^{\prime}$, such that for all blocks $j\in\{0,\dots,m\}$ we have
+$\mathbf{occ}_{\alpha,i}(j)=0\Longleftrightarrow\sum_{C_{\alpha,i^{\prime}}\in S^{\prime}}\mathbf{occ}_{\alpha,i^{\prime}}(j)=0$
+and
+$\mathbf{occ}_{\alpha,i}(j)\geq\sum_{C_{\alpha,i^{\prime}}\in S^{\prime}}\mathbf{occ}_{\alpha,i^{\prime}}(j).$
+Then all words generated in each block individually by $C_{\alpha,i}$ can also be generated by the variables in $S^{\prime}$ and we may remove all variables of that equivalence class without changing the language.
+We may consider each block individually since $c$ does not occur as a terminal, we may assign the letter $c$ to all variable occurrences.
+By iteratively removing dominated classes we obtain a minimal pattern of $\alpha$.
+For a removed class $i$, we set $\mathbf{occ}_{\alpha,i}(j)=0$ for all $j\in\{0,\dots,m\}$.
+We do the same reduction step for $\beta$.
+Now, the languages are equal if and only if the minimal patterns are identical, i.e., for all $i\in\{1,\dots,k_{\alpha}\}$ there exists $i^{\prime}\in\{1,\dots,k_{\beta}\}$ such that $\mathbf{occ}_{\alpha,i}=\mathbf{occ}_{\beta,i^{\prime}}$.
+We now argue that this condition is correct.
+Suppose the minimal patterns differ.
+Then there exists a class $C$ for one pattern that has no matching class in the other pattern but is also not dominated.
+Because $c$ is unused as a terminal, construct a word that assigns all variables of $C$ the letter $c$ and sets all other variables to $\varepsilon$.
+Due to the distinct number of occurrences, this gives in at least one block a word length that can not be matched by the other pattern.
+Thus, the languages differ.
+The block decomposition and counting the occurrences both require linear time.
+Testing the dominance between classes takes at most $O(k_{\alpha}^{2}m)$, and respectively $O(k_{\beta}^{2}m)$ comparisons.
+Comparing the minimal patterns is also possible in linear time Hence, the complete procedure runs in polynomial time.
 ∎
-Proof of Proposition˜2:
-Proof
-We show this, also, by reducing from the $3\mathtt{-SAT}$ problem.
-So, as before, let $\varphi$ be a $3\mathtt{-SAT}$ formula and let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the set of the used variables in $\varphi$ for some $m\in\mathbb{N}$.
-Let $C=\{C_{1},...,C_{n}\}$ be the set of all clauses in $\varphi$ that are in conjunction to each other for some $n\in\mathbb{N}$.
-For each clause $C_{j}\in C$ for $j\in[n]$, refer to its three variables by $X_{j,1}$, $X_{j,2}$, and $X_{j,3}$.
-W.l.o.g.
-assume that no variable occurs two times in the same clause (a variable and its reversal, however, may still occur together in one clause).
-Define $n+1$ functions $f_{j^{\prime}}$ with $j_{j^{\prime}}\in[n]_{0}$ such that $f_{j^{\prime}}:\mathcal{X}\rightarrow X$ with $f_{j^{\prime}}(X_{i})=u_{i,j^{\prime}}$ and $f_{j^{\prime}}(\overline{X_{i}})=v_{i,j^{\prime}}$.
-Assume $R=\{r\}$ for $r\in\{r_{com^{+}},r_{com^{*}}\}$.
-Construct the word
-$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
-with
-$s_{i}$ $=1\#1\text{, and}$
-$t_{j}$ $=1^{10}\#1^{10}\#1^{10}$
-for $i\in[m]$ and $j\in[n]$ as well as the relational pattern $(\alpha,r_{\alpha})$ such that $\alpha$
-$\alpha=\#\#\ \alpha_{1,1}\ \#\#\ ...\ \#\#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\#\ ...\ \#\#\ \alpha_{2,n}\ \#\#$
-for $i\in[m]$ and $j\in[n]$ with
-$\alpha_{1,i}$ $=f_{0}(X_{i})f_{0}(\overline{X_{i}}),\text{ and}$
-$\alpha_{2,j}$ $=z_{j,1}\ f_{j}(X_{j,1})\ z_{j,2}\ f_{j}(X_{j,2})\ z_{j,3}\ f_{j}(X_{j,3})\ z_{j,4},$
-and the set of related variables $r_{\alpha}=\{\ (r,u_{i,j},u_{i,j^{\prime}}),(r,v_{i,j},v_{i,j^{\prime}})\ |\ i\in[m],j,j^{\prime}\in[n+1]\ \}$.
-Hence, notice that two variables with the same base name $u_{i,j}$ and $u_{i^{\prime},j^{\prime}}$ (resp.
-$v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$) are related to each other if and only if $i=i^{\prime}$.
-Also notice, that each variable $z_{i,j}$ is independent and may be substituted to anything.
-Now, assume there exists an assignment of the variables in $\mathcal{X}$ such that $\varphi$ is satisfied.
-We will now construct a $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$.
-For each $i\in[m]$, if $X_{i}$ is true (hence, $\overline{X_{i}}$ is false), set $h(u_{i,j})=1$ and $h(v_{i,j})=\#1$, otherwise set $h(u_{i,j})=1\#$ and $h(v_{i,j}=1)$ for $j\in[n+1]$.
-This results in $h(\alpha_{1,i})=1\#1$ for all $i\in[m]$.
-As there is at least one variable assigned to true for each clause $j\in[n]$, we know that $|h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))|_{\#}\leq 2$.
-As there is at most two $1$’s between two occurrences of #\# and at most $3$ occurrences of $1$’s before or after the first or last occurrence of #\# in $h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))$, we always find a non-empty substitution of all variables $z_{j,1}$ to $z_{j,4}$ such that $h(\alpha_{2,j})=\#\#\ 1^{10}\#1^{10}\#1^{10}$.
-All substitutions done by $h$ result in related variables to be substituted equally, hence, this satisfies the relation $r$, and by that there exists some $r_{\alpha}$ valid $h\in H$ such that $h(\alpha)=w$.
-Now assume $w\in L_{NE}(\alpha,r_{\alpha})$.
-Then there exists some $r_{\alpha}$ valid substitution $h\in H$ such that $h(\alpha)=w$.
-We know that each occurrence of the factor ##\#\# in $w$ must be mapped by one occurrence of the factor ##\#\# in $h(\alpha)$.
-As the same number of ##\#\# occur as terminal letters in $\alpha$ as there are occurrences of it in $w$, we know that these occurrences must align.
-Hence, we get that for all $i\in[m]$ and $j\in[n]$ we have
-$h(\alpha_{1,i})=h(u_{i,1}v_{i,1})=1\#1=s_{i}$
-and
-$h(\alpha_{2,j})=h(z_{j,1}\ f_{j}(X_{j,1})\ z_{j,2}\ f_{j}(X_{j,2})\ z_{i,3}\ f_{j}(X_{j,3})\ z_{j,4})=1^{10}\#1^{10}\#1^{10}=t_{j}.$
-By the first equality, we obtain that either $h(u_{i,1})=1$ and $h(v_{i,1})=\#1$ or $h(u_{i,1})=1\#$ and $h(v_{i,1})=1$.
-Notice by the definition of $r$, in this case, that for all $i\in[m]$ and $j\in[n]$ we have that $h(u_{i,j})\in\{1\}^{*}$ if and only if $h(u_{i,0})=1$.
-The same holds analogously for all $v_{i,j}$.
-Now select an assignment for the variables in $\mathcal{X}$ for the $3\mathtt{-SAT}$ formula by setting $X_{i}=true$ (hence $\overline{X_{i}}=false$) iff $h(u_{i,1})=1$ and otherwise $\overline{X_{i}}=true$ (hence $X_{i}=false$) iff $h(v_{i,1})=1$.
-By the argument from before, this is a valid assignment of the variables.
-If there were only falsely assigned variables in a single clause, using the previous observations, there existed $j\in[n]$ such that $|h(\alpha_{2,j})|_{\#}\geq 3$, which is not the case.
-Hence, there is at least one variable assigned to true per clause and by that we obtain that $\varphi$ must be satisfiable.
+0.A.6 Proof of Proposition˜1
+Proof.
+For $r\in\{r_{*},r_{ssq}\}$, reduce from equivalence of erasing pattern languages over $r_{=}$ as follows.
+Given $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r_{=}\}$, construct $(\alpha,r^{\prime}_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r\}$ by replacing every $(r_{=},x_{i},x_{j})\in r_{\alpha}$ with both $(r,x_{i},x_{j})$ and $(r,x_{j},x_{i})$.
+Since symmetric $r_{*}$-constraints and symmetric $r_{ssq}$-constraints force equality of substituted words, this preserves exactly the valid substitutions.
+Hence the equivalence problem for erasing pattern languages reduces directly to equivalence for relational patterns over $r_{*}$ and $r_{ssq}$.
+For $r=r_{rev}$, let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r_{=}\}$. Define
+$\overset{\leftrightarrow}{\alpha}=\alpha\overset{\leftarrow}{\alpha},$
+where $\overset{\leftarrow}{\alpha}$ is the reversal of $\alpha$ in which each variable $x_{i}$ is replaced by a fresh variable $x_{i^{\prime}}$.
+Let $r_{\overset{\leftrightarrow}{\alpha}}$ contain $(r_{rev},x_{i},x_{i^{\prime}})$ for every $x_{i}\in\operatorname{var}(\alpha)$ and, for every $(r_{=},x_{i},x_{j})\in r_{\alpha}$, also $(r_{rev},x_{i},x_{j^{\prime}})$ and $(r_{rev},x_{j},x_{i^{\prime}})$.
+Then every valid substitution satisfies $h(x_{i^{\prime}})=h(x_{i})^{R}$, and the additional constraints imply $h(x_{i})=h(x_{j})$ iff $(r_{=},x_{i},x_{j})\in r_{\alpha}$.
+Consequently, $L_{E}(\overset{\leftrightarrow}{\alpha},r_{\overset{\leftrightarrow}{\alpha}})=\{ww^{R}\mid w\in L_{E}(\alpha,r_{\alpha})\}.$ Thus, for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ over $r_{=}$, $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ if and only if $L_{E}(\overset{\leftrightarrow}{\alpha},r_{\overset{\leftrightarrow}{\alpha}})=L_{E}(\overset{\leftrightarrow}{\beta},r_{\overset{\leftrightarrow}{\beta}}).$ Hence equivalence of erasing pattern languages reduces to the equivalence problem for relational patterns over each $r\in\{r_{ssq},r_{*},r_{rev}\}$.
 ∎
-Proof of Lemma˜2:
-Proof
-Again, the original proof in [Jiang1994] for the case of equality between variables uses a reduction from $3\mathtt{-SAT}$ to the membership problem.
-The construction is essentially the same as in [Angluin1980] just with different sizes of each part of the word $w$ to construct.
-Here, we have for all $i\in[m]$ and $j\in[n]$ that $s_{i}=1$, $t_{j}=1^{7}$, and $w_{j}=1^{2}$.
-By that, as variables are allowed to be substituted by the empty word, the reduction works in exactly the same way.
-Consider [Angluin1980, Jiang1994] for the formal proof of this construction.
-As the reasoning about the construction works exactly in the same way as before, and as the considered relations don’t change in their general behavior if the empty word is allowed, we can use the same arguments as in Lemma 1 and obtain NP-hardness for all these cases here as well.
-∎
-Proof of Proposition˜3:
-Proof
-Consider the construction from Proposition 2.
-As mentioned before, because of the definition of $r_{com^{+}}$, two related variables may never be substituted to the empty word, as $\varepsilon\notin\{y\}^{+}$ for all $y\in\Sigma^{*}$.
-As only the variables with the base name $z$ are independent, only those may be substituted emptily.
-Those are only used to fill up remaining #\#’s or $1$’s in the encoded clauses in the case of valid assignments for the $3\mathtt{-SAT}$ formula $\varphi$.
-Thus, the main argument stays the same and can be applied here in exactly the same way.
-∎
-Proof of Proposition˜4:
-Proof
-Again, we reduce from $3\mathtt{-SAT}$.
-Let $\varphi$ be a $3\mathtt{-SAT}$ formula, let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the variables in $\varphi$ for some $m\in\mathbb{N}$, let $C=\{C_{1},...,C_{n}\}$ be the set of clauses in $\varphi$ for some $n\in\mathbb{N}$.
-For each clause $C_{j}\in C$, refer to its three variables by $X_{j,1}$, $X_{j,2}$, and $X_{j,3}$ for $j\in[n]$.
-W.l.o.g.
-assume that no variable occurs two times in the same clause (a variable and its reversal, however, may still occur together in one clause).
-Define $n+1$ functions $f_{j}$ with $j\in[n]_{0}$ such that $f_{j}:\mathcal{X}\rightarrow X$ with $f_{j}(X_{i})=u_{i,j}$ and $f_{j}(\overline{X_{i}})=v_{i,j}$ for $i\in[m]$.
-Now, let
-$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
-with
-$s_{i}$ $=1\#1,\text{ and}$
-$t_{j}$ $=1$
-for $i\in[m]$ and $j\in[n]$ and construct a relational pattern $(\alpha,r_{\alpha})$ by
-$\alpha=\#\#\alpha_{1,1}\ \#\#\ ...\ \#\#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\#\ ...\ \#\#\ \alpha_{2,n}\ \#\#$
-with
-$\alpha_{1,i}$ $=f_{0}(X_{i})f_{0}(\overline{X_{i}})\text{, and}$
-$\alpha_{2,j}$ $=f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3})$
-and the set of related variables $r_{\alpha}=\{\ (r,u_{i,j},u_{i,j^{\prime}}),(r,v_{i,j},v_{i,j^{\prime}})\ |\ i\in[m],j,j^{\prime}\in[n+1]\ \}$.
-Hence, notice that two variables with the same base name $u_{i,j}$ and $u_{i^{\prime},j^{\prime}}$ (resp.
-$v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$) are related to each other if and only if $i=i^{\prime}$.
-First, assume there exists a satisfying assignment for $\varphi$.
-Let $h\in H$ such that for all $i\in[m]$ and $j\in[n]_{0}$ we have $h(u_{i,0})=1$ and $h(v_{i,0})=\#1$ iff $X_{i}$ is true (hence $\overline{X_{i}}$ is false) and otherwise $h(u_{i,0})=1\#$ and $h(v_{i,0})=1$ iff $X_{i}$ is false (hence $\overline{X_{i}}$ is true).
-As the assignment for $\varphi$ is true, we have for each clause that there is at least one variable $u_{i,j}$ (or $v_{i,j}$) such that $(r_{com^{*}},u_{i,j},u_{i,0})\in r_{\alpha}$ (or $(r_{com^{*}},v_{i,j},v_{i,0})$) and $h(u_{i,0})=1$ (or $h(v_{i,0})=1$).
-Hence we may set $h(u_{i,j})=1$ (or $v_{i,j}=1$) and for all other variables $z$ in that clause set $h(z)=\varepsilon$.
-As the empty word commutes with any other word by the definition of $r_{com^{*}}$, this substitution is $r_{\alpha}$-valid.
-As we can do that for each $j\in[n]$, we get can find a $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$, concluding this direction.
-For the other direction assume $w\in L_{E}(\alpha,r_{\alpha})$.
-Then there exists some $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$.
-For each $j\in[n]$, we know that $h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))=1$.
-Hence, w.l.o.g.
-there exists one variable $u_{i,j}$ for $i\in[m]$, such that $h(u_{i,j})=1$ and $(r_{com^{*}},u_{i,1},u_{i,j})\in r_{\alpha}$.
-By the definition of $r_{com^{*}}$ and $w$ we know that $h(u_{i,0})\in\{\varepsilon,1\}$ for $h$ to be $r_{\alpha}$-valid.
-In that case we know that $h(v_{i,0})\in\{1\#1,\#1\}$ and by that we have $h(v_{i,j})\neq 1$ for any $j\in[n]$ by the definition of $r_{\alpha}$.
-So, we can set an assignment of $X$ for $\varphi$ by setting $X_{i}=true$ iff $h(u_{i,1})\in\{\varepsilon,1\}$ and setting $\overline{X_{1}}=true$ iff $h(v_{i,1})\in\{\varepsilon,1\}$.
-As for each clause $j$ we find either a variable $u_{i,j}$ or $v_{i,j}$ substituted to $1$, we know that the assignment for $X$ is satisfying for $\varphi$.
-That concludes the reduction.
-∎
-0.A.1 The Membership Problem under the One-Sided Star and One-Sided Subsequence Relations
-In this subsection, we briefly discuss the special cases mentioned in the end of [Angluin1980], which considers one-sided relations of potential interest.
-These cases cannot be expressed directly using the framework of this paper.
-Consider, for example, the pattern $\alpha=\mathtt{a}x\mathtt{b}x^{*}\mathtt{c}$ that assumes that $h(x^{*})\in\{h(x)\}^{*}$ needs to hold for some valid substitution $h\in H$.
-This can be mapped using the relation $r_{*}$, but the way Angluin described that case, we should not be able to have the $r_{*}$ hold in both directions between $x$ and $x^{*}$, which, however, is allowed in relational pattern languages over $r_{*}$.
-Hence, these special cases of one-sided relations are considered for the subsequence and the star relation at the end of this section.
-For simplicity, only in this subsection, we denote patterns similar to the notation used in [Angluin1980], i.e., for the one-sided star we write, for example, $\alpha=\#x\#x^{*}\#$ instead of introducing some variables $x_{1}$ and $x_{2}$ separately and relating them externally by some relation $r$.
-In particular, in this subsection, given $x$ and $x^{*}$, a substitution $h$ is valid if and only if $h(x^{*})\in\{x\}^{*}$, and, given $x$ and $x^{s}$, a substitution $h$ is valid if and only if $h(x^{s})\in\operatorname{SubSeq}(x)$.
-For the erasing case, we obtain that the same reduction can be used for both cases, one-sided star and one-sided subsequence.
-Proposition 6
-Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{*}$ (resp.
-$x^{s}$) may occur multiple times, it is NP-complete to decide whether $w\in L_{E}(\alpha)$.
-Proof
-NP-containment follows from the fact that we can guess a valid substitution $h$ that results in $|h(\alpha)|=|w|$ and check whether $h(\alpha)=w$ holds, all in polynomial time.
-To establish NP-hardness, we reduce from $3\mathtt{-SAT}$ as before and use the same formal framework as used in the reductions before.
-Assuming the number of variables in the formula is $m\in\mathbb{N}$ and assuming the number of clauses of the formula is $n\in\mathbb{N}$, the word $w\in\Sigma^{*}$ will be set to
-$w=\#\#\underbrace{\ 1\ \#\ ...\ \#\ 1\ }_{m\text{ blocks}}\#\#\underbrace{\ 1\ \#\ ...\ \#\ 1\ }_{n\text{ blocks}}\#\#.$
-The corresponding pattern $\alpha$ is given by
-$\alpha=\#\#\ \alpha_{1,1}\ \#\ ...\ \#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\ ...\ \#\ \alpha_{2,n}\ \#\#$
-where $\alpha_{1,i}=u_{i}v_{i}$, for $i\in[m]$ and $u_{i},v_{i}$ being new pattern variables, and $\alpha_{2,i}=f(X_{i,1})f(X_{i,2})f(X_{i,3})$, where $f(X_{i,j})=u_{k}^{*}$ if $X_{i,j}=X_{k}$ (resp.
-$f(X_{i,j})=v_{k}^{*}$ if $X_{i,j}=\overline{X_{k}}$) in the one-sided star framework, and $f(X_{i,j})=u_{k}^{s}$ if $X_{i,j}=X_{k}$ (resp.
-$f(X_{i,j})=v_{k}^{s}$ if $X_{i,j}=\overline{X_{k}}$) in the one-sided subsequence framework.
-The correctness of the reduction comes from the fact that, for $w\in L_{E}(\alpha)$ to hold, exactly one variable in each $\alpha_{1,i}$, for $i\in[m]$, must be set to $1$, either $h(u_{i})=1$ or $h(v_{i})=1$.
-A variable $u_{i}^{*}$ or $v_{i}^{*}$ (resp.
-$u_{i}^{s}$ or $v_{i}^{s}$), for $i\in[m]$, that occurs in some $\alpha_{2,j}$, for $j\in[m]$, can only be substituted by $1$ if the corresponding variable in $u_{i}$ or $v_{i}$ in $\alpha_{1,i}$ is substituted by $1$.
-Bringing these two facts together immediately results in that the $3\mathtt{-SAT}$ formula must have an accepting configuration if and only if $w\in L_{E}(\alpha)$, as we can only substitute all $\alpha_{2,i}$, for $i\in[n]$, by $1$ if and only if there exists some configuration of the variables where either $X_{k}$ or $\overline{X_{k}}$ is set to true and each clause contains at least one of these variables that are set to true.
-∎
-For the non-erasing cases, we can again use similar reductions as for all the other problems before, but the configuration is slightly different for one-sided star and one-sided subsequence, respectively.
-Proposition 7
-Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{*}$ may occur multiple times, it is NP-complete to decide whether $w\in L_{NE}(\alpha)$.
-Proof
-NP-containment follows from the fact that we can guess a valid substitution $h$ that results in $|h(\alpha)|=|w|$ and check whether $h(\alpha)=w$ holds, all in polynomial time.
-To establish NP-hardness, we reduce from $3\mathtt{-SAT}$ as before, again, assuming that a given formula $\phi$ has $n$ clauses and uses $m$ distinct variables.
-We set $w\in\Sigma^{*}$ to
-$w=\#\#\underbrace{\ 1^{3}\ \#\ ...\ \#\ 1^{3}\ }_{m\text{ blocks}}\#\#\underbrace{\ 1^{6}\ \#\ ...\ \#\ 1^{7}\ }_{n\text{ blocks}}\#\#$
-The pattern $\alpha$ is basically a prefix of the construction used in Lemma 1, i.e.,
-$\alpha=\#\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\#,$
-where $p_{1}$ to $p_{m}$ and $q_{1}$ to $q_{m}$ are constructed as in the proof of Lemma 1, just with the only difference that every occurrence of $u_{i}$ or $v_{i}$, for $i\in[m]$, in some $q_{1}$ to $q_{n}$, is replaced by $u_{i}^{*}$ and $v_{i}^{*}$ respectively.
-Then, each $u_{i}$ and $v_{i}$ without a start occurs only once, that is in $p_{i}$.
-All other occurrences are with a star.
-The variables $z_{i}$, for $i\in[n]$, are unchanged in $q_{i}$ to $q_{n}$.
-To show correctness of this reduction, notice the fact that each $p_{i}$ is still to be substituted to $1^{3}$, resulting in one of $u_{i}$ or $v_{i}$ being substituted by $1$ while the other is substituted by $11$.
-Assume that a substitution to $1$ represents a true assignment while a substitution by $11$ represents a false assignment.
-For each clause $i\in[n]$, we know that $q_{i}=x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*}z_{i}$, assuming each variable $x_{i,j}$ is either some $u_{k}$ or $v_{k}$, for $k\in[m]$.
-We know that $z_{i}$ must be substituted by a word of at least length $1$.
-Hence, for $h(q_{i})=1^{6}$ to be the case, we then need $|h(x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*})|\leq 5$.
-If $h(x_{i,j})=11$, for all $j\in[3]$, then we know that $|h(x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*})|\geq 6$, as $h(x_{i,j})=\varepsilon$ is not valid in the non-erasing case.
-So, for at least one variable $x_{i,j}$ we need $h(x_{i,j})=1$ for $h(q_{i})=1^{6}$ to be possible.
-As this holds for all clauses, we can only find a substitution $h$ for which $h(\alpha)=w$ if and only if the given $3\mathtt{-SAT}$ is satisfiable.
-∎
-As we only do slight changed to the constructed word and pattern, we give the proof of the following result in a very short form and assume the same formal framework as in Proposition 7.
-Proposition 8
-Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{s}$ may occur multiple times, it is NP-complete to decide whether $w\in L_{NE}(\alpha)$.
-Proof
-NP-containment follows as before, e.g. as in Proposition 7. For NP-hardness, we reduce from $3\mathtt{-SAT}$ as before. The word $w\in\Sigma^{*}$ is given by
-$w=\#\#\underbrace{\ 1^{3}\ \#\ ...\ \#\ 1^{3}\ }_{m\text{ blocks}}\#\#\underbrace{\ 1^{4}\ \#\ ...\ \#\ 1^{4}\ }_{n\text{ blocks}}\#\#.$
-The pattern $\alpha$ is given by
-$\alpha=\#\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\#,$
-where $p_{i}=u_{i}v_{i}$, for $i\in[m]$ and new variables $u_{i}$ and $v_{i}$, and we have $q_{i}=f(X_{i,1})f(X_{i,2})f(X_{i,3})$ for $f$ being defined as before as well, just using $u_{j}^{s}$ and $v_{j}^{s}$ instead of $u_{j}$ and $v_{j}$, for $j\in[m]$ and $i\in[n]$.
-Now notice, again, that for each $p_{i}$, for $i\in[m]$, one of $u_{i}$ and $v_{i}$ must be substituted by $1$ while the other one is substituted by $11$.
-Here, we consider a substitution by $11$ as a true assignment of that variable configuration.
-For all other variable substitutions we know that they must be a subsequence of either some $u_{i}$ or $v_{i}$.
-As we are in the non-erasing case, for a substitution $h$ to be valid, that results in w.l.o.g.
-$h(u_{i}^{s})=1$ if $h(u_{i})=1$ and $h(u_{i}^{s})\in\{1,11\}$ if $h(u_{i})=11$, for $i\in[m]$ (respectively for $v_{i}$).
-So, we can only obtain $h(q_{i})=1^{4}$, for all $i\in[n]$, if and only if there is at least one variable $u_{i}^{s}$ (or $v_{i}^{s}$) in $q_{i}$, where $h(u_{i})=11$ (or $h(v_{i})=11$).
-By the same arguments as in all results before, we can only find a substitution $h\in H$ for which $h(\alpha)=w$ if and only if there exists a satisfying assignment of the variable for the given $3\mathtt{-SAT}$ formula, concluding this reduction.
-∎
-This concludes all results necessary for Theorem 3.2 to hold.
-Appendix 0.B Additional Content for Inclusion Problem
-0.B.1 Definition of Nondeterminstic 2-Counter Automata without Input
-A nondeterministic 2-counter automaton without input (see e.g.
-[Iberra1978]) is a 4-tuple $A=(Q,\delta,q_{0},F)$ which consists of a set of states $Q$, a transition function $\delta:Q\times\{0,1\}^{2}\rightarrow\mathcal{P}(Q\times\{\-1,0,+1\}^{2})$, an initial state $q_{0}\in Q$, and a set of accepting states $F\subseteq Q$.
-A configuration of $A$ is defined as a triple $(q,m_{1},m_{2})\in Q\times\mathbb{N}\times\mathbb{N}$ in which $q$ indicates the current state and $m_{1}$ and $m_{2}$ indicate the contents of the first and second counter.
-We define the relation $\vdash_{A}$ on $Q\times\mathbb{N}\times\mathbb{N}$ by $\delta$ as follows.
-For two configurations $(p,m_{1},m_{2})$ and $(q,n_{1},n_{2})$ we say that $(p,m_{1},m_{2})\vdash_{A}(q,n_{1},n_{2})$ if and only if there exist $c_{1},c_{2}\in\{0,1\}$ and $r_{1},r_{2}\in\{-1,0,+1\}$ such that
-  1. 1.
-    if $m_{i}=0$ then $c_{i}=0$, otherwise if $m_{i}>0$, then $c_{i}=1$, for $i\in\{1,2\}$,
-  2. 2.
-    $n_{i}=m_{i}+r_{i}$ for $i\in\{1,2\}$,
-  3. 3.
-    $(q,r_{1},r_{2})\in\delta(p,c_{1},c_{2})$, and
-  4. 4.
-    we assume if $c_{i}=0$ then $r_{i}\neq-1$ for $i\in\{1,2\}$.
-Essentially, the machine checks in every state whether the counters equal 0 and then changes the value of each counter by at most one per transition before entering a new state.
-A computation is a sequence of configurations.
-An accepting computation of $A$ is a sequence $C_{1},...,C_{n}\in(Q\times\mathbb{N}\times\mathbb{N})^{n}$ with $C_{1}=(q_{0},0,0)$, $C_{i}\vdash_{A}C_{i+1}$ for all $i\in\{1,...,n-1\}$, and $C_{n}\in F\times\mathbb{N}\times\mathbb{N}$ for some $n\in\mathbb{N}$ with $n>0$.
-We encode configurations of $A$ by assuming $Q=\{q_{0},...,q_{e}\}$ for some $e\in\mathbb{N}$ and defining a function $\operatorname{enc}$ :: $Q\times\mathbb{N}\times\mathbb{N}\rightarrow\{0,\#\}^{*}$ by
-$\operatorname{enc}(q_{i},m_{1},m_{2}):=0^{x+i}\#0^{c_{1}+y_{2}\cdot m_{1}}\#0^{c_{2}+y_{2}\cdot m_{2}}$
-for some numbers $x,c_{1},y_{2},c_{1},y_{2}\in\mathbb{N}$.
-The values for these numbers depend on the construction of the respective proofs and are not specified here.
-This is extended to encodings of computations by defining for every $n\geq 1$ and every sequence $C_{1},...,C_{n}\in Q\times\mathbb{N}\times\mathbb{N}$
-$\operatorname{enc}(C_{1},...,C_{n}):=\#\#\ \operatorname{enc}(C_{1})\ \#\#\ ...\ \#\#\ \operatorname{enc}(C_{n})\ \#\#.$
-Encodings of this kind are used to prove different undecidability results regarding the inclusion problem for erasing relational pattern languages for various considered relations, in particular the results in Subsection 4.2.
-For some nondeterministic 2-counter automaton without input $A$, define the set of encodings of accepting computations
-$\mathtt{ValC}(A):=\{\operatorname{enc}(C_{1},...,C_{n})\ |\ C_{1},...,C_{n}\text{ is an accepting computation of }A\}$
-and let $\mathtt{InvalC}(A)=\{0,\#\}^{*}\setminus\mathtt{ValC}(A)$.
-The emptiness problem for deterministic 2-counter-automata is undecidable (cf.
-e.g.
-[Iberra1978, Minsky1961]), thus it is also undecidable whether a nondeterministic 2-counter automaton without input has an accepting computation [Freydenberger2010, Jiang1995].
-That the emptiness problem for universal Turing machines is undecidable is a known fact.
-0.B.2 Definition of the Universal Turing Machine $U$
-Here, we define the universal Turing machine $U$ used in the referenced proof from [BREMER201215] and referred to in the proofs for the undecidability of the inclusion problem for NE relational pattern languages under $r_{ab}$ and rev in Proposition 5 and Corollary 2.
-Let $U=(Q,\Gamma,\delta)$ be the universal Turing machine $U_{15,2}$ wih 2 symbols and 15 states as described by Neary and Woods [Neary2009].
+Appendix 0.B Additional Content for the Inclusion Problem
+0.B.1 Proof Sketch of Theorem˜4.1
+0.B.1.1 (A Specific Universal Turing Machine):
+Here, we define the universal Turing machine $U$ used in the referenced proof from [5] and referred to in the proofs for the undecidability of the inclusion problem for NE relational pattern languages under $r_{ab}$ and rev in this section.
+Let $U=(Q,\Gamma,\delta)$ be the universal Turing machine $U_{15,2}$ wih 2 symbols and 15 states as described by Neary and Woods [31].
 This machine hat the states $Q=\{q_{1},...,q_{15}\}$ and the tape alphabet $\Gamma=\{0,1\}$.
 The transition function $\delta:\Gamma\times Q\rightarrow(\Gamma\times\{L,R\}\times Q)\cup\mathtt{HALT}$ is given in Table 2.
-We follow with the definition of encodings of computations of $U$ as depicted in [BREMER201215].
+We follow with the definition of encodings of computations of $U$ as depicted in [5].
 The following conventions are needed to discuss configurations of $U$.
@@
 It is valid if $C_{1}=I$ (II being some initial configuration), $C_{n}$ is a halting configuration, and $C_{i+1}$ is a valid successor configuration of $C_{i}$, for $i\in[n-1]$, as defined by $\delta$.
-In [BREMER201215], the notion is adopted that any possible configuration where both tape sides have a finite value under $e$ is a valid successor configuration of a halting configuration.
+In [5], the notion is adopted that any possible configuration where both tape sides have a finite value under $e$ is a valid successor configuration of a halting configuration.
 The encoding of computations of $U$ is given analogously to the definition in the case of nondeterministic 2-counter automata without input, i.e., for some computation $\mathcal{C}=(C_{1},...,c_{n})$, we have
@@
 11 $(1,L,q_{10})$ $\mathtt{HALT}$ $(1,R,q_{14})$ $(1,R,q_{12})$ $(1,R,q_{12})$ $(0,R,q_{15})$ $(1,R,q_{14})$
-Table 4: Transition table of $U$, i.e., definition of $\delta$, as it is given in [BREMER201215] or [Neary2009].
-0.B.3 Undecidability of the Inclusion Problem for Several Basic Relations
-Proof of Corollary˜1:
-Proof
-Consider the construction of Freydenberger and Reidenbach [Freydenberger2010] and the construction of Bremer and Freydenberger [BREMER201215], respectively.
-The main differences between E-pattern languages $L_{E}(\alpha)$ (resp.
-NE pattern languages $L_{NE}(\alpha)$) of a pattern $\alpha\in Pat_{\Sigma}$ in [Freydenberger2010, BREMER201215]and relational pattern languages $L_{E}(\alpha,r_{\alpha})$ (resp.
-$L_{NE}(\alpha,r_{\alpha})$) that we consider are
-  1. 1.
-    In relational patterns each variable may occur only once whereas in classical patterns each variable may occur any number of times.
-  2. 2.
-    In relational pattern languages the substitution must be $r_{\alpha}$-valid whereas in pattern languages there is no relation.
-Thus, if we use the pattern given in relational pattern languages to simulate identical variables, there is no significant difference between relational pattern languages and classical pattern languages.
-With this, we can immediately refer to the proofs in [Freydenberger2010, BREMER201215].
-First, consider $r_{=}$.
-We can use the proofs in [Freydenberger2010, BREMER201215].
-For each variable $x$, we replace each multiple occurrence of $x$ by new variables $x_{i},i\in\mathbb{N}$ with $(r_{=},x,x_{i})\in r_{\alpha}$.
-Since $(x,x_{i})\in r_{\alpha}$, we get $h(x)=h(x_{i})$ for each $r_{\alpha}$-valid substitution $h$.
-As $x_{i}$ is a new variable, we do not change anything else.
-Thus, this slightly adapted proof works in exactly the same way as the original proof and we get that the inclusion problem for relational pattern languages with the relation $r_{=}$ is undecidable.
-Now, we show that we can do the same with the relations $r_{ssq}$ and $r_{*}$.
-For $r_{ssq}$, we get $(r_{ssq},x,y),(r_{ssq},y,x)\in r_{\alpha}$ implies $h(x)=h(y)$ for any two variables $x,y$ and any $r_{\alpha}$-valid substitution $h$ by the definition of $r_{ssq}$.
-For each variable $x$, we replace each multiple occurrence of $x$ by new variables $x_{i},i\in\mathbb{N}$ with $(r_{ssq},x,x_{i}),(r_{ssq},x_{i},x)\in r_{\alpha}$.
-With this, the proof works analogously to the one of [Freydenberger2010, BREMER201215].
-For $r_{*}$, we get the same observation: $(r_{*},x,y),(r_{*},y,x)\in r_{\alpha}$ implies $h(x)=h(y)$ for any two variables $x,y$ and any $r_{\alpha}$-valid substitution $h$.
-Thus, the proof works here as well.
-So, in total, we get that the inclusion problem for relational pattern languages with the relation $\tilde{r}$ for $\tilde{r}\in\{r_{=},r_{ssq},r_{*}\}$ is undecidable.
-0.B.4 Undecidability of the Inclusion Problem for Non-Erasing Relational Pattern Languages under Abelian Equivalence
-Proof
-As this proof works almost identical to the proof of Bremer and Freydenberer in [BREMER201215], we only give an extended sketch and focus on the formal intricacies that occur as we are now considering abelian equivalence instead of the equivalence relation.
-Let $\Sigma=\{0,\#\}$ and $R:=\{r_{ab}\}$.
+Table 2: Transition table of $U$, i.e., definition of $\delta$, as it is given in [5] or [31].
+0.B.1.2 (The Full Proof Idea):
+Proof.
+As this proof works almost identical to the proof of Bremer and Freydenberer in [5], we only give an extended sketch and focus on the formal intricacies that occur as we are now considering abelian equivalence instead of the equivalence relation.
+Let $\Sigma=\{0,\#\}$ and w.l.o.g. $R:=\{r_{ab}\}$. The arguments work analogously for $r_{rev}$ as both, $r_{rev}$ and $r_{ab}$, force equality on unary words and, in particular, single letters.
 To show undecidability of this problem, we reduce the problem of deciding whether $\mathtt{UnivValc}_{NE}(I)=\emptyset$ for some initial configuration $I$ of $U$ to the problem of deciding whether $L_{NE}(p_{1},\alpha_{R})\subseteq L_{NE}(p_{2},\beta_{R})$ for two relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$.
@@
 Using that, we say $t:=\psi(r_{1}\hat{\beta}_{1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$.
-We see that both patterns are basically constructed in the same way as in [BREMER201215], just using the formalism of relational patterns, i.e.
-all variables which are equal in the construction in [BREMER201215], are in this construction distinct variables that are, however, related to each other through the sets $r_{\alpha}$ and $r_{\beta}$.
-The following lemma follows directly from the proof in [BREMER201215] which can be converted directly to the formalism using substitutions over relational patterns.
-Lemma 6
-[BREMER201215] All $\psi(\hat{\beta}_{i})$ with $1\leq i\leq\mu$ and $t$ begin and end with the terminal letter 0 and do not contain the terminal word $^{4}$ as a factor.
+We see that both patterns are basically constructed in the same way as in [5], just using the formalism of relational patterns, i.e.
+all variables which are equal in the construction in [5], are in this construction distinct variables that are, however, related to each other through the sets $r_{\alpha}$ and $r_{\beta}$.
+The following lemma follows directly from the proof in [5] which can be converted directly to the formalism using substitutions over relational patterns.
+Lemma 6.
+[5] All $\psi(\hat{\beta}_{i})$ with $1\leq i\leq\mu$ and $t$ begin and end with the terminal letter 0 and do not contain the terminal word $^{4}$ as a factor.
 We recall, that $H_{r_{\alpha}}$ denotes the set of all $r_{\alpha}$-valid substitutions.
@@
 We proceed with more technical preperations that we need for the easy and dynamic construction of predicates.
-Once this is settled, as the constructed predicates from [BREMER201215] also work in this case, we are done.
+Once this is settled, as the constructed predicates from [5] also work in this case, we are done.
 We say that a substitution $h\in H_{r_{\alpha}^{+}}$ is of bad form if $h(0\alpha_{1}0)$ contains a factor $^{3}$ or if $h(0\alpha_{2}0)$ contains the letter #\#.
-We use the predicates $\pi_{1}$ and $\pi_{2}$ to handle all such substitutions of bad form and construct them exactly the same way as they are constructed in [BREMER201215] by saying
+We use the predicates $\pi_{1}$ and $\pi_{2}$ to handle all such substitutions of bad form and construct them exactly the same way as they are constructed in [5] by saying
 $\upgamma_{1}:=y_{1,1}\#^{k}y_{1,2}$ $\upgamma_{2}:=0y_{2}0$
@@
 for new variables $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2}\in X$ that are not related to any other variable in $\beta$.
-As they are not related to any other variable, no intricacies regarding abelian equivalence in comparison to normal equality occur and we directly get the following lemma from [BREMER201215] as their proof needs no adaptation.
-Lemma 7
-[BREMER201215] A substitution $h\in H_{r_{\alpha}^{+}}$ is of bad-form if and only if $h$ satisfies $\pi_{1}$ or $\pi_{2}$.
+As they are not related to any other variable, no intricacies regarding abelian equivalence in comparison to normal equality occur and we directly get the following lemma from [5] as their proof needs no adaptation.
+Lemma 7.
+[5] A substitution $h\in H_{r_{\alpha}^{+}}$ is of bad-form if and only if $h$ satisfies $\pi_{1}$ or $\pi_{2}$.
 The following key lemma allows us to focus only on predicates from now on and is independent of the shape of the specific predicates $\pi_{3}$ to $\pi_{\mu}$.
-Lemma 8
+Lemma 8.
 For every non-erasing substitution $h\in H_{r_{\alpha}^{+}}$, we have that $h(\alpha)\in L_{NE}(\beta,r_{\beta})$ if and only if $h$ satisfies some predicate $\pi_{1}$ to $\pi_{\mu}$.
-Proof
-Due to the different formalism, we need to rewrite certain parts of the proof from [BREMER201215].
-However, the general idea is the same.
-We just have to show that it also works for abelian equivalence.
+Proof.
+Due to the different formalism, we need to rewrite certain parts of the proof from [5]. However, the general idea is the same. We just have to show that it also works for abelian equivalence.
 ($\Rightarrow:$) Assume some $h\in H_{r_{\alpha}^{+}}$ satisfies some predicate $\pi_{i}$ for $i\in\{1,...,\mu\}$.
@@
 By that, we can always define $\tau^{\prime}$ that way.
-Also note, that the only significant difference to [BREMER201215] lies in the definition of $\tau^{\prime}(x_{j,k})$, $\tau^{\prime}(a_{k})$, and $\tau^{\prime}(b_{k})$ as no variables can occur multiple timesin relational patterns; the rest remains the same.
-As in [BREMER201215] we obtain $\tau^{\prime}(\upgamma_{i})=\tau(\gamma_{i})=h(0\alpha_{1}0)$, $\tau^{\prime}(\updelta_{i})=\tau(\updelta_{i})=h(0\alpha_{1}0)$, and
+Also note, that the only significant difference to [5] lies in the definition of $\tau^{\prime}(x_{j,k})$, $\tau^{\prime}(a_{k})$, and $\tau^{\prime}(b_{k})$ as no variables can occur multiple timesin relational patterns; the rest remains the same.
+As in [5] we obtain $\tau^{\prime}(\upgamma_{i})=\tau(\gamma_{i})=h(0\alpha_{1}0)$, $\tau^{\prime}(\updelta_{i})=\tau(\updelta_{i})=h(0\alpha_{1}0)$, and
 $\tau^{\prime}(a_{1}b_{1}\#^{5}a_{2}x_{1,1}\cdots x_{\mu,1}b_{2}\#^{5})$ $=0^{\mu-i+1}0^{i}\#^{5}0^{\mu-i+1}0^{i-1}\#0^{\mu-i}0^{i}\#^{5}$
@@
 $t$ $=\tau(r_{j+1}\hat{\beta}_{j+1}\cdots r_{\mu}\hat{\beta}_{\mu}r_{\mu+1})$
-As these equations must hold and especially as $h(0\alpha_{1}0)=\tau(\upgamma_{j})$ and $h(0\alpha_{2}0)=\tau(\updelta_{j})$, we know that $h$ satisfies the predicate $\pi_{j}$. ∎
+As these equations must hold and especially as $h(0\alpha_{1}0)=\tau(\upgamma_{j})$ and $h(0\alpha_{2}0)=\tau(\updelta_{j})$, we know that $h$ satisfies the predicate $\pi_{j}$. ∎∎
 By that, we can now select appropriate predicates $\pi_{1}$ to $\pi_{\mu}$ such that we obtain $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$.
@@
 For technical reasons, we define a predicate that sets a lower bound to the length of $h(\alpha_{2})$ if unsatisfied for a substitution $h\in H_{r_{\alpha}^{+}}$ of good form.
-As in [BREMER201215], the predicate $\pi_{3}$ is defined by
+As in [5], the predicate $\pi_{3}$ is defined by
 $\upgamma_{3}$ $:=y_{3,1}\hat{y}_{3,1}y_{3,2}\hat{y}_{3,2}y_{3,3}\hat{y}_{3,3}y_{3,4}$
@@
 for pairwise non-related variables $y_{3,1},...,y_{3,4}$ and $\hat{y}_{3,1},...,\hat{3,3}$.
-As in [BREMER201215], if some $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{3}$, then $h(\alpha_{2})$ is a concatenation of three non-empty factors of $h(\alpha_{1})$.
+As in [5], if some $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi_{3}$, then $h(\alpha_{2})$ is a concatenation of three non-empty factors of $h(\alpha_{1})$.
 In other words, if some $h$ does not satisfy any of $\pi_{1}$$\pi_{2}$, and $\pi_{3}$, then $h(\alpha_{2})\in 0^{+}$ has to be longer than the $3$ longest non-overlapping and non-touching factors of 0’s in $h(\alpha_{1})$.
 This restriction allows for a simpler way to construct all other predicates $\pi_{4}$ to $\pi_{\mu}$.
-As we will see, the next part regarding simple predicates is also very similar to the one defined and used in [BREMER201215] but had to be adapted formally to also work in the context of relational patterns.
+As we will see, the next part regarding simple predicates is also very similar to the one defined and used in [5] but had to be adapted formally to also work in the context of relational patterns.
 Let $X_{s}:=\{\hat{x}_{1,1},\hat{x}_{2,1},\hat{x}_{3,1},\hat{x}_{1,2},\hat{x}_{2,2},\hat{x}_{3,2},...\}\subset X$ and always assume that $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ in any set of relational constraints $r_{p}$ of any relational pattern $(p,r_{p})\in\mathtt{RelPat}_{\Sigma,R}$ if any variable of $X_{s}$ is used in $p$.
@@
 For any relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ we define $S(p_{s},r_{p_{s}}):=\{s(p_{s})|s\in S\}$.
-The next definition is exactly the same as in [BREMER201215], just adapted to the notation of relational patterns.
-Definition 1
-[BREMER201215] A predicate $\pi:G_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ is called a NE-simple predicate for $0\alpha_{1}0$ if there exists a relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ and $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p_{s}}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ and languages $L_{1}\in\{0\Sigma^{*},\{0\}\}$ and $L_{2}\in\{\Sigma^{*}0,\{0\}\}$ such that any non-erasing substitution $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi$ if and only if $h(0\alpha_{1}0)\in L_{1}S(p_{s},r_{p_{s}})L_{2}$.
+The next definition is exactly the same as in [5], just adapted to the notation of relational patterns.
+Definition 1.
+[5] A predicate $\pi:G_{r_{\alpha}^{+}}\rightarrow\{0,1\}$ is called a NE-simple predicate for $0\alpha_{1}0$ if there exists a relational pattern $(p_{s},r_{p_{s}})$ with $p_{s}\in(\Sigma\cup X_{s})^{*}\cap\mathtt{RelPat}_{\Sigma,R}$ and $(r_{ab},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\in r_{p_{s}}$ for $i\in\{1,2,3\}$ and $j,j^{\prime}\in\mathbb{N}$ and languages $L_{1}\in\{0\Sigma^{*},\{0\}\}$ and $L_{2}\in\{\Sigma^{*}0,\{0\}\}$ such that any non-erasing substitution $h\in H_{r_{\alpha}^{+}}$ satisfies $\pi$ if and only if $h(0\alpha_{1}0)\in L_{1}S(p_{s},r_{p_{s}})L_{2}$.
 If $L_{1}=0\Sigma^{*}$ and $L_{2}=\Sigma^{*}0$, then we call $\pi$ an infix-predicate.
@@
 It shows that for each simple predicate $\pi_{s}$, we can construct a predicate $\pi$ defined over some $\upgamma$ and $\updelta$ such that each substitution of good form that would have satisfied $\pi_{s}$ now satisfies either $\pi_{3}$ or $\pi_{s}$ which suffices the demands of the construction.
-As we can see, the argument is again very similar to the one in [BREMER201215] and had to be adapted mainly to the new formalism.
-Lemma 9
+As we can see, the argument is again very similar to the one in [5] and had to be adapted mainly to the new formalism.
+Lemma 9.
 For every NE-simple predicate $\pi_{s}$, there exists a predicate $\pi$ defined by relational patterns $(\upgamma,r_{\upgamma})$ and $(\updelta,r_{\updelta})$ which can be embedded in $(\beta,r_{\beta})$ such that for all non-erasing substitutions $h\in G_{r_{\alpha}^{+}}$ we have
@@
     if $h$ satisfied the newly constructed $\pi$, then $h$ also satisfies $\pi_{s}$.
-Proof
+Proof.
 First, consider the case of $L_{1}=0\Sigma^{*}$ and $L_{2}=\Sigma^{*}0$.
@@
 So, $h$ must satisfy the simple predicate $\pi_{s}$.
-The other cases, where $L_{1}$ or $L_{2}$ is chosen to be the language containing only the word 0, are handled analogously by omitting $y_{1}$ or $y_{2}$ and keeping the rest of the construction. ∎
+The other cases, where $L_{1}$ or $L_{2}$ is chosen to be the language containing only the word 0, are handled analogously by omitting $y_{1}$ or $y_{2}$ and keeping the rest of the construction. ∎∎
 So, in general, if $h$ is of good form but does not satisfy $\pi_{3}$, then $h(\alpha_{2})\in 0^{+}$ is long enough to provide building blocks for NE-simple predicates that use variables from $X_{s}$.
-Lemma 9 shows us that the construction of NE-simple predicates in [BREMER201215] can be used in the setting of relational patterns under abelian equivalence in exactly the same way.
+Lemma 9 shows us that the construction of NE-simple predicates in [5] can be used in the setting of relational patterns under abelian equivalence in exactly the same way.
 Mainly, this is caused by the fact that each variable $\hat{x}_{i,j}\in X_{s}$ is substituted by a unary string.
-By that, each related variable has to be substituted exactly in the same way as they are substituted in [BREMER201215], i.e.
+By that, each related variable has to be substituted exactly in the same way as they are substituted in [5], i.e.
 equally.
-Having shown that the same simple predicates from [BREMER201215] also work in this setting in exactly the same way, we define $\pi_{4}$ to $\pi_{\mu}$ exactly as they are defined in [BREMER201215], as all of those predicates are given described as simple predicates.
+Having shown that the same simple predicates from [5] also work in this setting in exactly the same way, we define $\pi_{4}$ to $\pi_{\mu}$ exactly as they are defined in [5], as all of those predicates are given described as simple predicates.
 We set $\alpha_{1}:=\#\#enc_{NE}(I)\#\#x\#0^{6}0^{1}0\#\#$ and $\alpha_{2}:=y00$ for pairwise distinct variables $x$ and $y$ (notice, that $x$ and $y$ are the only variables in the pattern $\alpha$, hence for $(\alpha,r_{\alpha})$ we have $r_{\alpha}=\emptyset$).
-By the construction of $\pi_{1}$ to $\pi_{\mu}$ and essentially the same argument over the embedding of all invalid computations in $(\beta,r_{\beta})$ from [BREMER201215], we obtain that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$.
+By the construction of $\pi_{1}$ to $\pi_{\mu}$ and essentially the same argument over the embedding of all invalid computations in $(\beta,r_{\beta})$ from [5], we obtain that $L_{NE}(\alpha,r_{\alpha})\setminus L_{NE}(\beta,r_{\beta})=\emptyset$ if and only if $\mathtt{ValC}_{U}(I)=\emptyset$.
 This concludes the case $|\Sigma|=2$.
 For the case of larger alphabets, i.e.
-alphabets of size $|\Sigma|\geq 3$, the exact construction in [BREMER201215] can also be adapted to the formality of our setting and works by the same arguments as in [BREMER201215].
+alphabets of size $|\Sigma|\geq 3$, the exact construction in [5] can also be adapted to the formality of our setting and works by the same arguments as in [5].
 The relational patterns $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ need to be slightly adapted to work with larger alphabets.
@@
 Remember, that we have $(r_{ab},\tilde{x}_{1,1},\tilde{x}_{1,2})\in r_{\tilde{\beta}}$ and $(r_{ab},\tilde{x}_{2,1},\tilde{x}_{2,2})\in r_{\tilde{\beta}}$.
-Now, using these constructions, we obtain the exact argument as in [BREMER201215] to obtain the final result.
+Now, using these constructions, we obtain the exact argument as in [5] to obtain the final result.
 The main argument why this construction can be used by the same argument is that it works over unary substitutions of $\tilde{x}_{1,1},\tilde{x}_{1,2},\tilde{x}_{2,1},$ and $\tilde{x}_{2,2}$ of length $1$.
 Hence, using substitutions of this type, abelian equivalence behaves exactly the same as equality between variables.
-Thus, we omit the details at this point and refer to [BREMER201215] for a more detailed proof.
+Thus, we omit the details at this point and refer to [5] for a more detailed proof.
 This concludes the case of larger alphabets of size $|\Sigma|\geq 3$.
@@
 ∎
-0.B.5 Undecidability of the Inclusion Problem for Erasing Relational Pattern Languages under the Reversal Relation
-Proof of Theorem˜4.1:
-Proof
-The proof is based on the construction by Freydenberger and Reidenbach [Freydenberger2010].
-For better understanding for those who don’t know the construction and for a really detailed verification, we copied the proof of Theorem 2 of [Freydenberger2010] and adapted it at several points.
+0.B.2 Proof of Lemma˜4
+Proof.
+The proof relies on the proof of Lemma˜3.
+(1:) As shown in the first part of proof of Lemma˜3, the language of accepted words of $\alpha$ is
+$L_{E}(\alpha,r_{\alpha})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}$
+and the language of accepted words of $\beta$ is
+$L_{E}(\beta,r_{\beta})=\{\varepsilon\}\cup\{w\mid|w|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
+Thus, $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$ if and only if $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$.
+This property can be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
+(2:) Assume $|\Sigma|=1$ and $a\in\Sigma$.
+In order to be included, $\alpha$ must have at least as many terminals as $\beta$.
+Otherwise we could set all variables in $\alpha$ to $\varepsilon$ and obtain a word that is not in the language of $\beta$.
+Let $\ell_{\alpha}$ be the number of terminals in $\alpha$ and $\ell_{\beta}$ be the number of terminals in $\beta$.
+As shown in the second part of proof of Lemma˜3, the language of accepted words of $\alpha$ is
+$L_{E}(\alpha,r_{\alpha})=\{a^{\ell_{\alpha}}\}\cup\{w\mid|w|\geq\ell_{\alpha}+\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\}$
+and the language of accepted words of $\beta$ is
+$L_{E}(\beta,r_{\beta})=\{a^{\ell_{\beta}}\}\cup\{w\mid|w|\geq\ell_{\beta}+\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|\}.$
+Thus, $L_{E}(\alpha,r_{\alpha})\subseteq L_{E}(\beta,r_{\beta})$ if and only $\ell_{\alpha}\geq\ell_{\beta}$ and $\min_{C_{\alpha,i}\in\mathcal{C}_{\alpha}}|C_{\alpha,i}|\geq\min_{C_{\beta,i}\in\mathcal{C}_{\beta}}|C_{\beta,i}|$.
+This can be tested in time $O(n_{\alpha}+n_{\beta})$ by scanning once over the patterns and counting the sizes of the equivalence classes.
+∎
+0.B.3 Proof Sketch of Theorem˜4.3
+0.B.3.1 (Nondeterministic 2-Counter Automata Without Input):
+A nondeterministic 2-counter automaton without input (see e.g.
+[21]) is a 4-tuple $A=(Q,\delta,q_{0},F)$ which consists of a set of states $Q$, a transition function $\delta:Q\times\{0,1\}^{2}\rightarrow\mathcal{P}(Q\times\{\-1,0,+1\}^{2})$, an initial state $q_{0}\in Q$, and a set of accepting states $F\subseteq Q$.
+A configuration of $A$ is defined as a triple $(q,m_{1},m_{2})\in Q\times\mathbb{N}\times\mathbb{N}$ in which $q$ indicates the current state and $m_{1}$ and $m_{2}$ indicate the contents of the first and second counter.
+We define the relation $\vdash_{A}$ on $Q\times\mathbb{N}\times\mathbb{N}$ by $\delta$ as follows.
+For two configurations $(p,m_{1},m_{2})$ and $(q,n_{1},n_{2})$ we say that $(p,m_{1},m_{2})\vdash_{A}(q,n_{1},n_{2})$ if and only if there exist $c_{1},c_{2}\in\{0,1\}$ and $r_{1},r_{2}\in\{-1,0,+1\}$ such that
+  1. 1.
+    if $m_{i}=0$ then $c_{i}=0$, otherwise if $m_{i}>0$, then $c_{i}=1$, for $i\in\{1,2\}$,
+  2. 2.
+    $n_{i}=m_{i}+r_{i}$ for $i\in\{1,2\}$,
+  3. 3.
+    $(q,r_{1},r_{2})\in\delta(p,c_{1},c_{2})$, and
+  4. 4.
+    we assume if $c_{i}=0$ then $r_{i}\neq-1$ for $i\in\{1,2\}$.
+Essentially, the machine checks in every state whether the counters equal 0 and then changes the value of each counter by at most one per transition before entering a new state.
+A computation is a sequence of configurations.
+An accepting computation of $A$ is a sequence $C_{1},...,C_{n}\in(Q\times\mathbb{N}\times\mathbb{N})^{n}$ with $C_{1}=(q_{0},0,0)$, $C_{i}\vdash_{A}C_{i+1}$ for all $i\in\{1,...,n-1\}$, and $C_{n}\in F\times\mathbb{N}\times\mathbb{N}$ for some $n\in\mathbb{N}$ with $n>0$.
+We encode configurations of $A$ by assuming $Q=\{q_{0},...,q_{e}\}$ for some $e\in\mathbb{N}$ and defining a function $\operatorname{enc}$ :: $Q\times\mathbb{N}\times\mathbb{N}\rightarrow\{0,\#\}^{*}$ by
+$\operatorname{enc}(q_{i},m_{1},m_{2}):=0^{x+i}\#0^{c_{1}+y_{2}\cdot m_{1}}\#0^{c_{2}+y_{2}\cdot m_{2}}$
+for some numbers $x,c_{1},y_{2},c_{1},y_{2}\in\mathbb{N}$.
+The values for these numbers depend on the construction of the respective proofs and are not specified here.
+This is extended to encodings of computations by defining for every $n\geq 1$ and every sequence $C_{1},...,C_{n}\in Q\times\mathbb{N}\times\mathbb{N}$
+$\operatorname{enc}(C_{1},...,C_{n}):=\#\#\ \operatorname{enc}(C_{1})\ \#\#\ ...\ \#\#\ \operatorname{enc}(C_{n})\ \#\#.$
+Encodings of this kind are used to prove different undecidability results regarding the inclusion problem for erasing relational pattern languages for various considered relations, in particular the results in Subsection 4.2.
+For some nondeterministic 2-counter automaton without input $A$, define the set of encodings of accepting computations
+$\mathtt{ValC}(A):=\{\operatorname{enc}(C_{1},...,C_{n})\ |\ C_{1},...,C_{n}\text{ is an accepting computation of }A\}$
+and let $\mathtt{InvalC}(A)=\{0,\#\}^{*}\setminus\mathtt{ValC}(A)$.
+The emptiness problem for deterministic 2-counter-automata is undecidable (cf.
+e.g.
+[21, 29]), thus it is also undecidable whether a nondeterministic 2-counter automaton without input has an accepting computation [13, 23].
+That the emptiness problem for universal Turing machines is undecidable is a known fact.
+0.B.3.2 (The Full Proof Idea for $r_{rev}$):
+Proof.
+The proof is based on the construction by Freydenberger and Reidenbach [13].
+For better understanding for those who don’t know the construction and for a really detailed verification, we copied the proof of Theorem 2 of [13] and adapted it at several points.
 The main difference is the definition of $u$ and the definition of the delimiter.
-Why these adaptations work is shorter described in the proof of Theorem˜4.2 for those who already are familiar with the construction.
-The adaptations of Theorem˜4.1 and Theorem˜4.2 are very similar.
+Why these adaptations work is shorter described in the proof part for $r_{ab}$ for those who already are familiar with the construction.
+The adaptations for $r_{rev}$ and $r_{ab}$ are very similar.
+For now, assume $R=\{r_{rev}\}$.
 We begin with the case $|\Sigma|=2$, so let $\Sigma:=\{0,\#\}$.
@@
 $\beta_{A}:=x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}\#^{6}\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}\#^{6}\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}},$
-with, for all $i\in\{1,\ldots,\mu\}$, $\hat{\beta}_{i}:=x_{i_{1}}\ \gamma_{i}\ x_{i_{2}}\ \delta_{i}\ x_{i_{3}}$ and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{1}},x_{1_{2}},x_{1_{3}},x_{1_{4}},x_{1_{5}},\ldots,x_{\mu},x_{\mu^{\prime}},x_{\mu_{1}},x_{\mu_{2}},x_{\mu_{3}},x_{\mu_{4}},x_{\mu_{5}}$ are distinct variables with $(x_{i},x_{i^{\prime}}),(x_{i},x_{i_{1}}),(x_{i},x_{i_{2}}),(x_{i},x_{i_{3}}),(x_{i},x_{i_{4}}),(x_{i},x_{i_{5}})\in R$ and all
-$\gamma_{i},\delta_{i},\eta_{i}\in X^{*}$ are terminal-free patterns. The patterns $\gamma_{i}$ and $\delta_{i}$ shall be defined later; for now, we only mention:
+with, for all $i\in\{1,\ldots,\mu\}$, $\hat{\beta}_{i}:=x_{i_{1}}\ \gamma_{i}\ x_{i_{2}}\ \delta_{i}\ x_{i_{3}}$ and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{1}},x_{1_{2}},x_{1_{3}},x_{1_{4}},x_{1_{5}},\ldots,x_{\mu},x_{\mu^{\prime}},x_{\mu_{1}},x_{\mu_{2}},x_{\mu_{3}},x_{\mu_{4}},x_{\mu_{5}}$ are distinct variables with $(r_{rev},x_{i},x_{i^{\prime}})$, $(r_{rev},x_{i},x_{i_{1}})$, $(r_{rev},x_{i},x_{i_{2}})$, $(r_{rev},x_{i},x_{i_{3}})$, $(r_{rev},x_{i},x_{i_{4}})$, $(r_{rev},x_{i},x_{i_{5}})\in r_{\beta_{A}}$ and all $\gamma_{i},\delta_{i},\eta_{i}\in X^{*}$ are terminal-free patterns.
+The patterns $\gamma_{i}$ and $\delta_{i}$ shall be defined later; for now, we only mention:
   1. 1.
@@
 Karhumäki et al.
-[DBLP:journals/jacm/KarhumakiMP00]), this means that $\sigma$ satisfies $\pi_{i}$ if and only if the system consisting of the three equations $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ has a solution $\tau$.
+[24]), this means that $\sigma$ satisfies $\pi_{i}$ if and only if the system consisting of the three equations $\tau(\gamma_{i})=\sigma(x)$, $\tau(\delta_{i})=\sigma(y)$ and $\tau(\eta_{i})=u$ has a solution $\tau$.
 Later, we shall see that $L_{E}(\alpha_{A},r_{\alpha_{A}})\setminus L_{E}(\beta_{A},r_{\beta_{A}})$ exactly contains those $\sigma(\alpha_{A})$ for which $\sigma$ does not satisfy any of $\pi_{1}$ to $\pi_{\mu}$, and choose these predicates to describe $\mathtt{InvalC}(A)$.
@@
 $\delta_{1}:=\hat{y}_{1},$ $\delta_{2}:=\hat{y}_{2,1}\ \hat{z}_{2_{6}}\ \hat{y}_{2,2}$
-where $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2},\hat{z}_{1_{6}},\hat{z}_{1_{7}},\hat{z}_{1_{8}}$ and $\hat{z}_{2_{6}}$ are pairwise distinct variables with $(\hat{z}_{1_{1}},\hat{z}_{1_{6}}),(\hat{z}_{1_{1}},\hat{z}_{1_{7}}),(\hat{z}_{1_{1}},\hat{z}_{1_{8}}),(\hat{z}_{2_{1}},\hat{z}_{2_{6}})\in R$.
-Recall that
-$\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ for all $i$.
+where $y_{1,1},y_{1,2},y_{2},\hat{y}_{1},\hat{y}_{2,1},\hat{y}_{2,2},\hat{z}_{1_{6}},\hat{z}_{1_{7}},\hat{z}_{1_{8}}$ and $\hat{z}_{2_{6}}$ are pairwise distinct variables with $(r_{rev},\hat{z}_{1_{1}},\hat{z}_{1_{6}}),(r_{rev},\hat{z}_{1_{1}},\hat{z}_{1_{7}}),(r_{rev},\hat{z}_{1_{1}},\hat{z}_{1_{8}}),(r_{rev},\hat{z}_{2_{1}},\hat{z}_{2_{6}})\in r_{\beta_{A}}$.
+Recall that $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ for all $i$.
 It is not very difficult to see that $\pi_{1}$ and $\pi_{2}$ characterize the morphisms that are of bad form:
-Lemma 10
+Lemma 10.
 A substitution $\sigma\in H$ is of bad form if and only if $\sigma$ satisfies $\pi_{1}$ or $\pi_{2}$.
-Proof
+Proof.
 We begin with the only if direction.
@@
 Consequently, $\tau(\hat{z}_{1_{6}})=\tau(\hat{z}_{1_{7}})=\tau(\hat{z}_{1_{8}})=\#$ and thus, $\sigma(x)$ contains $^{3}$, and $\sigma$ is of bad form.
-Analogously, if $\sigma$ satisfies $\pi_{2}$, then $\sigma(y)$ contains the letter #\#, and $\sigma$ is of bad form.
+Analogously, if $\sigma$ satisfies $\pi_{2}$, then $\sigma(y)$ contains the letter #\#, and $\sigma$ is of bad form. ∎
 This allows us to make the following observation, which serves as the central part of the construction and is independent from the exact shape of $\pi_{3}$ to $\pi_{\mu}$:
-Lemma 11
+Lemma 11.
 For every substitution $\sigma\in H$, $\sigma(\alpha_{A})\in L_{E}(\beta_{A},r_{\beta_{A}})$ if and only if $\sigma$ satisfies one of the predicates $\pi_{1}$ to $\pi_{\mu}$.
-Proof
+Proof.
 We begin with the if direction.
@@
 Therefore, $\sigma$ satisfies the predicate $\pi_{i}$.
+∎
 Thus, we can select predicates $\pi_{1}$ to $\pi_{\mu}$ in such a way that $L_{E}(\alpha_{A},r_{\alpha_{A}})\setminus L_{E}(\beta_{A},r_{\beta_{A}})=\emptyset$ if and only if $\mathtt{ValC}(A)=\emptyset$ by describing $\mathtt{InvalC}(A)$ through a disjunction of predicates on $H$.
@@
 $\delta_{3}\$$:=\hat{\gamma}_{3,1^{\prime}}\hat{\gamma}_{3,2^{\prime}}\hat{\gamma}_{3,3^{\prime}},$
-where all of $\gamma_{3,1}$ to $\gamma_{3,4}$ and $\hat{\gamma}_{3,1}$ to $\hat{\gamma}_{3,3}$ and $\hat{\gamma}_{3,1^{\prime}}$ to $\hat{\gamma}_{3,3^{\prime}}$ are pairwise distinct variables with $(\hat{\gamma}_{3,1},\hat{\gamma}_{3,1^{\prime}}),(\hat{\gamma}_{3,2},\hat{\gamma}_{3,2^{\prime}}),(\hat{\gamma}_{3,3},\hat{\gamma}_{3,3^{\prime}})\in R$.
+where all of $\gamma_{3,1}$ to $\gamma_{3,4}$ and $\hat{\gamma}_{3,1}$ to $\hat{\gamma}_{3,3}$ and $\hat{\gamma}_{3,1^{\prime}}$ to $\hat{\gamma}_{3,3^{\prime}}$ are pairwise distinct variables with $(r_{rev},\hat{\gamma}_{3,1},\hat{\gamma}_{3,1^{\prime}}),(r_{rev},\hat{\gamma}_{3,2},\hat{\gamma}_{3,2^{\prime}}),(r_{rev},\hat{\gamma}_{3,3},\hat{\gamma}_{3,3^{\prime}})\in r_{\beta_{A}}$.
 Clearly, if some $\sigma\in H$ satisfies $\pi_{3}$, $\sigma(y)$ is a concatenation of three (possibly empty) reversed factors of $\sigma(x)$.
@@
 For relational patterns $(p,v_{R})$ with $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ and $v_{R}:=\{(r_{rev},\hat{x}_{i,j},\hat{x}_{i,j^{\prime}})\ |\ i\in\{1,2,3\};j,j^{\prime}\in\mathbb{N}\}$, we define $S(p,v_{R}):=\{\rho(p)\ |\ \rho\in S\cap H_{v_{R}}\}$.
-Definition 2
+Definition 2.
 A predicate $\pi:G_{v_{R}}\rightarrow\{0,1\}$ is called a simple predicate if there exists a pattern $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ and languages $L_{1},L_{2}\in\{\Sigma^{*},\{\epsilon\}\}$ such that $\sigma$ satisfies $\pi$ if and only if $\sigma(x)\in L_{1}\ S(p,v_{R})\ L_{2}$.
@@
 Using $\pi_{3}$, our construction is able to express all simple predicates:
-Lemma 12
+Lemma 12.
 For every simple predicate $\pi_{S}$ over variables from $X^{\prime}$, there exists a predicate $\pi$ defined by terminal-free patterns $\gamma,\delta,\eta$ such that for all substitutions $\sigma\in G_{v_{R}}$:
@@
     if $\sigma$ satisfies $\pi$, then $\sigma$ also satisfies $\pi_{S}$.
-Proof
-We first consider the case of $L_{1}=L_{2}=\Sigma^{*}$.
-Assume that $\pi_{S}$ is a simple predicate, and $p\in(\Sigma\ \cup\ X^{\prime})^{*}$ is a pattern such that $\sigma\in G_{v_{R}}$ satisfies $\pi_{S}$ if and only if $\sigma(x)\in L_{1}\ S(p,v_{R})\ L_{2}$.
-Then define $\gamma:=y_{1}p^{\prime}y_{2}$, where $p^{\prime}$ is obtained from $p$ by replacing all $k,k\in\mathbb{N}_{0}$ occurrences of 0 with new variables $\tilde{z}_{j},j\in[k]$ with $(\tilde{z}_{1},\tilde{z}_{j})\in R$ for all $j\in[k]$ and all $g,g\in\mathbb{N}_{0}$ occurrences of #\# with different variables $\ddot{z}_{j},j\in[g]$ with $(\ddot{z}_{1},\ddot{z}_{j})\in R$ for all $j\in[g]$, while leaving all present elements of $X^{\prime}$ unchanged.
-Furthermore, let $\delta:=\hat{x}_{1,j}\hat{x}_{2,j}\hat{x}_{3,j}\hat{y}$ with $j\in\mathbb{N}$ such that $\hat{x}_{1,j},\hat{x}_{2,j},\hat{x}_{3,j}$ are new variables and (in order to stay consistent with the $\eta_{i}$ appearing in $\beta_{A}$) $\eta:=\tilde{z}\ \ddot{z}_{g+1}\ \ddot{z}_{g+2}\ \ddot{z}_{g+3}\ \ddot{z}_{g+4}\ \ddot{z}_{g+5}\ \tilde{z}^{\prime}$ with $(\tilde{z}_{1},\tilde{z}),(\tilde{z}_{1},\tilde{z}^{\prime}),(\ddot{z}_{1},\ddot{z}_{g+1}),(\ddot{z}_{1},\ddot{z}_{g+2}),(\ddot{z}_{1},\ddot{z}_{g+3}),(\ddot{z}_{1},\ddot{z}_{g+4}),(\ddot{z}_{1},\ddot{z}_{g+5})\in R$.
+Proof.
+[…]
 Now, assume that $\sigma\in G_{v_{R}}$ satisfies $\pi_{S}$.
@@
 […]
-$(\tilde{z}_{1},\tilde{z}),(\tilde{z}_{1},\tilde{z}^{\prime}),(\ddot{z}_{1},\ddot{z}_{g+1}),(\ddot{z}_{1},\ddot{z}_{g+2}),(\ddot{z}_{1},\ddot{z}_{g+3}),(\ddot{z}_{1},\ddot{z}_{g+4}),(\ddot{z}_{1},\ddot{z}_{g+5})\in R$ and u=0#####0u=0\#\#\#\#\#0, $\tau(\tilde{z})=\tau(\tilde{z}^{\prime})=0$ and $\tau(\ddot{z}_{j})=\#$ for all $j\in\{g+1,\ldots,g+5\}$ must hold.
-By definition $\tau(y_{1}),\tau(y_{2})\in\Sigma^{*}$.
-If we define $\rho(\hat{x}_{i,j}):=\tau(\hat{x}_{i,j})$ for all $i\in\{1,2,3\},j\in\mathbb{N}$ such that $\tau(\hat{x}_{i,j})$ is defined and $\rho(\hat{x}_{i,j}):=\rho(\hat{x}_{i,j^{\prime}})$ for all other $i\in\{1,2,3\},j\in\mathbb{N}$ for one arbitrary, specific $j^{\prime}\in\mathbb{N}$ such that $\rho(\hat{x}_{i,j^{\prime}})$ is defined, we see that $\sigma(x)\in L_{1}S(p,v_{R})L_{2}$ holds.
-Thus, $\sigma$ satisfies $\pi_{S}$ as well.
-The other three cases for choices of $L_{1}$ and $L_{2}$ can be handled analogously by omitting $y_{1}$ or $y_{2}$ as needed. Note that this proof also works in the case $p=\epsilon$.
+The other three cases for choices of $L_{1}$ and $L_{2}$ can be handled analogously by omitting $y_{1}$ or $y_{2}$ as needed. Note that this proof also works in the case $p=\epsilon$. ∎
 Roughly speaking, if $\sigma$ does not satisfy $\pi_{3}$, then $\sigma(y)$ (which is in $0^{*}$, due to $\sigma\in G_{v_{R}}$) is long enough to provide building blocks for simple predicates using variables from $X$.
@@
 But this problem can be avoided by choosing $\sigma(y)$ long enough to leave $\pi_{3}$ unsatisfied, and the following holds:
-Lemma 13
+Lemma 13.
 A word $w\in\Sigma^{*}$ is of good structure if and only if there exists a substitution $\sigma\in H_{v_{R}}$ with $\sigma(x)=w$ such that $\sigma$ satisfies none of the predicates $\pi_{1}$ to $\pi_{13}$.
-Proof
+Proof.
 We begin with the if direction.
@@
 By looking at the cases used above to define $\pi_{4}$ to $\pi_{13}$, we see that none of these predicates can be satisfied.
+∎
 For every $w$ of good structure, there exist uniquely determined
@@
 We conclude that $A$ has an accepting computation iff $L_{E}(\alpha_{A},r_{\alpha_{A}})$ is not a subset of $L_{E}(\beta_{A},r_{\beta_{A}})$.
-Therefore, any algorithm deciding the inclusion problem of Theorem˜4.1 can be used to decide whether a nondeterministic 2-counter automaton without input has an accepting computation.
-As this problem is known to be undecidable, the inclusion problem of Theorem˜4.1 is also undecidable.
+Therefore, any algorithm deciding the inclusion problem under $r_{rev}$ can be used to decide whether a nondeterministic 2-counter automaton without input has an accepting computation.
+As this problem is known to be undecidable, the inclusion problem under $r_{rev}$ is also undecidable.
 This proof can be extended to larger (finite) alphabets.
@@
 instead of $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$, where all $z_{i},z_{i^{\prime}},\hat{z}_{i_{j}},\ddot{z}_{i,k_{j}}$ are pairwise different variables with
-$(z_{i},z_{i^{\prime}}),(\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(\hat{z}_{i_{1}},\hat{z}_{i_{3}}),(\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(\hat{z}_{i_{1}},\hat{z}_{i_{5}}),(\ddot{z}_{i,k_{1}},\ddot{z}_{i,k_{2}})\in R.$
+$(r_{rev},z_{i},z_{i^{\prime}}),(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{2}}),(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{3}}),$
+$(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{4}}),(r_{rev},\hat{z}_{i_{1}},\hat{z}_{i_{5}}),(r_{rev},\ddot{z}_{i,k_{1}},\ddot{z}_{i,k_{2}})\in r_{\beta_{A}}.$
 Referring to the new shape of $u$, we can make the following observation:
-Lemma 14
+Lemma 14.
 Let $n\geq 2$, $\{x_{1},x_{1}^{\prime},x_{2_{1}},x_{2_{2}},x_{2_{3}},x_{2_{4}},x_{2_{5}},x_{3_{1}},x_{3,2},\ldots,x_{n_{1}},x_{n_{2}}\}\subset X$ and $\{a_{1},\ldots,a_{n}\}\subseteq\Sigma$. If
 $\alpha=x_{1}\ x_{2_{1}}\ x_{2_{2}}\ x_{2_{3}}\ x_{2_{4}}\ x_{2_{5}}\ x_{3_{1}}\ x_{3_{2}}\ldots x_{n_{1}}\ x_{n_{2}}\ x_{1}^{\prime}$
-with $(x_{1},x_{1}^{\prime}),(x_{2_{1}},x_{2_{2}}),(x_{2_{1}},x_{2_{3}}),(x_{2_{1}},x_{2_{4}}),(x_{2_{1}},x_{2_{5}}),(x_{i_{1}},x_{i_{2}})\in R$ for all $i\in\{3,\ldots,n\}$ and there is a morphism $\sigma:X^{*}\rightarrow\Sigma^{*}$ with
-$\sigma(\alpha)=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$, then $\sigma(x_{1})=\sigma(x_{1}^{\prime})=a_{1}$ and $\sigma(x_{i_{j}})=a_{i}$ for each $i\in\{1,\ldots,n\}$.
-Proof
+with $(r_{rev},x_{1},x_{1}^{\prime})$, $(r_{rev},x_{2_{1}},x_{2_{2}})$, $(r_{rev},x_{2_{1}},x_{2_{3}})$, $(r_{rev},x_{2_{1}},x_{2_{4}})$, $(r_{rev},x_{2_{1}},x_{2_{5}})$, $(r_{rev},x_{i_{1}},x_{i_{2}})\in r_{\beta_{A}}$ for all $i\in\{3,\ldots,n\}$ and there is a morphism $\sigma:X^{*}\rightarrow\Sigma^{*}$ with $\sigma(\alpha)=a_{1}(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}a_{1}$, then $\sigma(x_{1})=\sigma(x_{1}^{\prime})=a_{1}$ and $\sigma(x_{i_{j}})=a_{i}$ for each $i\in\{1,\ldots,n\}$.
+Proof.
 Assume
@@
 $\sigma(x_{2_{1}}\ x_{2_{2}}\ x_{2_{3}}\ x_{2_{4}}\ x_{2_{5}}\ x_{3_{1}}\ x_{3_{2}}\ldots x_{n_{1}}\ x_{n_{2}})=(a_{2})^{5}(a_{3})^{2}\ldots(a_{n})^{2}$
-must hold. Now, $\sigma(x_{2_{j}})=a_{2}$ must hold for every $j\in\{1,2,3,4,5\}$ since
-$(x_{2_{1}},x_{2_{2}}),(x_{2_{1}},x_{2_{3}}),(x_{2_{1}},x_{2_{4}}),(x_{2_{1}},x_{2_{5}}),(x_{i_{1}},x_{i_{2}})\in R$ for all $i\in\{3,\ldots,n\}$.
+must hold.
+Now, $\sigma(x_{2_{j}})=a_{2}$ must hold for every $j\in\{1,2,3,4,5\}$ since $(r_{rev},x_{2_{1}},x_{2_{2}})$, $(r_{rev},x_{2_{1}},x_{2_{3}})$, $(r_{rev},x_{2_{1}},x_{2_{4}})$, $(r_{rev},x_{2_{1}},x_{2_{5}})$, $(r_{rev},x_{i_{1}},x_{i_{2}})\in r_{\beta_{A}}$ for all $i\in\{3,\ldots,n\}$.
 Thus, we also get that $\sigma(x_{i_{1}})=\sigma(x_{i_{2}})=a_{i}$ for every $i\in\{3,\ldots,n\}$.
+∎
 Lemma˜14 allows $\pi_{\mu+1}$ to $\pi_{\mu+2n}$ to be analogously constructed to $\pi_{2}$. To this end, we define
@@
 for each $i\in[n]$.
-Again, all $y_{j,k},\hat{y}_{j,k},\ddot{z}_{j,k}^{\prime}$ are pairwise different variables with $(\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i-1,i}^{\prime}),(\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i,i}^{\prime})\in R$.
-Now Lemma˜10 applies (mutatis mutandis) as for binary alphabets, and since all substitutions of good form behave for $\Sigma$ as for the binary alphabet, we can use the very same predicates and the same reasoning as before to prove undecidability of the inclusion problem of Theorem˜4.1.
-This concludes the proof of Theorem˜4.1.
-Appendix 0.C Additional Content for Equivalence Problem
-Proof of Lemma˜4:
-Proof
-Assume $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$.
-First of all, we know $|\alpha|=|\beta|$ as otherwise, if w.l.o.g.
-$|\alpha|>|\beta|$, then no shortest word $w\in L_{NE}(\beta,[r_{\beta}])$ obtained by substitutions $h$ where variables are replaced with letters can be obtained from $(\alpha,r_{\alpha})$.
-So, from now on, assume $|\alpha|=|\beta|$.
-Suppose $\alpha\neq\beta$.
-Then there exists a first position $i\in[|\alpha|]$ such that $\alpha[i]\neq\beta[i]$.
-First, if $\alpha[i],\beta[i]\in\Sigma$ are both terminal letters, then using a substitution $h$ that replaces all variables with letters in $\alpha$ and $\beta$ always results in words $h(\alpha)$ and $h(\beta)$ for which $h(\alpha)[i]\neq h(\beta)[i]$, a contradiction.
-So, assume w.l.o.g.
-that $\alpha[i]\in\Sigma$ is a terminal letter and $\beta[i]\in X$ is a variable.
-Then, again, any substitution $h\in H_{r_{\beta}}$ that replaces all variables in $\beta$ with letters and sets $h(\beta[i])$ to a letter which is not $\alpha[i]$ produces a word that cannot be in $L_{NE}(\alpha,[r_{\alpha}])$, another contradiction to $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$.
-As we assume all variables to be introduced in an increasing manner, $\alpha=\beta$.
+Again, all $y_{j,k},\hat{y}_{j,k},\ddot{z}_{j,k}^{\prime}$ are pairwise different variables with $(r_{rev},\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i-1,i}^{\prime}),(r_{rev},\ddot{z}_{i,1_{1}},\ddot{z}_{\mu+2i,i}^{\prime})\in R$.
+Now Lemma˜10 applies (mutatis mutandis) as for binary alphabets, and since all substitutions of good form behave for $\Sigma$ as for the binary alphabet, we can use the very same predicates and the same reasoning as before to prove undecidability of the inclusion problem under $r_{rev}$.
+This concludes the proof for $r_{rev}$. ∎
+0.B.3.3 (The Full Proof Idea for $r_{ab}$)
+Proof.
+This proof is analogous to the one for $r_{rev}$.
+Therefore, we only describe how to get to the analogy and refer for a detailed proof to the above part.
+Let $A$ be a nondeterministic 2-counter automaton.
+Consider
+$\alpha_{A}:=v\ v\ \#^{6}\ v\ x\ v\ y\ v\ \#^{6}\ v\ u\ v$
+where $x,y$ are distinct variables, $v=0\#^{3}0$ and $u=0\#^{5}0$ and
+$\beta_{A}:=x_{1}x_{1^{\prime}}\ \ldots x_{\mu}x_{\mu^{\prime}}\#^{6}\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}\#^{6}\ddot{\beta_{1}}\ldots\ddot{\beta_{\mu}}$
+with, for all $i\in\{1,\ldots,\mu\}$, and $\ddot{\beta_{i}}:=x_{i_{4}}\ \eta_{i}\ x_{i_{5}}$, where $x_{1},x_{1^{\prime}},x_{1_{4}},x_{1_{5}},\ldots$ are distinct variables with $(r_{ab},x_{i},x_{i^{\prime}})$, $(r_{ab},x_{i},x_{i_{1}})$, $(r_{ab},x_{i},x_{i_{2}})$, $(x_{i},x_{i_{3}}),(r_{ab},x_{i},x_{i_{4}})$, $(r_{ab},x_{i},x_{i_{5}})\in r_{\beta_{A}}$ and all $\eta_{i}\in X^{*}$ are terminal-free patterns for some $\mu$ and some $\hat{\beta}_{i}$ as for $r_{rev}$.
+We have $\eta_{i}:=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ with distinct variables $z_{i},z_{i^{\prime}},\hat{z}_{i_{1}},\hat{z}_{i_{2}},\hat{z}_{i_{3}},\hat{z}_{i_{4}},\hat{z}_{i_{5}}$ and $(r_{ab},z_{i},z_{i^{\prime}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{2}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{3}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{4}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in r_{\beta_{A}}$.
+The main concept of the proof is to show that iff $L_{E}(\alpha_{A},r_{\alpha_{A}})\subseteq L_{E}(\beta_{a},r_{\beta_{A}})$ then $A$ has no accepting computation.
+All possible non accepting computations are listed in the $\hat{\beta}_{1}\ldots\hat{\beta}_{\mu}$ part where each $\hat{\beta}_{i},i\in\{1,\ldots,\mu\}$ encodes one non accepting computation together with selecting variables $x_{i_{j}}$ for $j\in{1,2,3}$.
+The part of $\beta_{A}$ and $\alpha_{A}$ before the first $^{6}$ ensures that exactly one pair $x_{i},x_{i}^{\prime},i\in\{1,\ldots,\mu\}$ is selected, resulting in that exactly one non accepting computation is selected and that exactly one $\ddot{\beta}_{i},i\in\{1,\ldots,\mu\}$ is selected.
+With this, we get that $\eta_{i},i\in\{1,\ldots,\mu\}$ matches $u$ and thus, that $z_{i}=0$ and $\hat{z}_{i_{1}}=\#$.
+We use these two variables to encode the 0 and the #\# in the non accepting computation encodings.
+If we can show that this selecting process also works in the setting $R=r_{ab}$ and that we get two variables with which we can encode the letters 0 and #\#, then this proof also works in this setting.
+For two variables $x,y$, we have if $x=0$ (or x=#x=\#) and $(x,y)\in r_{ab}$, then $y=0$ (or y=#y=\# respectively).
+Thus, we can model with the relation $r_{ab}$ identical variables if we know a one letter substitution of one variable.
+[…]
+$x_{i}$ $x_{i^{\prime}}$ works?
+0 0 $00\neq 0\#$
+0#0\# 0#0\# or #0\#0 $0#$, $0\#,\ 0\#^{2}0\neq 0\#^{3}$
+$0\#^{2}$ $0\#^{2}$ or #0#\#0\# or ##0\#\#0 $0\#^{2}0\#^{2},\ 0\#^{3}0\#,\ 0\#^{4}0\neq 0\#^{3}00$
+$0\#^{3}$ $0\#^{3}$ or $0\#^{2}$ or $^{2}0\#$ or $^{3}0$ $0\#^{3}0\#^{3},\ 0\#^{3}\#0\#^{2},\ 0\#^{5}0\#,\ 0\#^{6}30\neq 0\#^{3}00\#^{2}$
+$0\#^{3}0$ $0\#^{3}0$ $0\#^{3}00\#^{3}0=v\ v$
+Table 3: Possibilities for the first match with vvv\ v
+Thus, the only possibility is that for exactly one $i\in\{1,\ldots,\mu\}$ the variables $x_{i}$ and $x_{i^{\prime}}$ get substituted with $v$ and all other variables $x_{j},x_{j^{\prime}}$ for $j\neq i,j\in\{1,\ldots,\mu\}$ get substituted with $\epsilon$.
+This together with the matching from vuvv\ u\ v to $\ddot{\beta}_{1}\ldots\ddot{\beta}_{\mu}$ implies that $\eta_{i}$ needs to match $u$.
+For this matching between $\eta_{i}=z_{i}\ \hat{z}_{i_{1}}\ \hat{z}_{i_{2}}\ \hat{z}_{i_{3}}\ \hat{z}_{i_{4}}\ \hat{z}_{i_{5}}\ z_{i^{\prime}}$ and $u=0\#^{5}0$, we get the following possibilities.
+Remember that $(r_{ab},z_{i},z_{i^{\prime}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{2}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{3}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{4}})$, $(r_{ab},\hat{z}_{i_{1}},\hat{z}_{i_{5}})\in r_{\beta_{A}}$.
+$z_{i}$ $\hat{z}_{i_{1}}$ works?
+$\epsilon$ 0 or #\# $0^{5},\ \#^{5}\neq 0\#^{5}0$
+0 #\# $0\#^{5}0=0\#^{5}0=u$
+0#0\# #\# $0\#\#^{5}\#0\neq 0\#^{5}0$
+0##0\#\# #\# $0\#^{2}\#^{5}\#^{2}0\neq 0\#^{5}0$
+0###0\#\#\# $\epsilon$ $0\#^{6}0\neq 0\#^{5}0$
+Table 4: Possibilities for the match with $u$
+Thus, the only possibility is that $z_{i}$ and $z_{i^{\prime}}$ get substituted with 0 and $\hat{z}_{i_{k}},k\in\{1,2,3,4,5\}$ get substituted with #\#.
+With this, we can encode the 0 and the #\# and thus, the proof for $r_{rev}$ works in this setting analogously.
 ∎
-Proof of Lemma˜5:
-Proof
-($\Leftarrow$): Assume $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$. Then $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$. So, we trivially get $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$.
-($\Rightarrow$) Assume $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$.
-By Lemma 4 we have $\alpha=\beta$.
-The relational constraints $r_{\alpha}]$ and $r_{\beta}]$ imply blocks of equivalently substituted variables under $\tilde{r}$.
-Suppose $r_{\alpha}]\neq[r_{\beta}]$.
-Let $\alpha=\alpha_{p}x_{i}\alpha_{s}=\beta$ for some $\alpha_{p},\alpha_{s}\in\mathtt{RelPat}_{\Sigma,R}$ such that $x_{i}$ is the first variable for which there is a difference in the relations, i.e., we get that w.l.o.g.
-$x_{i}\in\operatorname{var}(\alpha)$ is related to some other variable $x_{j}\in\operatorname{var}(\alpha)$ in $r_{\alpha}]$ but not in $r_{\beta}]$ (note that $\operatorname{var}(\alpha)=\operatorname{var}(\beta)$).
-So, there exists $(\tilde{r},x_{i},x_{j})\in[r_{\alpha}]$ such that $(\tilde{r},x_{i},x_{j})\notin[r_{\beta}]$.
-Select $h\in H_{r_{\beta}}$ such that $h$ substitutes all variables in $\beta$ to some letter, i.e., $|h(x)|=1$ for all $x\in\operatorname{var}(\beta)$.
-In particular, let $h(x_{i})=h(x_{k})=\mathtt{a}$, for some $\mathtt{a}\in\Sigma$ and all $(\tilde{r},x_{i},x_{k})\in[r_{\beta}]$.
-Finally, set $h(x_{j})=h(x_{k^{\prime}})=\mathtt{b}\neq\mathtt{a}=h(x_{i})$, for all other variables $x_{k^{\prime}}\in\operatorname{var}(\beta)$ which are not already set to $h(x_{i})=\mathtt{a}$ and some letter $\mathtt{b}\in\Sigma$ with $\mathtt{a}\neq\mathtt{b}$.
-Then, we have that $h(\beta)\in L_{NE}(\beta,[r_{\beta}])$ but $h(\beta)\notin L_{NE}(\alpha,[r_{\alpha}])$ as there cannot exist any substitution $h^{\prime}\in H_{r_{\alpha}}$ with $h^{\prime}(x_{i})\neq h^{\prime}(x_{j})$ as $(\tilde{r},x_{i},x_{j})\in[r_{\alpha}]$ and by the assumption that $\tilde{r}$ is an equivalence relation.
-This is a contradiction to $L_{NE}(\alpha,[r_{\alpha}]])=L_{NE}(\beta,[r_{\beta}])$.
-So, we must have $r_{\alpha}]=[r_{\beta}]$.
-∎
-Proof of Theorem˜5.1:
-Proof
-($\Rightarrow$): Assume $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
-As $\tilde{r}$ is an equivalence relation, we get by Lemma 3 that $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$ and $L_{NE}(\beta,r_{\beta})=L_{NE}(\beta,[r_{\beta}])$.
-Hence, $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$, and we get by Lemma 5 that $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
-Thus, we have $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
-($\Leftarrow$): Assume $(\alpha,[r_{\alpha}])=(\beta,[r_{\beta}])$.
-Then, by Lemma 5 we have $L_{NE}(\alpha,[r_{\alpha}])=L_{NE}(\beta,[r_{\beta}])$ and by Lemma 3 we get that $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ using the same arguments from the other direction.
-∎
-Proof of Corollary˜3:
-Proof
-First, we notice that all relations in $\{r_{=},r_{ab},r_{con^{+}}\}$ are equivalence relations that are antisymmetric on elements of length $1$.
-Now, we normalize $(\alpha,r_{\alpha})$ and $(\beta,r_{\beta})$ such that all variables are introduced with an index in increasing manner, starting with $x_{1}$ in both relational patterns.
-Then, we check in linear time whether $\alpha=\beta$.
-If that is the case, we check whether $r_{\alpha}]=[r_{\beta}]$ by constructing the sets representing reflexive, symmetric, and transitive closure of related variables.
-This can be done in linear time in the lengths of the patterns by scanning them and filling up the resulting sets.
-The total number of elements in these sets will not exceed the total number of variables in the patterns.
-If these sets are equal for both relational patterns, we return true, applying Theorem˜5.1.
-In any other case, we return false.
-∎
+Appendix 0.C Additional Content for the Membership Problem
+0.C.1 Proof of Lemma˜5
+Proof.
+This result follows from known reductions, e.g., [28], for all considered cases in this Lemma.
+For reference, we recall the construction formally in the relational pattern framework and state why it translates over.
+First, we recall positive 1-in-3-SAT.
+A formula $\varphi$ in positive 3-CNF (conjunctive normal form) is a set of clauses $\varphi=\{C_{1},...,C_{n}\}$, for some $n\in\mathbb{N}$, where each clause $C_{i}$ is a set of 3 non-negated literals $C_{i}=\{X_{i,1},X_{i,2},X_{i,3}\}$, where each $X_{i,j}\in X_{p}$ ($X_{p}$ being the set of all propositional variables).
+A positive 3-CNF formula $\varphi$ is 1-in-3 satisfiable if there exists an assignment $\beta:X_{p}\rightarrow\{0,1\}$ that sets exactly one literal per clause to $1$.
+Assume an arbitrary positive 3-CNF formula $\varphi=\{C_{1},...,C_{n}\}$. Construct the word $w=(1\#)^{n-1}1$ over the binary alphabet $\{1,\#\}$. Next, construct the pattern
+$\alpha=x_{1,1}x_{1,2}x_{1,3}\#...\#x_{n,1}x_{n,2}x_{n,3},$
+for independent variables $x_{i,j}\in X$ where $i,j\in\mathbb{N}$.
+Finally, construct the relational constraints $r_{\alpha}$ by adding, for each $X_{i,j}$ and $X_{i^{\prime},j^{\prime}}$ occurring in $\varphi$ where $X_{i,j}=X_{i^{\prime},j^{\prime}}$ are the same variable, the pairs $(x_{i,j},x_{i^{\prime},j^{\prime}}),(x_{i^{\prime},j^{\prime}},x_{i,j})\in r_{\alpha}$.
+Notice that $|w|_{\#}=|\alpha|_{\#}=n-1$.
+Hence, each $r_{\alpha}$-valid substitution $h$ must set all variables either to $1$ or to $\varepsilon$ to obtain $h(\alpha)=w$.
+Now, we see that if $\varphi$ is 1-in-3-satisfiable, then there exists an assignment of variables $\beta$ that sets exactly one variable per clause to $1$.
+Based on $\beta$, construct a substitution $h$ where $h(x_{i,j})=1$ iff $\beta{X_{i,j}}=1$.
+Otherwise set $h(x_{i,j})=\varepsilon$.
+As each clause only has one variable set to true by $\beta$, we obtain $h(\alpha)=w$.
+For the other direction, notice that we can analogously construct an assignment $\beta$ out of a $r_{\alpha}$-valid substitution $h$ with $h(\alpha)=w$.
+This concludes the erasing case.
+For the non-erasing case, notice that we can construct $(\alpha,r_{\alpha})$ in exactly the same manner and construct a word $w=(1\#)^{4}\#1^{4}$.
+As the non-erasing case forces each variable to have at least length 1 and all variables must be substituted to multiples of the letter $1$, we can use length $2$ substitutions to represent the positive literals.
+The argument follows analogously.
+  1. 1.
+    If $R=\{r_{|w|}\}$, then variables have to be substituted with equal length.
+As any substitution resulting in $w$ substituted each variable only over unary words over the letter $1$, having equal length also means being the equal word.
+This concludes this relation for the E and NE cases.
+  2. 2.
+    If $R=\{r_{ssq}\}$, then relating two variables $x$ and $y$ in both directions, i.e., putting $(r_{ssq},x,y)\in r_{\alpha^{\prime}}$ as well as $(r_{ssq},y,x)\in r_{\alpha^{\prime}}$ results in some substitution $h\in H$ to be valid if and only if $h(x)=h(y)$.
+This concludes this relation for the E and NE cases.
+  3. 3.
+    If $R=\{r_{ab}\}$ the same argument as for $r_{|w|}$ holds for the E and NE cases.
+  4. 4.
+    If $R=\{r_{perm}\}$ the same argument as for $r_{|w|}$ holds for the E and NE cases.
+  5. 5.
+    If $R=\{r_{rev}\}$ the same argument as for $r_{|w|}$ holds for the E and NE cases.
+  6. 6.
+    If $R=\{r_{*}\}$ the the same argument as for $r_{ssq}$ holds for the E and NE cases.
+  7. 7.
+    If $R=\{r_{com^{+}}\}$ then, in the erasing case, substituting one variable to the letter $1$, as $\varepsilon$ is not allowed, forces related variables to also be substituted to the letter $1$ for a $r_{\alpha}$-valid substitution $h$ to result in $h(\alpha)=w$.
+  8. 8.
+    If $R=\{r_{\varepsilon_{=}}\}$ the sane argument as for $r_{com^{+}}$ holds for the E case.
+This concludes this proof. ∎
+0.C.2 Proof of Theorem˜5.1
+Proof.
+Indeed, NP-containment of all considered cases follows from Proposition˜2.
+For most cases, NP-hardness has been confirmed in 5.
+What is left to show is that, first, in the case of non-erasing pattern languages, NP-hardness also holds for the relations $r_{com^{+}}$ and $r_{com^{*}}$.
+Indeed, in the NE case, these two relations result in the same languages, hence they can be considered together.
+Second, in the erasing case, the relation $r_{com^{*}}$ is still open.
+The rest of this proof provides two reductions, showing NP-hardness for the remaining two cases.
+(Non-Erasing ($r_{com^{+}}$$r_{com^{*}}$): W.l.o.g.
+we consider $R=\{r_{com^{+}}\}$.
+We reduce the general 3-SAT problem to the respective membership problem.
+Let $\varphi=\{C_{1},...,C_{n}\}$ be a 3-CNF formula, for some $n>0$, where $C_{i}=\{X_{i,1},X_{i,2},X_{i,3}\}$ for clauses $C_{i}$ and propositional variables $X_{i,j}\in\{X_{1},...,X_{m},\overline{X_{1}},...,\overline{X_{m}}\}$.
+Notice that in this variant of the problem, negated variables may occur as literals.
+First, we construct the word
+$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
+where $s_{i}=1\#1$ and $t_{j}=1^{10}\#1^{10}\#1^{10}$, for each $1<i\leq n$ and $1\leq j\leq m$. Next, we construct the regular pattern
+$\alpha=\#\#\ u_{0}v_{0}\ \#\#\ ...\ \#\#\ u_{m}v_{m}\ \#\#\ \alpha_{1}\ \#\#\ ...\ \#\#\ \alpha_{n}\ \#\#$
+where $\alpha_{i}=z_{i,1}\ x_{i,1}\ z_{i,2}\ x_{i,2}\ z_{i,3}\ x_{i,3}\ z_{i,4}$, for $1\leq i\leq n$ and variables $z_{i,1}$ to $z_{i,4}$ as well as $x_{i,1}$ to $x_{i,3}$.
+Notice that the pattern $\alpha$ is indeed regular and that each variable occurs only once.
+Finally, we construct the regular constraints $r_{\alpha}$.
+For all non-negated variables $X_{k}\in\{X_{1},...,X_{m}\}$ and literals equal to that variable $X_{i,j}=X_{k}$, we add the triple $(r_{com^{*}},u_{k},x_{i,j})\in r_{\alpha}$.
+Analogously, for all negated variables $\overline{X_{k}}\in\{\overline{X_{1}},...,\overline{X_{m}}\}$, we add the triple $(r_{com^{*}},v_{k},x_{i,j})\in r_{\alpha}$.
+In conclusion, there exist two parts.
+First, each factor $s_{k}$, $k\in[m]$, in $w$ has to be matched against $u_{k}v_{k}$ in $\alpha$, and each factor $t_{i}$, $\in[n]$ has to be matched against $\alpha_{i}$.
+We proceed by proving the arguments.
+($\Rightarrow$:) Assume there exists a satisfying assignment of variables $\beta$ for $\varphi$.
+Let $h$ be a substitution with the following properties: For every variable $X_{k}$, $k\in[m]$, where $\beta(X_{k})=1$ (thus $\beta(\overline{(}X_{k}))=0$), we set $h(u_{k})=1$, $h(v_{k})=\#1$ and, for every literal $X_{i,j}$ with $X_{i,j}=X_{k}$, we set $h(x_{i,j})=1$, and, for every literal with $X_{i,j}=\overline{(}X_{k})$, we set $h(x_{i,j})=\#1$.
+Analogously, if $\beta(X_{k})=0$ and $\beta(\overline{X_{k}})=1$, we set $h(u_{k})=1\#$ and $h(v_{k})=1$ and the $x_{i,j}$ variables correspondingly.
+As all variables are assigned some value in $\beta$, all variables of the form $x_{i,j}$ in $\alpha$ are substituted by some value (either $1$ or 1#1\# or #1\#1).
+For each $\alpha_{i}$, $i\in[n]$, we can now set the surrounding independent variables $z_{i,1}$ to $z_{i,4}$ to non-empty words such that $h(\alpha_{i})=1^{1}0\#1^{1}0\#1^{1}0$.
+As there is at least one variable $x_{i,j}$ in each $\alpha_{i}$ that is substituted by a word not containing #\# (as $\beta$ is a satisfying assignment), we will not obtain too many #\# by the substitutions for $x_{i,j}$s alone.
+Additionally, as there are enough $1$s in between each #\# in $w$, we can find a non-empty substitution for each $r_{i,j}$ variable.
+As all variables related by constraints in $r_{\alpha}$ are substituted by equal words, conforming the relation $r_{com^{+}}$, the chosen substitution is $r_{\alpha}$-valid.
+We notice that each $h(u_{k}v_{k})=1\#1$, each $h(\alpha_{i})=1^{1}0\#1^{1}0\#1^{1}0$, and therefore $h(\alpha)=w$.
+($\Leftarrow$:) Now assume there exists a $r_{\alpha}$-valid substitution $h$ such that $h(\alpha)=w$.
+We notice that no variable can be substituted to contain the factor ##\#\#, as otherwise $|h(\alpha)|_{\#\#}>|w|_{\#\#}$.
+Hence, we must have $h(u_{k}vk)=s_{k}=1\#1$ as well as $h(\alpha_{i})=t_{i}=1^{1}0\#1^{1}0\#1^{1}0$.
+As this is the non-erasing case, for each $k\in[m]$, either $h(u_{k})$ or $h(v_{k})$ contains the symbol #\# while the other is substituted only by $1$.
+As each $h(\alpha_{i})$ only contains two #\# symbols, there is at least one variable of the form $x_{i,j}$ which is substituted by a unary word in $\{1\}^{+}$.
+As each such variable is related to some $u_{k}$ or $v_{k}$ respectively, this can only be the case if both are substituted by unary words in $\{1\}^{+}$.
+Construct an assignment $\beta$ that sets, for each $k\in[m]$, $\beta(X_{k})=1$ and $\beta(\overline{X_{k}})=0$ if $h(u_{k})=1$ (thus $h(v_{k})=\#1$), and, otherwise, set $\beta(X_{k})=0$ and $\beta(\overline{X_{k}})=1$ if $h(v_{k})=1$ (thus $h(u_{k})=1\#$).
+By the argument from above, we now obtain an assignment that sets at least one variable per clause to $1$, satisfying $\varphi$.
+(Erasing ($r_{com^{*}}$): In this case, we can construct a simplified, but analogoue version of the reduction of the NE case.
+Assume $\varphi$ to be a formula of the form from above.
+Construct the word
+$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
+where, this time, $s_{i}=1\#1$ and $t_{j}=1$, for $1\leq i\leq m$ and $1\leq j\leq n$, and let
+$\alpha=\#\#\ u_{0}v_{0}\ \#\#\ ...\ \#\#\ u_{m}v_{m}\ \#\#\ \alpha_{1}\ \#\#\ ...\ \#\#\ \alpha_{n}\ \#\#$
+where $\alpha_{i}=x_{i,1}x_{i,2}x_{i,3}$ for variables $x_{i,1}$ to $x_{i,3}$.
+We set $r_{\alpha}$ analogously to the NE case.
+The first direction of the proof, assuming a satisfying assignment for $\varphi$, we can create a $r_{\alpha}$-substitution similarly to how it is done in the NE case.
+Set $u_{k}$ and $v_{k}$ analogously to the NE case.
+For each clause, choose one of the literal variables $x_{i,j}$ that is related to either some $u_{k}$ or $v_{k}$ substituted by $1$ and substitute it by $1$.
+The other direction of the proof follows by the same reasoning from the NE case that the ##\#\# factors have to align in $w$ and $h(\alpha)$, therefore forcing $h(u_{k}v_{k})=1\#1$ and $h(\alpha_{i})=1$.
+Hence, for each $i\in[n]$, some variable $x_{i,j}$ is substituted by $1$.
+This is only possible if either the related $u_{k}$ or $v_{k}$ is substituted by either $\varepsilon$ or $1$ due to the definition of $r_{com^{*}}$.
+From that, we can construct a satisfying assignment $\beta$ for $\varphi$, analogously to before, concluding this proof idea.
+Hence, the membership problem for all cases considered in Theorem˜5.1 is NP-complete, concluding this proof. ∎